A Bibliography of Publications in Science and Public Policy: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.03

Title word cross-reference

3 [Mah07].

19th [Joh00a].

2 [Cro07, LLS06]. 2001 [McN03]. 2002 [LE03]. 2004 [Ree05].

5 [PS05].

7th [DM07].

Aarhus [Rig07]. Abandon [Cha00]. abroad [Bie07]. absolute [Iai04]. absorptive [KL01]. academia [Per04]. Academic [Bra07, GH08, LLST09, Mor03, Fis07b, GG00, HJW03, KL08b, MR04, Pav00, Rod06, RH07, Sim04]. academically [JRKC08]. academics [VGP06]. academy [VC02]. academy-industry [VC02]. accept [Cha00]. acceptability [Hor08]. access [Fou01, dJS09]. accidental [Hei05b]. according [Ano01a]. accountability [Kul09, May03]. Accounting [Jas03a]. achievement [BS05a]. acid [Zeh05]. across [Lan05, OZ04]. acting [Dur07]. activitie [KMM09]. activities [BL07, GMS09b, GAP05, Lun07]. Actor [Lan02]. actors [MF08]. adapt [DW04, Lau06]. adaptive [Paa05]. Add [Kar02]. additionality [PS05]. administrations [Kin02]. advanced [Moe07]. advances [Gly01]. Advancing [AP08]. advantage [Kah06, ME00, Pap06]. advice [Cro03, FS02, MvZ01, Old06, Rav01, Wea01]. Advisory [Lei05, Buf01]. afar [Pre06]. affairs [Jas00]. afford [Col07]. Africa [Aye07, KH08, Lor09, Mab03, NM09, Pou05, TZ05]. African [Kah06, TZ05]. After [RD08]. again [Maz02]. agbiotech [LC05]. age [Axo00, Ray03, Res05]. aged [DM06]. Ageing [vdS08]. Agency [JCR05, Spi00]. agenda [CFBT^{+}06, Dal06, RH07]. agent [BG03, FC07, Gus03, Mor03, vdM03]. agents [Atl08, Cas03, Sho03]. Agoras [DL05]. agreements [Rod07]. agri [ARvT04, BG05, LCW05, Tof05, TB05]. agri-biotechnology [BG05, LCW05, Tof05, TB05]. agri-food [ARvT04]. Agricultural [LCS02, Car02, Dal06, KKA09, LM01, LOAJ02]. agriculture [JWHJ06, MR08, VCDM^{+}08]. Agro [Aru02, Men07, TCW02]. Agro-biotechnology [Aru02, TCW02]. agro-food [Men07]. agrochemical [BT02, GGB02, JL02]. ahead [vE01]. aided [WK01]. AIDS [Mah08]. Algeria [SZM08]. alignment [AD09]. alliance [SV00]. alliances [KY00]. allocation [PR07]. alternative [Smi03, Vog08a]. alternatives [BG05]. America [CT01, Tas02, Vel05]. American [Bru05a, Bru05b, CT01, Mil05, Ram08]. among [Ano01b, PR00, Swe07]. analyses [Lar01]. analysis [And06, Bor05, Coo07, DW04, DR01, GG06, Hou03a, Joh09, KKA09, Lep06, LHP04, McN03, Mic06, Uya07]. Analyzing [Pel06]. ancient [DL05]. Andersson [Jac03]. animal [KKBdCB07]. animals [Whi03]. ante [BC06, DM07]. anti [Sch02]. anti-modernization [Sch02]. anticipatory [Gus08, NGH08]. any [Iai04]. anybody [Mar03]. applicants [MD06]. applied [ACW06, Han06]. Applying [Fos04, Nie05]. approach [AC09, Ano00b, BL03, But07, Duf02, Edq04, Eki05, Gre05, GHB09, Hei08b, Koc08, Old08, Sim05, Thu07, Xie07]. Approaches [Mic06, Bal00, Ble05, CP06, MB08, Seg00]. approved [MD06]. Arctic [Mar04]. Area [Per02]. Argentina [Sch02]. arms [Ric02]. art [Lau06]. article [Sim04]. articulation [KO04]. Asia [Lun07]. Asking [McC07]. aspects [Wei03]. assess [VD07]. Assessing [Blo07, BFS09, BS05a, But07, FF03, Mah07, McN03, MR08, PTW08]. assessment [BA00, BA02, DM07, Don07a, EHvM06, Gru03, Gru06, GAP05, Han06, McN03, Mic06, Pou05, Ray03, SK01, Ree05]. assessments [BC06, Mat07, Pre06, Vog08b]. assisted [Etz06]. asthma [CFBT^{+}06]. asymmetries [Coo07]. Attempts [Dre02, Inz02]. attitudes [FS02]. audience [Lan03a]. Australia [Bel06, Ros07]. Australian [Moh02]. Austria [SLD07, TB05]. Austrian [Bie07, Kne08]. Authority [Fin00, Col00, KL08a]. autonomy [Koc08, Sch09a]. Avoiding [Geo07].

Back [Gel01, CCSM05]. bad [Bir01]. balance [Ram07]. balanced [But07]. balances [Mej09]. Balancing [Meh05]. Ballistic [Spi08]. Bangalore [Sco07]. based [AC09, BAV07, CBD^{+}05, Din07, Fin00, HJM07, Hér05, HH01, MH05, PR07, Spi08]. Basic [Mes00, AT01, SQ09]. Basin [RT06]. basis [Woo00]. battle [Maz02]. Battles [Mon04]. be [ACW06, Buc03, Gus08, LE03, McC03, Pav00]. beasts [SR00]. become [AP08]. bed [MPMA01]. behaviour [DW04]. behind [AFS08]. being [Moh02]. beliefs [GJ06]. benchmark [Tsi01]. Benchmarking [EB01, GFH01, Lar01, PRB06, PGS^{+}01, Kas01, Paa05, Sen01]. beneath [Geo07]. benefits [Rap08]. best [Kas00a, Kin02]. better [JRKC08]. Between [JL09, Len06, SLD07, CF08, Dyk04, GLR09, HvLS09, KK09, Keu04, LCW05, Ley01, Mej09, NM09, PR09, RL09, SW00a, Val02, Wat05]. Beyond [Leh04, SW00a, AFKTUO09, GMS09a, Per02]. Bibliometric [CP03, RG02]. bibliometrics [SS09]. bid [AP08]. big [Ast08, BD09b]. biodiversity [EW04, Mun00]. Bioethical [SQ09]. Bioethics [Lei05, Lóp04]. Biography [Ots06]. biological [EW04, Lit08, Rod07, Suk08]. biomedicine [Fil07, SQ09]. Biosciences [Coo04]. Biosecurity [MF08, Suk08, Vog08a]. Biosphere [Jap09]. Biotech [Sco07, Vog08a]. Biotechnology [CS04, ARvT04, Aru02, BS00, BT02, BG05, BS05a, CB04, Coo07, EvdGK04, Gan04, GGB02, HS07, JWHJ06, KKA09, KKBdCB07, LRS04, LM01, LOAJ02, LCW05, Man04, Men07, PF06, Rap08, Ras03, RHL04, TCW02, Tof05, TB05, WR01]. biothreat [Atl08]. bioweapons [Vog08b]. birth [JWHJ06]. black [Gon04, Lor09, THP08]. Blood [Bro01]. blue [Rus06]. Blurred [Edl04]. Bodies [Res05]. Book [Hol06, Pro06]. books [Meh05]. borders [Edl04]. both [Wei03]. bother [LF03]. Bottom [Ble05, FF08]. Bottom-up [Ble05, FF08]. Boundaries [BA02, Hal05, JCR05, Lan05, Lya07, Ram05, Rod06, Swe07, Wat05]. Boundary [HJ03, KL08a, Lei05, PR09, Swe07, Tur09, Zeh05]. boundary-work [Swe07]. bow [AG01b]. box [Gon04, Lor09, THP08]. brain [Par02]. brands [ARvT04]. Brazil [FF03, LRdCT00, Pel09, SdV^{+}09]. Brazilian [MdM09]. Breaking [THP08, Hei08a, MdM09]. breast [CP07]. Breeding [BA00]. Bridging [Bor09b, MY07a]. Bringing [CCSM05, Sha09]. brings [AG00c]. Britain [Mit06]. British [Buf01]. Broadband [Fer07]. BSE [Buf01, FS02, MvZ01, Seg00]. Bt [Cho01]. Buckinghamshire [Fin00]. Building [Ade02, Die00, EZ08, KL01, Sal08, PPMP09]. Bulgaria [Sim06]. Bundestag [Gru03].

Burying [Dur07]. business

[DBFK07, Kar02, MJ01, MMvPdlP08].

bypassing [Tof05].

Cajal [CCSM05]. California [BK02]. Can [Lan07a, LD07, LE03, Tsi01, Wei03, Col07]. Canada [And06, BA00, Duf02, RG02]. Canadian [Dur06, HW05, Wil06]. cancer [CP07]. candidate [MW01]. cannot [Gus08]. capabilities [ARvT04, Har02, PRJN02]. capability [Ade02, MH05, PPMP09, SV00, THP08]. capacity [KL01, PTW08]. capital [Bru09d, Die00]. capture [PS05]. Carcinogens [Gus03]. care [DM06]. career [Ora04, Pri03]. careers [MR03]. Careful [Bor05]. Caribbean [Vel05]. Carnegie [Bru07]. cartographers [CP07]. Case [GMG09, AD09, AO09, AFS08, APS^{+}09, Bas00, BS00, CFBT^{+}06, Dal06, Die00, Fre07, FS02, Hei05a, IM09, Jas03b, KH08, Kin02, Kul09, Lar04, Lep06, LM01, MU09, MMvPdlP08, Mey08, MR06, Nis05, Rod07, SM04, Sch09b, She05, Spi08, Sta06, TK08, Tur09, TZC09, VCDM^{+}08]. cases [Han06]. catastrophe [Bee05]. Catching [Chu02]. causes [CP07]. ceiling [Rod06]. cell [Hor08, Kul09, LHP04, SQ09]. cells [Spe04]. census [BD09a]. centers [Fuk08]. Central [KY00, LMJ^{+}09, RL09]. centre [Bie07]. centres [Muc00]. century [Daw02, Joh00a, San07]. Chain [Moh02]. Challenge [Sha07, MH05, vE01]. challenges [GMS09a, HS08, Lar03, Mil05]. Change [JL02, BS00, ER08, JGB09, Keu03, LvdBD^{+}07, Mar04, MW01, Nie05, PR00, Sha03, Zeh05]. Changes [Mes00]. Changing [Bar07, Kri01, Lya07, Mej09, PR07, Sen01, Smi05, TLL07, Bun00a, GP08, HvLS09, Sen00]. Characterising [Buf01]. characteristics [ACW06]. Characterizing [BD09a]. charisma [Len06]. Check [BK02]. chemicals [ARvT04]. China [AP08, Cho01, FF03, HY09, HS08, SQ09, She05, TZC09, WK01, ZP08]. chips [Lip02]. choice [BG05]. Choosing [BL03]. citation [CA07]. Citizen [Nis05, PB09, BD09b, Dou04, Fin00, Mar04, Ras03]. citizen-based [Fin00]. citizens [CM02, Kim06]. citizenship [EW04, Ile04, Jas04, Mej09]. civil [FP09, KMM09]. claims [Swe07]. clarify [God03]. classification [BD09a]. clean [Ros02]. cleaner [Muc00]. climate [JGB09, LD07, Nie05, Zeh05]. cloning [BM06]. close [MLG06]. closer [OZ04]. Closing [Dyk04]. cluster [Wil04, Wil06]. clusters [Coo07, CSJP08, Uya07]. Co [MR08, RT06, BR00, Dur07, Tof05, Tsi04, Val02]. co-evolution [Tsi04]. Co-existence [MR08, Tof05]. co-operative [BR00, Val02]. co-option [Dur07]. Co-production [RT06]. coalitions [LRdCT00]. cocoa [AO09]. cogent [Kee01]. cognitive [GD06]. cohesion [Meh05, Tsi04]. collaboration [Ama02, Cso06, ER08, Jap09, MPMA01, OZ04, RP03, Rog06, Thu07, TZC09]. collaborations [AP08, CP03]. collaborative [GG00, RP03, Son03]. collection [Cro03]. Collective [New03, Paa05, CSJP08, Lar01]. combination [Moe07]. Combining [ARvT04]. come [vHG03]. Comment [Sim04]. Comments [HH01, Mir04]. commerce [Bru05a]. Commercialisation [EvdGK04]. Commission [Lei05, Cro03]. commissions [JWHJ06]. committees [Buf01]. Communicating [Bri06]. communication [Ben03, HS09, LT04, Par02]. communications [Fri08]. communities [Zit01]. Community [Rog06, Ste03, Gul05]. Comparative [Bar07, Zeh05, Coo07, Dyk00, KKA09, Lar01, LHP04, ME00]. Comparing [LvdBD^{+}07, LMJ^{+}09, Smi01]. comparison [CS04, Han06, Hei08a, HAPN09, Jap09, LWYP08]. competence [Bie07, Mej09]. Competition [MU09, Hem03]. Competitive [Die00, HHH08, Pap06, VCDM^{+}08]. competitively [DW04]. competitiveness [Tsi01, Tsi04]. complex [Hou03a, KK01, Rya08]. complexity [Koc04, Lan07b]. compliance [Sti05]. Comprehensive [Fla00]. computer [WK01]. computer-aided [WK01]. computers [Teu03a]. computing [Ora04]. conceive [McC03]. concentration [LHP04]. concept [FF08]. conceptions [Wat05]. concepts [Dem08, Edl04, God03, IR08, TB05]. conceptual [BC06, Lep07a, Seg00, Son03]. Concerned [Bel06]. concerns [FM06, Leh03]. Concise [Kee01]. conclusion [Els09a]. concrete [GD06]. conditions [Lau06, PGS^{+}01]. conducted [GAP05]. Conference [LE03, Moh02, Nis05, Sae02]. conferences [BD09b]. conflation [IM09]. conflict [LCS02, Men07]. Conflicting [Lip04]. Congressional [Pay02]. conjectures [MS01]. connections [Vis04]. Consensus [Moh02, Nis05, Sae02, TB05, JWHJ06]. consequences [KK09, PM09]. conservation [Tak01]. considered [LE03]. consolidation [GD09]. constitution [Bog06]. constitutive [Gru06]. constraining [Suk08]. constraints [Sen05]. constructed [Kah06]. constructing [MB08]. construction [Bru06b, HJ03, Nav07]. constructivist [Fis07b, Gre05]. consultation [Wei03]. consumer [Dou04]. consumer-citizen [Dou04]. consumption [Ile04]. Containing [Suk08]. contemporary [ER09, Gan04, GD09]. contention [JWHJ06]. contested [Sco07]. Context [Bru05a, Gib00, WMWW06]. context-sensitive [Gib00]. contexts [Ley01]. Contextual [Phi02]. continuity [JL02]. Contract [SW00b, HvLS09]. contracted [Koc08]. contracts [Cas03]. Contradictory [FM06]. contrasting [LLS06]. contribute [CSJP08]. Contributing [Sti03]. Contribution [AT01, Kas01, EHvM06, FSWR00, Tsi01, Wil06, Wol07]. contributions [MS01]. contributors [Ano05a]. Control [Dur07, Atl08, GS03, Gug06, Lar08]. Controlling [Sti05]. controversy [LOAJ02, MJRB05]. Convergence [PR00, ERB^{+}08, KKA09]. Conversations [Lan05, LDDM08]. Cooperation [Dra09, HH01, Gün04, Inz00, Sta06]. Cooperative [LRdCT00]. Coordination [EK08, Bra08a, Bra08b, GB08, PTW08]. COPD [CFBT^{+}06]. coping [MR06]. coproduction [Sch09b]. Coronary [Buf01]. cotton [Cho01, TK08]. Council [Pel06, Sch09b, Gul05, SLD07, vdM03, vHG03]. councils [Sko00a]. counseling [Bir01]. Countering [Len08]. countries [Ade02, CS04, HAPN09, Kim00, LMJ^{+}09, MW01, ME00, SZM08, She05, Sko00b, Tho00, dJM01]. Country [SM04, Jas03b, Mey08, Pov02, Sha09]. counts [CA07]. coverage [Fla00]. craze [Wil06]. Creating [Joh00a, Lew02]. creation [Pre06]. creationists [Maz02]. Creativity [DAZ09, KI09]. creativus [KI09]. credits [Hei05a]. crime [BM06]. criminal [Col07]. crisis [ER09, Kah06, Seg00]. Critical [Wat00, Wol07]. critique [AH03, Jac03]. crop [Ros07]. crops [BA00, ML06b, Nis05, Ore05, Sch05]. Cuba [OJ09]. Cultural [Mab03]. Culture [Cha09, Bra07, DAZ09, RD08]. cultures [GH08, Kri01]. current [Cha00, EB01, HS09, Sko00b]. customer [ACW06]. cyber [Kee03]. cyber-goddess [Kee03]. Cyberspace [Bel02]. Cyborg [Kee03]. Cyprus [MH05].

D [God06]. Dale [Bru07]. dangerous [Mun00]. dangers [Jak02]. Danish [GLR09]. Data [Jak02, PR07]. database [LLST09]. day [Gha09]. death [Vog08b]. debate [DL05, Fou01, FSWR00, GJ06, Hor08, Kas00b, KKBdCB07, SH04, Sti03, Wei03, vdS08]. decade [MR04, Sim04]. decentralised [Pov02]. decision [CM02, Fou01, Ray03, Ste03]. decision-making [Fou01, Ste03]. decisions [Phi02]. decreasing [PR00]. dedicated [EvdGK04]. deep [Inz00]. Defence [JCR05, Spi00, Spi08, Spi00]. defense [Whi03]. defined [Daw02]. Defining [BA02, GHB09, O’R09]. definition [DFG^{+}03]. Delegation [KL08a, Bra03]. deliberation [KKBdCB07]. deliberations [Nis05, Zeh05]. deliberative [Fin00, Fou01, vE01]. Delivering [Gum01]. Demand [KO04, AD09, EHvM06, Fou01]. Democracy [Ray03, Fee03, Fou01, Gru03, Leh04, LF03, Mil05, vE01]. democratic [GS03, Wea01, Sae02]. democratisation [dJM01]. Democratising [LF03, Now03, Gru03, MY07a]. Democratizing [Skl04, Chr03, TW03]. Denmark [LLST09, Mej09, SGL03, Tof05]. Department [Fre07]. derivatives [ML06b]. Deserves [Lan03a]. design [WK01]. Designing [Mah08, BC06]. despite [TB05]. Devastating [Sla01]. developed [Ade02, Kim00]. Developing [GG06, Gre05, SZM08, Sha09, She05, Wei03]. Development [Cho01, MD06, TZ05, CL03, CBD^{+}05, Die00, DBFK07, DFG^{+}03, FM06, GD06, God06, HTCR06, KH08, Mab03, Mic06, MY07a, MY07b, MD09, OJ09, PR09, Rue07, Smi03, Tsi01, Voß07, HH01]. developmental [Gun06]. Developments [SK01, Blo07, EB01, Sha06b]. device [Gol03]. dialogue [Dou04]. Diana [Ora04]. Differences [KK09]. different [ARvT04]. differentiation [PR00]. difficulties [Lov07, Tsi04]. Diffusion [KWC08, WK01]. Dilbert [Bru07]. dilemma [She05]. dimension [Per02]. Dimensions [Uya07]. direct [CF08, Dyk04, Inz00, PPMP09, WR01]. directions [TZ05]. disappointing [Tur07]. disarmament [Lit08]. disciplinary [KK09]. discipline [HJW03]. disciplines [Fis07b, LLS06]. discourse [HJ03, Sti05]. discoveries [Hei05b]. discussion [SR00]. Disease [BM00, Buf01]. disenfranchised [Whi02]. disparity [Val09]. dissemination [KK09]. distinct [Edl04]. Distributional [CBD^{+}05]. Distributive [Coz07]. disturbing [HW01]. divergent [TB05]. diverse [HJM07]. Diversification [Spi00]. Diversity [Sko00a, HHH08, Smi01]. dividend [TZC09]. DNA [Col07]. Do [EvdGK04, Pay02, Bru06a, FP09, Moh02, Swe07]. do-it-yourself [Bru06a]. Does [CSJP08, HHH08, Sri06, FW04, Gol03, LF03, MH05, Tho00]. Doing [Ast08]. domestic [Duf02]. dominant [GJ06]. Don’t [AG02, Bel00a]. drain [Par02]. drama [Bel01]. drawbacks [Sri04]. Drawing [Cro03]. driven [Lar08]. driving [AFS08]. drought [Bel06]. drowning [SS08]. drug [Cro07, Rap01]. Drugs [Bel00a, ARvT04]. Ds [Mah07]. dual [Atl08]. dual-use [Atl08]. Duality [Bas00]. dynamic [THP08]. dynamics [For04, GD06, LNB06, Mey08, SGL03]. dynamism [IM09]. dystopian [Tay02].

early [IR08, MR03, MvZ01]. earmarks [Pay02]. East [Etz00, Gün04]. Eastern [Dyk04, KY00, LMJ^{+}09, RL09]. easy [Lan04]. ecological [Tur09]. econometrics [Ste09]. Economic [Dyk00, HH01, AFKTUO09, CBD^{+}05, Cun08, ER09, FF03, Kas01, MR08, MW01, MD09, OJ09, SZ02, Smi01, Smi05, TA01, vTN04]. Economics [Lan03b, AC09]. economies [Dyk00, HH01, Mes00]. economy [Din07, Eub07, HY09, Hem05, Hér05, HS08, KH08, MdM09, MH05, Phi09, Rav01, SQ09]. edition [Blo07]. Editors [Ano05a]. education [BL07, GMS09a, Hei08b, HHH08, Kid04, Lar03, LRdCT00, McN03, MW01, OJ09, PTW08, RD08, Sko00b]. effective [AD09, Hou03a, Sal08, Sae02]. effectiveness [Nie05, Tur09]. effects [CBD^{+}05, MR08]. efforts [And06]. eight [MD06]. Einstein [San07]. electrical [Fri08]. electron [Ots06]. Electronic [Fin00, RG04]. elements [DAZ09]. elusive [Wag02]. emancipation [MJRB05]. embeddedness [Thu07]. embryonic [Kul09]. emergence [Gib00]. emergency [Gre05]. emergent [RT06, Son03]. emerges [EHvM06]. emerging [KH08, PRJN02]. Emigration [CL03]. emissions [Voß07]. emphasis [Jac02]. Empirical [MR04, SZ02, BH06, FR09, HS09, PR07]. Empiricism [Kas00a]. employment [Aru02, WR01]. enabler [Cha09]. encounters [MLG06]. encourage [HHH08]. Encouraging [Bru06a]. endeavors [KL08b]. Energy [WMWW06, Joh09, Joh00a, NHS08, Ros02, Wei02]. engage [JRKC08]. engagement [RG04]. Engaging [Bal06]. engine [TA01]. engineered [BA00, TK08]. Engineering [Rap03, LRdCT00, Lóp04, Por03, Sri06, Sto08]. enhance [Nie05]. enhanced [Hei08b]. enlargement [MW01, vTN04]. enterprise [Huf04, LWYP08]. enterprises [BAV07, GGB02, HY09]. Entertaining [Bel01]. enticed [Bir01]. entity [Cre07, Hou03a]. entrepreneur [KGR07]. entrepreneurial [EZ08, ERB^{+}08, GH08, LWYP08, Rod06, ZP08]. entrepreneurship [Bra07, Hem05]. entry [HY09, Lar01]. environment [DAZ09, Kim00, KGR07]. Environmental [BA00, EW04, BK02, Ble05, EHvM06, ER08, HS09, Jap09, Leh03, Mat07, Ren01, Skl04, Sla01]. environments [Nis05, SGL03]. envisage [Fon07]. epidemic [Mah08]. epidemics [Mah08]. Epistemic [Zit01, LDDM08]. EPO [Sch09a]. equal [Col07, Teu03b]. equipment [GFH01]. Equity [LD07, Dow05]. ERA [Bor05]. erode [PM09]. essay [Hal06, Huf04]. establishments [LCS02, Spi00]. Estonia [MU09]. ethical [Sch05]. ethics [BM06, Cox04, Fis07a]. ethnographer [Ora04]. ethnographic [JG09]. ethos [Maa03]. EU [BD09b, EHvM06]. Europe [BK09, CB04, Dyk04, FLS04, For04, Gon04, HS09, Kas01, KY00, Lar01, LM01, LCS02, Man04, Mid02, MWC08, OZ04, RL09, Sch09a, Sen01, TCW02]. European [ACM05, Bon07, BD09a, Bor04a, Bor04b, Car02, Chr03, CP06, Coo07, Cro03, DM07, Els09b, Els09a, Etz00, Eva06, FSWR00, vHG03, GAP05, HS08, LRS04, LMJ^{+}09, LOAJ02, LCW05, LC05, Luu01, Luu07, MPMA01, MMvPdlP08, MW01, OV05, Pap06, Pav00, Per02, PR00, Rap01, Ree05, RHL04, SS08, SW00a, Sen00, Sha06a, SSA05, Ste02b, Ste04, Tsi04, VC02, Von00, Zit01, dJM01, vTN04]. Evaluating [CA07, Fuk08, OV05, TK08, BL03, BC06, God03, HJM07]. Evaluation [JCR05, RG02, ACM05, BK02, Bre07, Don07b, Fel07, Geo07, Kat00, KK09, Mic06, Moe07, Ree05, Rue07, Sae02, Sch05, THP08, VSZ07]. evaluations [Lan02]. everyone [Hol08]. everywhere [Ric05]. Evidence [Cro07, Fuk08, Kim00, SdV^{+}09, VSZ07, BH06, LLST09, WK01, RH07]. Evolution [Lan06c, TCW02, FMV09, LvdBD^{+}07, Man04, Tsi04]. Evolutionary [Eki05, VC02]. evolutionists [Maz02]. evolving [MR06]. Ex [DM07]. Ex-ante [DM07]. examination [JG09, Sha06b]. examined [Ora04]. Examining [Ora04]. example [WR01]. excellence [HBHN00, LNB06, RL09]. exception [Pel09]. exceptionalism [Mit06]. Exercise [McN03]. existence [MR08, Tof05]. existing [Seg00]. expect [FP09]. expectations [BH06, Bor09b]. experience [BR00, Chu02, Cro03, DM07, JGB09, Lit08, MWC08, Old06]. experienced [Wri03b]. Experiences [KL08a, Ren01, Gün04, SZM08, KMM09]. experiment [HW05]. Experimental [Die00]. experimentation [Kas03]. experiments [ML06b]. Expert [LLS06, Buf01, Cro03, Gus08, MvZ01]. expertise [Chr03, Dur06, Fau01, Gru03, Jas03a, LCW05, LC05, LF03, MJRB05, Now03, Ray03, TW03]. Expertising [Gru03, LF03]. Experts [Bed02, Dem08, RH07, Wei03]. Explaining [Bon07]. exploit [EB01]. Exploitation [TS01]. Exploration [Ber05, PR07]. Exploring [Eub07, IM09, KKBdCB07, Nav07, O’R09]. explosive [Fos04]. Export [PPMP09]. extension [GMS09b, NM09].

face [Bun00a]. faced [Sha07]. faces [MR03]. facing [Lar03]. fact [Hea01]. facto [CT01]. factor [KO04]. factory [GP08]. Faculty [Sko00b]. failure [BS05b]. Fairer [Kin06]. familiarity [BA00]. Fascinating [Lan04]. faster [Kin06]. favoured [APS^{+}09]. federal [BW01, EK08, FM06, Rue07]. feedback [Ros07]. few [PPMP09]. fiction [MB08]. fiction-inspired [MB08]. filled [Gus08]. finance [VCDM^{+}08]. financial [ML06b]. financing [BAV07]. Finding [Sti04, MPMA01]. Findings [MS01, Moh02, PS05]. Finland [Pel06, PTW08, SK01]. Finnish [Ras03]. firms [DG01, Hei05a, Lor09, MWC08, Oss04, Sha07, Val02]. first [BD09b, MR04, Moh02, Ros07, Sae02, Sim04]. Five [Ree05, GAP05, HAPN09, Sko00a]. Five-Year [Ree05, GAP05]. floor [GP08]. flows [GD06]. fly [BD09b]. Focal [Bre07]. focusing [WT03]. Food [Moh02, ARvT04, And06, Men07, Ros07]. forces [AFS08]. forecasting [LD07]. Foreign [Ama02, CF08, Inz00, Dyk04, Mil05, NS06, PPMP09, Ste02b, Wag02]. Foreign-owned [CF08]. foreknowledge [Mah08]. Foresight [LT04]. fork [Duf02]. form [Son03]. formation [SV00]. Forsythe [Ora04]. Forum [PB09, HS08]. forums [Pre06]. foundation [Hem05]. foundations [HJW03]. Four [Har02, Len06]. fourth [LE03, LE03]. FP5 [GAP05]. fragile [Sti05]. fragmentation [EK08]. Framework [ACM05, DM07, Luu01, OV05, BC06, DL05, DR01, HJM07, Han06, Har07, LWYP08, Mab03, Mah07, PGS^{+}01, TZ05, THP08, VD07, GAP05, PS05]. frameworks [KKBdCB07, Lar04, TK08]. Framing [Vog08a, Dou04]. France [BL07, Fou01, MR08, OZ04, TLL07, Wei03]. French [MR03, MJRB05]. friendly [Bru09a]. Friends [AG00a]. Frontiers [VSZ07, Fel07]. fuel [LHP04]. fun [Bru07]. funded [DW04, DBFK07, Lau06]. funding [DLR03, Hei08a, HAPN09, HHH08, JL09, KL08a, Lau06, Lep06, Lep07b, LMJ^{+}09, MU09, Pav00, PR07, RL09, Sla05, TLL07]. funds [VCDM^{+}08]. fusion [LHJOS08]. Future [Don07a, Fro03, Des07, Don07b, Gel01, Moe07, Mun08, NGH08, PB09, Rob07, Sko00b, Tay02, TA01, Wri03a]. Futures [RT06, Bee05, Gus08, MB08, Vog08b].

Gambling [Bee05]. game [DW04]. Gandhi [Old08]. gap [Dyk04, NS07]. gaps [LRS04]. Gene [Moh02]. generation [KI09]. generative [Gun06]. genes [Teu03a]. Genetic [AG00b, Mit06, Fro03, Lóp04, Sri06, Sto08]. genetically [And06, BA00, BA02, MR08, ML06b, Pel09, Ros07, TK08]. Genetics [Cox04, Kid04, LDDM08, Mit06]. genomics [DDS05, PM09]. geography [Pov02]. Georgia [RT06]. Geosphere [Jap09]. Gerais [SdV^{+}09]. German [Coo07, EK08, Gru03, KMM09, Spe04, WR01]. Germany [BG05, FP09, Gün04, KL08b, MR08]. Get [Axo02, Mat07]. Getting [DDS05, Els09b, Els09a, Sha04, Lau06]. Global [Mar04, Mat07, AP08, Atl08, Cun08, ER08, EZ08, ERB^{+}08, GD06, HS08, IR08, Jap09, KKA09, Mah08, SQ09, Sen03, Sri04, Sta06, Ste02b, Zeh05]. Globalisation [Ste02a]. globalization [Ric06]. globalizing [Mil05]. globally [Dur07]. GM [MJRB05, Nis05, Ore05, Sch05]. GMO [Aye07]. GMOs [FP09]. go [Maz02]. goal [KL08a]. goddess [Kee03]. goes [Bog06]. gone [Por03]. Good [Kin02, Bir01]. got [God06]. Governance [EK08, Aye07, BH06, De 03, DR01, DDS05, FR09, FSWR00, Gon04, HS07, JN03, Koc08, LM01, Lov07, Luu07, LT04, Lya07, O’R09, SQ09, Suk08, Voß07, Whi07, Zit01]. Governing [BH08, Fis07a, MY07a, Sch09a, Win07, ML06a, MY07b, Pel06]. Government [Sch09b, BH08, Etz02, FM00, GG06, GMG09, Kim00, LE03, RG02, RP03]. governmental [Dou04, PEP04]. grant [MD06, dJS09]. grants [Ito04, VGP06]. Greece [Ama02, APS^{+}09, BS05a, CP06]. green [Bel06, LHP04, Kas03]. Greening [Bun00b]. ground [Hei08a, MD09, Spi08]. ground-based [Spi08]. ground-breaking [Hei08a]. groups [GH08]. grow [Lan07a]. growing [Fou01]. growth [PPMP09, TA01, Tsi01]. guidelines [Cro03]. guts [Bro01].

habitats [BS09]. hand [Etz06]. happy [Sto08]. Hard [Wei02]. Hard-to-shake [Wei02]. harmonization [Atl08]. Hayles [Lin07]. Health [Fin00, CFBT^{+}06, Geo04, JG09, PM09, Skl04, Sti05]. healthcare [FM00]. heart [Buf01]. heavens [Teu03a]. helix [LE03, EKHMW01, Etz02, EZ06, SZM08, VC02]. help [Wei03]. hemisphere [Ano00b]. Henrekson [Jac03]. Here [Maz02]. heritage [JL09]. heuristics [Len06]. Hidden [Sen05]. high [AG00a, OJ09]. high-tech [OJ09]. Higher [OJ09, PTW08, BL07, HHH08, Lar03, Sko00b]. higher-education [BL07]. historical [Fri08, JL09]. Historiography [Ano01a]. history [Axo02, Ber05, Chi01, MvZ01, Sha06a, Vis04]. Holistic [Sim05]. home [Fon07]. Homo [KI09]. Hong [Sha06b]. hopeful [Euc06]. horizontal [Man04]. house [ARvT04, Rig07]. hubris [Cox04]. human [CCSM05, Eki05, Fis07a, JWHJ06, Lóp04, Mah07, ME00, Sen01]. humanities [Teu03b]. Humboldt [SW00a]. Hungarian [Inz02]. Hungary [Inz08]. Hydrological [Jap09].

idea [vHG03]. ideas [Wei02]. Identification [Hei05a]. identity [Sko00a, Spe04]. If [Fre01]. illness [Sti04]. images [Mon04]. imagined [Vog08b]. immature [SdV^{+}09]. immigrant [Sch09b]. immigrants [Spe04]. immigration [Har07]. Impact [GG00, BFS09, Col07, DM07, Don07a, Els09c, GAP05, Lar08]. impacts [FF03, WR01]. Implementation [SSA05, PGB09]. Implementing [Fin00, Ren05]. implications [Bal00, Sen01]. importance [Tur07, TZC09]. important [Buc03]. Imposing [Gan04]. imprecision [HSA01]. improve [And06]. improving [Hic09]. in-house [ARvT04]. inadequate [Lan03b]. inclusive [Ste03]. increase [Pay02]. increased [Pav00]. Increasing [dJS09]. Incubation [Etz02, MJ01]. incubators [Etz02, MJ01]. independence [Gru06]. independent [Kat00]. Index [Ano00c, Ano01c, Ano02, Ano03, Ano04, Ano05b, Ano06, Ano07, Ano08, Ano09a]. India [Bas00, Cho01, KWC08, Kri01, Mah08, Par02, SV00, TK08]. Indicator [PR09]. Indicators [BL07, Bar01, HJM07, Kat00, Lep07b, Tur09]. indigenous [Kim00]. indirect [WR01]. Individual [Col00]. Individualized [Fis07a]. Industrial [FM00, AH03, CF08, EB01, Jac02, Par02]. industries [BT02, Eva06]. Industry [LOAJ02, PEP04, AD09, AO09, BW01, Bun00b, CP03, Etz02, Fil07, GP08, JL02, LE03, MS01, Muc00, PGS^{+}01, RP03, SV00, SdV^{+}09, TCW02, Thu07, VC02, Fre07]. industry-science [PGS^{+}01]. inequalities [Fis07a]. inequality [Eub07, Sen03]. inequities [WS07]. inequity [Wet07]. Inertia [BS00]. Influence [Bru02, HAPN09, Jas03b, Joh00c]. influenced [Sla05]. influences [GGB02]. influencing [BT02]. inform [HS09]. informal [Kwa06]. Information [Fou01, HTCR06, Ben03, Cso06, Eub07, MPMA01, Par02, Res05]. Informing [Hic09]. infrastructures [Smi05]. Initiative [GMG09]. Innovation [Car02, CP06, Eva06, Fau01, Gün04, HW05, Lar04, Rob07, SS08, SGL03, Tsi04, Voß07, Ade02, AO09, Ano01b, APS^{+}09, AT01, Aru02, Bas00, BS09, BT02, Blo07, Bor04a, Bor04b, Bra08a, Bra08b, Bru06a, Bru09a, CS04, CB04, CF08, CSJP08, DG01, Edq04, Etz02, Etz06, EZ06, ER09, FMV09, FR09, For04, Fre07, FM00, Fuk08, GG06, Gon04, GGB02, GB08, Gun06, HJM07, Har02, Har07, Hem03, HBHN00, Hic09, HTCR06, Inz02, Inz08, IM09, Jas03b, Kah06, Kin02, Koc08, Lar03, Lor09, Lov07, Lya07, Mes00, NHS08, Paa05, Pap06, Pel06, Pov02, Rol09, Rub06, SH04, SM04, Sha06b, Sha09, Sim05, Sim06, SdV^{+}09, Val09, VCDM^{+}08, WT03, WMWW06, Wol07, Xie07]. innovations [Fri08]. innovative [Sha07, THP08]. innovator [Bru09d]. innovators [Oss04]. ins [MdM09]. insiders [Mas01]. Insights [BL07, BK02, Lun07]. Inspiration [Ano01b]. inspired [MB08]. Inspiring [Lun07]. institute [HJW03]. institutes [BL07, PR00]. Institutional [Ram05, Ram08, JG09, PR00, WMWW06, Zei04]. Institutionalising [Dou04]. Institutions [Bon07, Hei08b, Rol09, Sal08, Vel05]. instrument [Paa05, Voß07]. instruments [Fre07]. insurance [Bee05, Mit06]. Integrated [Koc08, EHvM06, Pel06]. integrating [FM06]. integration [LT04, OZ04, Sch09b, Ste02b]. Intellectual [Cun08, Des07, Haw08b, Rod06, She05, Win07]. intelligent [Moe07]. intent [FM06]. inter [OZ04]. inter-regional [OZ04]. interaction [Ste02b]. interactions [Kwa06, SdV^{+}09]. Interactive [Bal00, Orm00, CS00a, CS00b, RT06, SR00, SW00b, Wat00, Woo00]. interdisciplinarity [Len06]. interdisciplinary [Koc04, Kwa06]. interest [Hei05a]. interface [Bor09a, JGB09, Rig07]. Intermediation [vdM03]. International [Jap09, Per02, Ama02, AP08, Ble05, Chu02, Cun08, Dal06, Dem08, Duf02, EB01, ER08, Gul04, OZ04, TZC09, Jap09]. internationalisation [AFS08]. Internet [Fin00, Sha06a]. interpretation [DAZ09]. interpretative [Jac02]. intersections [PPMP09]. intervention [Lar03]. interventions [HJM07]. Introducing [CS00a, Rya08, Muc00]. Introduction [Bor04a, CB04, Car02, Don07a, EZ08, GMS09a, Lep07b, MY07b, NGH08, Sen00, Suk08, VSZ07, Wet07, BS09, BG03, Els09b, RL09]. Inventions [Fri08, dJS09]. Inventive [LHP04]. inventors [GH08]. invents [Fil07]. investigation [FR09]. investment [CF08, Dyk04, Inz00, Kim00, PPMP09, Tsi01]. investments [VD07]. invisible [Vis04]. involvement [Inz08, Sko00b]. IPR [Sri04]. Iran [GG06]. Iranian [GMG09, Gha09]. Iraqi [Por03, Vog08b]. Ireland [CS04, CP06]. Irish [AT01]. ironies [Chi01]. irresponsibility [Hea01]. Islam [Huf04]. Israel [PF06]. issue [Bor04a, CB04, CS00b, Els09b, Els09a, EZ08, FP09, Lep07b, MY07b, Sen00, Suk08, VSZ07, Wet07]. issues [Fos04, Kas00b, SZ02, Tof05]. Italy [AD09]. ivory [Ano01b, GP08].

Jacobsson [AH03]. Japan [FM00, Fuk08, Ham06, IM09]. Japanese [DM06, Ito04, KO04, Nis05]. joint [CVI02, vBE04]. Joseph [Haw08a]. judgment [Gus08, RH07]. Jury [Fin00]. just [Bru09c]. justice [Col07, Coz07, Jen00, Tak01].

Katherine [Lin07]. KEINS [LLST09]. Kenya [HS07]. key [DAZ09, KO04]. Keynesian [ER09]. Kingdom [Dre02, Ore05]. know [Axo02, Leh04]. knowledge [ARvT04, Bal00, BAV07, Bra08a, Bra08b, Bru06b, Cro07, Des07, Din07, EK08, GD06, GB08, GS03, Gus08, HY09, Hei08b, Hér05, HW05, HS08, KWC08, LLS06, Maa03, MH05, NGH08, Now03, Pre06, Rav01, Sha03, Smi05, Ste04, vBE04]. knowledge-based [BAV07, Din07, Hér05, MH05]. knows [AG00c]. Kong [Sha06b]. Korea [Kim00]. Korean [BM06, Chu02]. Kostadinka [Mir04]. Kplus [Bie07]. Kuhn [Ram01].

lab [KKBdCB07]. laboratory

[Bog06, BW01, Hor08]. Laboring [Bru09b]. labour [KWC08]. Lack [Ram07, Gus08]. landscape [Smi05]. large [McC07]. Lasting [Bra03]. later [San07]. Latin [Vel05]. Latvia [AFKTUO09]. Lay [LHJOS08]. lead [AC09]. leader [AP08]. Learning [Abe05, Bie07, Lor09, CSJP08, JL09, KL01, Paa05, Rog06, SW09]. led [AP08, MJRB05]. left [Buc03, Sch02]. legacy [Haw08a]. legal [Chr03]. legalisation [JN03]. legislation [FM06]. legislators [Kas00b]. legitimation [Aye07]. leisure [KMM09]. Leo [Lan06c]. less [Ade02, APS^{+}09, Kim00]. Lessons [Bra08a, LM01, MvZ01, MWC08, GD06, IR08, Old06, PB09, FP09, She05]. level [CBD^{+}05, DR01, FR09, LT04, Lya07]. Liberalisation [Smi05]. lies [Geo07, VSZ07]. Life [RD08, Bee05, Car02, Gha09, Haw08a, Len08, MR03, MR06, TCW02]. light [DFG^{+}03, Eva05]. like [Wil04]. limitations [GKZ09]. limits [JN03, VCDM^{+}08]. linear [Etz06]. lines [Keu04]. Linkages [KMM09, Chu02, NM09]. Linking [GLR09]. links [LCW05]. Lisbon [Lar04]. List [Ano09b]. Listening [LDDM08]. literature [BFS09, Fil07, Jac02]. little [Mun00]. lively [Lan06a]. Local [GD06, KKA09, Fuk08, Kin02]. locally [Dur07]. lock [MdM09]. lock-ins [MdM09]. long [Lep06]. long-term [Lep06]. longer [MB08]. look [OZ04]. looking [Har02]. losing [AC09]. low [Hei05a]. Luxembourg [Mey08].

Machinic [Bel00b]. Macro [HJM07]. made [Vis04]. main [PS05]. mainstream [Sha09]. Mainstreaming [APS^{+}09]. maize [MR08]. major [Dra09, Lar03, MMvPdlP08]. majority [Whi03]. makers [LDDM08, NS07, PR09, Ver04]. makes [Gol08, Joh02]. Making [Gow03, Ile04, Bra03, CM02, EW04, EHvM06, Fou01, HS09, Inz08, Mah08, MJRB05, NGH08, Nie05, PRB06, Ray03, Ren01, Ste03]. Malaysia [SZM08]. Management [MR03, Wat00, Dur06, Dur07, HW05, HS09, JN03, KI09, Lov07, Ren05]. managers [JGB09]. Managing [Dur06, EKHMW01, Lit08, Spe04]. manifold [GMS09b]. Manual [Blo07, HH01]. manufacturing [HH01, TA01]. many [Mid02]. Mapping [Fel07, Hou03a]. maquila [PPMP09]. Marine [WMWW06]. mark [Jam04]. market [EHvM06, Kim00]. marriage [Sto08]. masculinity [Nav07]. mass [Sko00b]. masts [Geo04]. material [Rod07]. mathematics [Ber05]. matter [Duf09, EvdGK04, FW04, Tho00]. McLuhan [Gow03]. mean [LF03]. meanings [FF08]. measure [Lep07a]. measurement [Blo07, IR08]. measures [dJS09]. Measuring [PRJN02]. mechanism [Bre07]. Mechanisms [BAV07]. medical [Ram08, Sae02]. medically [Whi02]. medicine [Daw02]. medium [GGB02]. meet [GP08, MH05]. meets [EHvM06]. Member [LRS04, RHL04]. memory [New03]. mention [AG02]. merger [Sko00a]. merging [Ile04]. Merton [Rod07]. metaphors [Joh00b]. methodological [ACW06]. Metrics [Don07a, CA07, Moe07]. Mexican [CL03, VCDM^{+}08]. Mexico [CL06]. mezzo [DR01]. mezzo-level [DR01]. microdata [BD09a]. Microsoft [She05]. midcourse [Spi08]. migration [KWC08]. Minas [SdV^{+}09]. mind [Mon04]. ministers [Dre02]. miraculous [San07]. Mirskaya [Sim04]. missile [Spi08]. Missing [Jam04]. mission [PGB09]. mistakes [Abe05]. misuse [Len08]. Mixing [Joh00b]. MNEs [Lun07]. Mobile [Geo04, Sti05]. Mobiles [Ham06]. Mobility [NS06, Fon07]. Modal [AG01a]. Mode [Gib00, ML06a, Cro07, LLS06]. Mode-2 [LLS06]. model [Die00, Etz06, FC07, GMG09, God03, KL01, PPMP09, Pov02, SW09, Vog08a]. modeling [Lep07a]. modelling [AC09]. modelling-based [AC09]. models [AG01a, Bru09b, DG01, Mah08, PR07]. modern [DL05]. modernization [Sch02]. modes [LHJOS08, Rod07]. modified [And06, BA02, MR08, ML06b, Pel09, Ros07]. Monitoring [DBFK07]. moral [KKBdCB07]. motion [Res05]. motivation [IR08]. Moving [AFKTUO09]. Mrs [Old08]. much [Col07]. Multi [RP03, FR09, LT04, Lya07]. multi-level [FR09, LT04, Lya07]. Multi-sector [RP03]. multicriteria [BL03]. Multilevel [Hou03a]. Multinational [Gul04]. multinationals [ARvT04]. Multiple [JG09, GMS09b, KL08a, LHJOS08, TK08]. multiplicity [Fau01, Len08]. museum [Bro01]. must [Iai04, Ver04]. mysticisms [Bel00b]. myth [WT03]. mythical [SR00]. Myths [GJ06]. nano [IM09]. nano-innovation [IM09]. NanoI [IM09]. nanotech [AP08]. Nanotechnology [GMG09, AP08, BH08, FM06, IM09, Whi07]. narratives [Tay02]. National [Cun08, Hem05, Lei05, PB09, WT03, Ade02, AT01, Bas00, EB01, FMV09, GG06, GMG09, Gul04, Hal05, Har02, Har07, Kah06, LvdBD^{+}07, Mes00, Muc00, PRJN02, Xie07, GMG09]. nationalism [AG00b]. Nations [Old06, Sta06]. naturalists [EW04]. Nature [Tak01, Eki05, Gly01]. necessary [LWYP08, Muc00]. Need [Keu03, Joh09, May03, Par02, RD08, Sch09a]. needed [Kim00]. Needham [Ano01a]. needs [GP08]. neglected [CF08]. negotiation [DLR03]. neighborhood [Bru09a]. Nelson [Bru02]. Netherlands [CFBT^{+}06, Dre02, PM09, Sch05, Sch09b, vdS08]. Network [HW05, Hem05, LDDM08, Thu07]. networks [Blu08, Cro07, Etz02, Hei08b, KL08a, LRdCT00, LNB06, Lya07, MF08, Rya08]. never [VSZ07]. next [Lip02]. nexus [CF08]. niche [Lov07]. niches [Smi03]. Nigerian [AO09]. No [Jas03a, EHvM06]. non [Dou04, HH01, MF08]. non-governmental [Dou04]. non-manufacturing [HH01]. non-state [MF08]. nonlinear [ZP08]. nonsense [Bar01]. Normative [BH06]. norms [Bra07]. Norway [AFS08, FMV09, NS06, Sae02, Sko00a, SLD07]. Nuclear [Gus08, Dur06, Dur07, LHJOS08]. numbers [SS08].

objectives [Ano00a, BS05a]. objects [Tur09]. off [MWC08]. offices [Dra09]. official [Buf01]. offs [KL08b]. Old [Luu01]. One [Mid02, Bru07, Det06, Pro06, San07]. ones [Cha00]. opening [Gon04]. Operationalising [Han06]. operative [BR00, Val02]. opinion [Hou03b]. opinions [Lip04, Pro06]. Opportunities [Tsi04].

Opposite [Pro06]. optimal [Bre07]. option [Dur07, VD07]. Options [dJM01, BAV07]. Organics [Sto08]. Organisation [JN03, DG01, Dou04, Sen00, Sen01, HH01]. organisations [DW04, FP09, HJ03]. organised [Hea01]. Organising [Bra08b]. organism [BA02]. organisms [Pel09]. organization [KL08a, Lei05, LMJ^{+}09, Son03]. organizational [Son03]. organizations [BL07]. orientation [Von00]. Origin [Gal01]. orphan [Cro07]. Oscar [Sch02]. Oslo [Blo07, HH01]. ourselves [McC03]. outcomes [McN03]. outlook [Ren01]. output [CA07, Pay02, RG02]. Outsiders [Mas01]. oversight [Lew02, TK08]. overview [Vel05]. own [Wat05]. owned [CF08, HY09].

Painful [Ste09]. paper [AH03]. paradigm [Det06, Gol03, Lan03b]. paradigmatic [HJW03]. Park [APS^{+}09]. parks [BS09, KI09, MJ01]. parliamentary [Gru06, SK01]. Part [Bru07, Per04, Gru06, GAP05, HJM07]. Part-timers [Per04]. participant [Ora04]. participation [CFBT^{+}06, De 03, EW04, FF08, FP09, Mej09, Ras03, Ros07]. participative [Wat00]. Participatory [BH06, BC06, Han06]. particularities [Mes00]. partners [AG01b, MPMA01]. Partnership [TZ05, JCR05, Orm00, Wag02]. partnerships [BW01, SZ02]. passionate [Joh01]. Past [Des07, Mun08, TA01, Abe05, Old06]. Patent [BK09, Bru09c, Dra09, Els09b, Els09a, For04, Sch09a]. patented [HBHN00, dJS09]. patenting [Cre07, Els09c, LLST09]. Patents [KK01, LHP04]. paths [ZP08]. Pathways [ERB^{+}08, Don07a]. Patras [APS^{+}09]. Patronage [Orm00]. Pattern [Hei05b]. Patterns [DG01, ER08, LvdBD^{+}07, Mes00, TLL07]. payoff [FC07]. Peer [NS07, Don07a, Moe07]. penetration [Fer07]. people [DAZ09, Gum01, Kar02]. percent [MD06]. perceptions [Bor09b, LHJOS08, MD09]. perfect [Joh02]. perform [JRKC08]. Performance [IR08, RHL04, ACM05, Bon07, Dyk00, HAPN09, Mic06, SS09, SZ02, Smi01]. performers [Dur06]. perils [RG04]. periphery [CP06]. personal [Ros00]. perspective [Ast08, Cho01, Chr03, Dyk00, EZ08, Fri08, GD06, RP03, SGL03, Sch09a, TW03, Tod01]. perspectives [Bar07, Cun08, HTCR06, Ram05, Sko00b]. Pharmaceutical [PEP04, SV00]. phases [Kri01]. PhD [MR03]. PhDs [MR03]. Philippines [MJ01]. Philosophy [Hud00]. phone [Sti05]. phones [Geo04]. phones/masts [Geo04]. picking [ER09]. piracy [She05]. place [HS07]. places [AG00a]. Plan [PEP04]. planning [GMG09, HS09]. plant [GGB02]. plea [But07]. Plethora [Ano00a]. point [HY09, Lar01]. poison [Sla01]. Poland [JL09, Jas00, Kul09, Sim05]. policies [BT02, Bon07, Bra08a, Bra08b, Bra07, CVI02, CB04, CL06, CF08, EB01, EvdGK04, Fon07, GB08, Hei05a, Kas01, Kne08, LRS04, Lar01, LvdBD^{+}07, Man04, Paa05, PTW08, Pov02, SM04, WT03, WS07]. Policy [Ben03, GGB02, HS08, SSA05, dJM01, AFS08, Ama02, APS^{+}09, BAV07, BL03, Ble05, BFS09, BG05, BS05a, BS05b, Bra03, BG03, Cho01, CP06, Coo04, Coz07, CT01, Det06, DM06, Don07a, EK08, EW04, Els09b, Els09a, EHvM06, Etz06, Fre07, FS02, Gha09, GD09, GJ06, Gus03, HJM07, Hal05, Hem03, Hem05, Hic09, HAPN09, HS09, Hou03b, Inz08, JL09, JGB09, Jas00, Jas03b, Joh09, Kid04, KKA09, KL08b, Kri01, Kul09, Kwa06, Lar04, Lar08, LDDM08, Lep06, LCW05, LT04, Lya07, Mah08, MJRB05, Mil05, Mit06, Mor03, NHS08, NGH08, Nie05, NS07, Paa05, PRB06, Pel06, PTW08, Pre06, PR09, Ram05, Ras03, Ren01, Rol09, Rub06, SGL03, Sch02, Sch09b, Skl04, SLD07, Ste02a, Ste02b, Ver04, Von00, Voß07, Wag02, Wat05, Wea01]. policy [Wei03, Wei02, Wet07, WMWW06, Zeh05, vTN04, Sch09b]. policy-driven [Lar08]. policy-makers [LDDM08, NS07, PR09]. policy-making [Bra03, EHvM06, Inz08, Mah08, Nie05, PRB06]. policy-relevant [NGH08]. policymaking [Sal08]. political [Bir01, Bra08a, Bra08b, GB08, Lar08, TB05, vTN04, vdS08]. Politics [JN03, MvZ01, Bar07, Bel06, Bor09a, FS02, GD09, Lan06c, Mit06, Nut00, Sco07, Spi08]. Pollution [Swe07]. poor [Bon07, Jas00, LD07, Whi02]. Popular [Eub07]. popularization [Lan06b]. portfolio [Rue07]. portfolios [HJM07]. Portugal [CS04, GD09, Per02]. Portuguese [Fon07]. pose [Sri06]. positioning [KGR07]. possible [CP07]. Post [Sim04, HJW03, MR04, ME00, dJS09]. post-academic [HJW03]. post-grant [dJS09]. post-socialist [ME00]. Post-Soviet [Sim04, MR04]. postal [DG01]. posthumanity [Lin07]. Postscript [CS00b]. potential [Spe04]. power [Bun00a, Dem08, Lep07b]. Practice [Joh02, Bal00, Dow05, Kin02, Lov07, MY07a, Nie05, Rog06, Sko00b, TS01]. Practices [MLG06, HS09, KK09, Muc00, Vis04]. pragmatic [Bel06]. pre [Euc06]. pre-hopeful [Euc06]. precaution [BG05, Ore05, Sch05, TB05]. Precautionary [LC05, Chr03, LCW05, Mit06, TW03]. preliminary [BD09a]. presence [Smi01]. present [KO04, Mun08, Sha06b, TA01]. prestigious [MD06]. pretty [Bru09c]. Prevailing [Els09c]. prevention [CP07, PM09]. price [CL03]. Principal [BG03, Gus03, FC07, vdM03].

Principal-agent [BG03, Gus03, FC07, vdM03]. Principals [Cas03, Sho03, Dur06]. principle [Chr03, TW03]. Principles [Duf09]. priorities [GHB09, LCS02]. privacy [Dow05]. Private [Har02, Inz08, DBFK07, FR09, JCR05, Kim00, PRB06]. privatisation [JCR05]. privilege [LLST09]. Privileging [Hea01]. problem [Bra03, Pel06, Ras03, Sri00, vBE04]. Problematic [Gly01]. problems [Dur06]. proceduralisation [Gug06]. process [BM00, Bra05, Fin00, Hea01, Mit06, Ore05]. processes [DLR03, Voß07]. procurement [Rol09]. produced [Tas02]. production [Bal00, Cro07, LHP04, Muc00, RT06, vBE04]. productively [Leh03]. productivity [Bar01, Dyk04]. products [Har02, Pap06]. professor [LLST09]. profile [RG02]. Program [Die00, Jap09, BK02, FF03, IR08, Jap09]. Programme [ACM05, GAP05, Luu01, PS05, Bie07, CCSM05, MPMA01, DM07]. programmes [DM07, Muc00, Sho03, OV05]. programming [Kwa06]. programs [AP08, Fel07, Rue07]. Project [RT06, Lep07b, MU09, PR07]. projects [Bre07, Gug06]. Promise [RG04]. Promises [Sri04]. Promotion [BR00, PEP04]. properties [RT06]. property [Cun08, Des07, Haw08b, Rod06, She05, Win07]. proposal [Pav00]. Prospects [Spi00, GKZ09]. protection [Des07, Jak02]. Providing [Ros07]. provision [GFH01]. Provisional [Moh02]. Provocative [HW01, Bal06]. Public [BT02, BS05b, De 03, FR09, FS02, Kas00b, Lep06, RL09, Rip05, Rol09, Ros01, Ros07, Tho00, AD09, BL07, BS00, Dal06, DBFK07, Dou04, EB01, EvdGK04, Fel07, FF08, FC07, FLS04, Fou01, Fuk08, Han06, Hor08, Hou03b, JL09, JCR05, Kid04, Kin02, Kul09, Lar03, Leh04, Lep07b, LMJ^{+}09, LCS02, LE03, Man04, MU09, May03, Mej09, MD09, PRB06, PR00, PR07, PM09, Ray03, RG04, SGL03, Sen00, Sen01, Sti04, TLL07, VD07, Wea01, Wei03, Zit01]. public-research [BL07]. Public-sector [Tho00, BS00, Fel07, LCS02, Ray03, Sen00, Sen01]. publication [KK09]. publishing [Kin06]. pure [Nut00]. purity [Swe07]. pursuit [Joh01]. Puzzle [Koc04].

qualitative [Don07b]. Quality

[Pap06, Don07a, Gug06]. Quantitative [Rue07, Lep07a, Pou05]. Questions [Fee03, McC07].

r [MLG06]. R&D

[PEP04, Ama02, Bre07, Fel07, FW04,

God06, Inz00, Kim00, MH05, SV00, Son03]. Rabkin [Sim04]. radical [Pav00]. rain [Zeh05]. Ramón [CCSM05]. Random [CM02]. randomisation [Bre07]. rationalisation [EHvM06]. rationalities [JG09]. Re [Ley01]. Re-thinking [Ley01]. reach [Axo00]. Reaching [Tra09]. read [Lan04, Lan06a]. Reading [Teu03a]. Real [VD07, Hol08]. realities [CP07, MD09]. reality [GJ06]. realized [Vog08b]. reciprocal [MJRB05]. reconfiguration [KO04]. Reconstructing [Sri00, HJW03]. record [ACM05]. recording [EW04]. reducing [WS07]. Reengineering [LRdCT00]. referees [Ano09b]. reference [McN03]. Reflections [Bru06b, Edq04, Ray03]. reflexive [McC02, Sti02]. reform [BK09, Dre02, JL09, Ram08]. regard [Teu03b]. regime [Sri04]. regimes [Bar07, Smi03]. region [KL01]. regional [Coo04, Fuk08, GHB09, Gun06, Lar04, OZ04, THP08]. regions [APS^{+}09, VC02]. Regulating [Mit06, Sch05, Wri03a]. regulation [And06, BK02, Cun08, HS09, JG09, LCW05, LC05, Lóp04, Pel09, PF06, Rap01, Ros07, Tof05, TB05]. Regulatory [ML06b, DR01, FS02, Len08, Phi02, Sch09a]. reinforced [GD06]. rejected [MD06].

Relations [Teu03b, GD09, Ley01, LE03, PGS^{+}01, VC02]. relationship [Gul05, SW00a, vdM03]. relationships [Val02]. relevance [HvLS09, NS07, RL09]. relevant [NGH08]. renewable [NHS08]. renewed [Lar01]. repel [Ros02]. Reply [Mir04, vE01]. Report [Gus03, LE03]. reproduction [Nav07]. requirements [GMS09b]. Research [God06, vHG03, HBHN00, HW05, JCR05, KH08, McN03, Sim06, SSA05, SLD07, AD09, ACW06, AFKTUO09, AT01, BL07, BR00, BS00, BFS09, Bra03, BG03, Bru05b, But07, CVI02, CP03, CFBT^{+}06, CA07, CP07, CT01, Dal06, DLR03, DM07, DW04, Die00, DBFK07, DFG^{+}03, Don07a, Don07b, Dre02, EB01, ER08, EKHMW01, FC07, FM06, Fis07a, GG00, GH08, GMS09a, GHB09, Gug06, Gul05, GAP05, HJM07, Hei08a, HJW03, HAPN09, HS09, Hor08, HHH08, HTCR06, IM09, Ito04, JL09, Jap09, Kas01, Kat00, KK09, KL08a, Kne08, KL08b, Koc04, Kul09, KGR07, Kwa06, Lan07b, Lar01, Lep06, LvdBD^{+}07, Lep07b, LMJ^{+}09, LCS02, LRdCT00, MLG06, MU09, MMvPdlP08, Mic06, Moe07, MS01, MR06, OZ04, Ora04, Paa05, Pav00, Pay02, Per02, Por03, PR00, PR07]. research [Pou05, RL09, Ram08, Ree05, RT06, Rue07, SW00a, SGL03, SS09, Sch09b, Sen00, Sen01, Sen05, Sho03, SZ02, SW00b, Sko00a, Sko00b, Spi00, Suk08, TS01, TLL07, Tho00, Tsi01, VCDM^{+}08, VGP06, VD07, Wat05, Wat00, vdM03, Die00, OV05]. research-output [CA07]. research-policy [SGL03]. researcher [Gha09, KGR07]. researchers [Ito04, JGB09, Mor03, NS06]. resource [VGP06]. resources [CCSM05, FF03, Mah07, ME00, Sen01]. Response [Jac03, PEP04]. responses [LOAJ02]. responsibility [RD08]. responsible [Lew02]. responsiveness [SLD07, Wea01]. restructuring [MW01]. rests [Moe07]. results [DG01]. Rethinking [Haw08b, NM09]. retrospective [CA07, FC07]. return [Fon07]. Review [Hal06, Huf04, BFS09, BD09b, Don07a, Moe07, NS07, Val02]. reviews [Hol06]. Revising [DFG^{+}03, Leh04]. revisited [BH06, ER09, Gus03, Ram01]. revolution [Lan06c, Vog08a]. revolutions [Gal01]. rewards [OJ09]. Riddled [HSA01]. right [Duf02, Els09b, Els09a, Mat07, Moh02]. rights [Col00, Haw08b, Jen00, Win07]. rigorous [Joh09]. rise [Coo04, MF08]. Risk [BAV07, Gon04, Tod01, BA00, BA02, De 03, Dra09, Geo07, GD09, JN03, Sri06, Sti03, Tof05]. risks [Geo04, Rap08, Sti05]. risky [Kar02, OJ09, Pri03]. road [Duf02]. roads [PPMP09]. robust [Now03]. Role [Hem03, Ade02, BM06, Dyk04, GKZ09, Lya07, Pel06, PGS^{+}01, PR00, Ren01, Rol09, Smi03]. roles [BW01, GMG09, Gun06, Luu01, MW01, Ray03, vdM03]. rolling [CA07]. RTD [Kas01, Lar01, Sen01]. rule [FC07]. ruled [Fre01]. Russia [GKZ09, SW09]. Russian [Sim04, MR04].

S&T [Bar01, Gha09, God03, Luu07, ME00, SM04, Vel05]. safe [BA02]. saga [MvZ01]. same [TZ05]. sample [CA07]. satellite [FF03]. save [LD07]. Scale [Kat00]. Scale-independent [Kat00]. scandal [BM06]. Scandinavian [BS00, Sko00b]. scare [Hea01]. scenario [PM09]. schemes [Hei08a]. Schumpeter [Haw08a]. Science [AFS08, APS^{+}09, BK02, BS09, BM06, Buc03, CVI02, Chu02, DR01, FLS04, FSWR00, Gün05, Hal05, Jas04, Joh01, Kul09, LM01, May03, Mil05, Nie05, Nut00, Old06, PF06, Rav01, Sla05, Sta06, Ste02b, Teu03b, Wag02, Wea01, WS07, Ano01a, Ast08, Bal00, Bar07, Bir01, Bon07, BH06, Bor09a, BS05b, CL06, CS00a, CS00b, Coo04, CBD^{+}05, Coz07, CCSM05, Dal06, DL05, DW04, Don07a, Duf02, EB01, ER08, Etz06, FF08, FP09, Gib00, GG00, GD09, GJ06, Gum01, Gus03, Hei05b, HJ03, HJW03, HvLS09, HAPN09, Hol08, Iai04, Inz08, JGB09, KI09, Kar02, Kas00b, KK09, Keu04, KO04, Koc08, Kri01, Kwa06, Lan06b, Lan06c, Lar03, Lar08, Lee03, LDDM08, LD07, LCW05, Ley01, Lya07, MLG06, ML06a, Mab03, MJ01, Mah07]. science [Mar04, Mas01, MMvPdlP08, Mey08, MB08, Mor03, Ned09, Nis05, Old08, Par02, PTW08, PGS^{+}01, Pre06, Ram05, Ren01, Ric06, Rig07, Rip05, Rod07, Ros01, SQ09, Sch02, Sen03, Sha04, SR00, SW00b, Skl04, SLD07, Ste02a, Sti02, Swe07, TCW02, TS01, Wat00, Wet07, Woo00, dJM01, Pel06, TZ05]. science- [Lar08]. Science-policy [Hal05, GD09, JGB09, Kwa06, Pre06, Ram05]. science-society [DL05]. sciences [Car02, CL06, Len08, MR03, Orm00]. Scientific [Fon07, Gru06, Huf04, Maa03, BA02, Dem08, FS02, Gal01, GFH01, Gly01, KWC08, LLS06, MJRB05, Mej09, RG02, Sae02, Sch09b]. Scientist [MR06, Wat05]. Scientists [Fil07, JRKC08, LDDM08, Ric02, Bel06, Buc03, DR01, Fon07, Lau06, MR04, MR06, PR09, RD08, Rya08, Sim04, Swe07]. scientometric [CA07]. Scoreboard [SS08]. Scotland [Joh00a, Lya07]. screening [SS09]. scrutiny [Bed02, GJ06]. Scuppering [Ros02]. seafood [Ile04]. sealing [Rod06]. Seams [Zei04]. search [HvLS09]. Searching [OZ04]. section [BS09, RL09]. sector [AO09, BS00, Dal06, DBFK07, Fel07, Inz02, Inz08, KH08, LCS02, Mun08, PRB06, Par02, Ray03, RP03, Sen00, Sen01, TCW02, TA01, Tho00, WR01]. sectoral [Man04, Mes00]. sectors [Fri08, GGB02]. seed [BT02, GGB02]. seen [Moh02]. Segmenting [VGP06]. selection [CM02]. selective [Jac02]. self [McC02]. self-reflexive [McC02]. selves [Spe04]. Senior [Kim06]. Sense [Bar01]. sensitive [Gib00]. Separated [JWHJ06]. serendipity [KI09]. serve [ACW06]. service [DG01, Eva06, Inz02, Kin02]. services [DFG^{+}03, Gre05, HTCR06, KH08, Rub06].

Setting [Dal06, CFBT^{+}06, Dem08].

settings [RH07]. seven [CA07]. seven-year [CA07]. shake [Wei02]. shape [Wei02]. Shaping [Hou03b]. Shedding [Eva05]. Shift [Phi09]. Shifting [Von00]. Shifts [Fis07b]. Ships [Lip02, Tas02]. shops [FLS04]. should [Pav00]. sides [WT03]. significance [KL08b]. significant [Sri06]. silent [Whi03]. silvering [vdS08]. Simeonova [Mir04]. similar [Rod07]. simple [CA07]. since [LRS04]. Singapore [LWYP08, TA01]. Situating [HS07]. Six [Lar03]. size [Tho00]. Skin [Inz00]. Skin-deep [Inz00]. sky [Rus06]. Small [Cre07, Oss04, CS04, GGB02, MU09, Mey08, Tho00]. Smart [Bev07]. Smog [BK02]. smorgåsbord [Hol08]. snappy [Bev07]. Snapshot [Bra05]. Snapshots [Bru05b]. Social [Bir01, Bra07, CL06, GS03, Woo00, AFKTUO09, Bal00, BM00, CS00a, CS00b, Die00, Meh05, Nis05, Orm00, Ren01, RT06, Rya08, SW00b, Sti02, Tak01, TS01, Wat00, Wet07, vTN04]. socialist [ME00]. socially [Now03]. societal [FM06, Hei08b, KGR07, Sch05, WS07]. societal-ethical [Sch05]. societies [GS03]. society [Bor09b, DL05, FP09, FM00, Gib00, Gon04, GLR09, HvLS09, JRKC08, Keu04, KMM09, Maa03, MLG06, McC02, Rig07, Tra09, WT03]. socio [GD06, Kas01, KL01, Lov07, MD09, OJ09]. socio-cognitive [GD06]. socio-economic [Kas01, MD09, OJ09]. socio-technical [KL01, Lov07]. sociological [Ast08]. software [She05]. solutions [GMS09b]. Some [CP07, Old06, WK01, Wei03]. Something [Hol08]. soul [Inz00]. soul-deep [Inz00]. sound [RH07]. source [Cso06]. South [Kah06, KH08, Lor09, NM09, Pou05]. Southern [Ano00b]. Soviet [Sim04, MR04]. space [Gow03]. Spain [FP09]. Spanish [BR00, CCSM05, Hei05a, PEP04]. spanning [PR09]. special [BS09, Bor04a, CB04, CS00b, Els09b, Els09a, EZ08, Jac02, Lep07b, McN03, MY07b, RL09, Sen00, Suk08, VSZ07, Wet07]. specialisation [DBFK07]. specific [BS05a]. specificities [DFG^{+}03]. spin [KL08b, MWC08]. spin-off [MWC08]. spin-offs [KL08b]. sponsor [Hei08a]. SPP [Ano00c, Ano01c, Ano02, Ano03, Ano04, Ano05b, Ano06, Ano07, Ano08, Ano09a]. spread [ER08]. Staffan [AH03]. Stakeholder [CFBT^{+}06, CP07, FLS04, RP03]. stakeholders [MF08]. standard [Dem08]. standard-setting [Dem08]. Standardising [Dem08]. standards [Cha00, Dow05]. start [LWYP08, MdM09]. start-ups [LWYP08, MdM09]. State [Pel09, AP08, CBD^{+}05, Har02, Jas00, MF08, NGH08]. state-led [AP08]. States [RHL04, CBD^{+}05, Ile04, LRS04, Mic06, Von00]. stem [Hor08, Kul09, SQ09, Spe04]. stem-cell [Kul09]. step [May03]. stewardship [Ste03]. STI [Sen03]. still [ACW06]. Stimulate [Die00]. Stimulating [NHS08]. story [TZ05]. stranger [MPMA01]. Strategic [KY00, KGR07, SZ02, Lov07, Luu01]. strategies [Car02, CBD^{+}05, GG06, GGB02, MR06, vdM03]. strategy [Fuk08, OJ09, Par02, SZM08, She05]. strengthen [VC02]. strengthening [GG06]. Strengths [Seg00]. strong [RT06]. structural [Fis07a]. structure [Gus03, LT04, Sen00]. students [MR03]. studies [Hol08, Kin02, Lan06b, Ned09, SM04, Sha04, Wol07]. study [GMG09, HS09, Jac02, MMvPdlP08, MR04, Nis05, She05, TK08]. styles [Hal05, Len06]. subjects [Fis07a]. subordinate [Col00]. subsidiaries [CF08]. subversion [Lan06c]. success [BW01]. such [vBE04]. Suggestions [VC02, NS07]. superministry [Koc08]. supply [AD09, EHvM06, WT03]. support [FR09, Man04]. supported [Hei05a]. supranational [Sch09a].

Supremacy [Lee03]. survey [DG01, GAP05, Inz02]. surveys [Ric06, SH04]. survival [PF06]. Suspicious [CP07]. sustainability [Bri06, EZ06, vTN04]. sustainable [Ile04, Smi03]. Sweden [Jac02, LWYP08, MD06]. Swedish [GJ06, PGB09]. Swiss [Lep06]. Switzerland [GB08, SLD07]. SWOT [GG06]. Sydney [Hea01]. synergy [Jas04]. synopsis [EB01]. System [Bor04b, Ade02, AO09, AT01, BD09a, Bor04a, Bru09a, CT01, DW04, EK08, Els09b, Els09a, FMV09, For04, GG06, Hei08b, Hem03, IR08, Kah06, Keu03, KO04, Lew02, Lor09, MU09, Mes00, MR06, PGB09, Sch09a, Sen03, Sha06b, SdV^{+}09, Ste04, Xie07]. Systems [HW05, Bas00, BS00, Bor05, CS04, CP06, Dre02, Edq04, FR09, FM00, Fuk08, Gun06, HJM07, Har02, Har07, HTCR06, JL09, Lar04, Lov07, PR00, RL09, SW00a, Sen00, WMWW06]. Szilard [Lan06c].

Tackling [Leh03, Par02]. Tagus [Lar04]. Take [AG01b]. Taking [Duf02]. tale [Dre02]. talent [CL03]. Tanzania [MD09]. targeting [VGP06]. tasks [AFKTUO09]. teaching [AFKTUO09, NM09, SW00a]. tech [OJ09]. Technical [BW01, KL01, Lov07]. Technological [Sha03, Ade02, ARvT04, CM02, GD06, Har02, JL02, MW01, PPMP09, RH07, Ros00, Smi03, SW09]. technologies [Ben03, Fro03, JG09, KK01, NHS08, WR01]. Technology [Etz00, Gru03, HS07, Jas00, MJ01, Moh02, Pel06, PB09, Spi00, Von00, WK01, vTN04, AD09, BL07, BS09, BL03, BC06, CVI02, Cho01, Chu02, Col07, CBD^{+}05, Coz07, CT01, CCSM05, Dal06, DM06, Duf02, Eub07, FF08, Fuk08, Gru06, Gün05, Han06, Hei05a, Hem05, Hol08, HTCR06, Hud00, Joh09, KI09, Kas00b, Kim06, Kne08, KL08b, KO04, Koc08, Kri01, KY00, Lar03, LWYP08, LHP04, Mab03, Mah07, MPMA01, McC07, MB08, MY07a, MY07b, Ned09, Nis05, Old06, Old08, Par02, PTW08, PRJN02, Pov02, Ric06, Ros01, SK01, Sco07, Sen03, Sha04, Smi03, Sta06, Ste02a, Ste02b, TK08, TZ05, WT03, Win07, MJ01]. technology-based [CBD^{+}05]. telemedicine [FM00]. tells [ACM05]. Ten [NS07, MWC08]. Tensions [GD09, Gul05, Bra03]. term [Lep06, MB08]. terror [Rap03]. tertiary [Hei08b]. Testing [JCR05, Spi08]. texts [Ley01]. their [AC09, Axo00, Bal00, DW04, Lau06, Nis05, Wat05]. theoretical [DG01, Val02]. theory [Bor04b, BG03, DW04, Gus03, HJM07, MY07a, Nie05]. theory-based [HJM07]. there [MB08, Ste04]. Thin [Keu04]. thing [Moh02, Mun00]. Thinking [Blu08, MB08, Ley01]. third [Axo00, Blo07, KI09, Lin07, PGB09]. thoughtful [Ros07]. threat [Vog08b]. threats [Rap08]. three [Han06]. time [vHG03, KMM09, Wri03b]. timers [Per04]. tobacco [Lar08]. Token [KL08b]. tolerance [Rod06]. tomorrow [AG00c]. Too [BD09b, Buc03]. tool [PRB06, Rol09, SS09]. toolbox [Lan07b]. top [AC09, MD06]. Total [CA07]. tower [GP08]. towers [Ano01b]. track [Duf02]. Trade [Fre07, JN03, Cun08]. trading [Voß07]. Traditional [Des07, AO09]. trajectories [JL02]. trajectory [Mej09]. trans [ER09]. trans-Keynesian [ER09]. Transatlantic [Men07]. Transdisciplinarity [ML06a, Pre06]. transdisciplinary [Gug06, KK09, MLG06, Pre06]. transfer [AD09, BL07, Etz00, GMS09b, Jas00, KL08b, KY00, MJ01, Rod07, Spi00]. Transformation [Din07, AH03, Bra05, Jac02]. Transforming [Smi03]. transgenic [Cho01]. Transition [Hér05, Kne08, Bar07, Dyk00, Etz00, Gün05, Jas03b, Mes00, Ren05, Tra09]. Transnational [Dow05, Bar07]. transparent [Moe07]. Transport [Pou05, Mun08]. Trapped [Chi01]. Trends [LRS04, Ben03, CP03, Els09c, Kri01, Mic06]. trial [ML06b]. tried [Ros02]. Triple [EZ06, LE03, EKHMW01, Etz02, SZM08, VC02]. trust [Dra09, Hei08b, Ros01]. truthful [ACW06]. Trying [PS05]. turn [Sti02]. Turning [MMvPdlP08]. twins [EZ06]. Two [LLS06, CS04, Dre02, Duf02, HJM07, MR03]. two-track [Duf02].

UK [CP03, Coo07, EW04, Fin00, Fre07, HTCR06, JCR05, JWHJ06, Lit08, May03, McN03, MR06, Seg00, Spi00, Sti05]. Ultra [Sch02]. Ultra-left [Sch02]. uncertainty [BK02, Mit06]. under-evaluation [Geo07]. Underscoring [Joh09]. Understanding [Har07, Hor08, LNB06, McC03, CP07, DAZ09, LHJOS08, Rip05]. Undisciplined [Gug06]. UNESCO [Sta06, Ric06]. uneven [ER08]. unfolding [Koc04]. unicorns [SR00]. Unilever [Dou04]. unintended [PM09]. Union [TZ05, ACM05, Bor04b, DM07, FSWR00, GAP05, HS08, LCW05, LC05, Luu01, MW01, OV05, Pav00, Rap01, Ree05, Von00, Zit01, vTN04]. unique [Cso06]. United [CBD^{+}05, Dre02, Ile04, Mic06, Old06, Ore05, Sta06, Von00]. universe [Eva05]. Universities [AH03, Axo00, Jac02, MD09, AFKTUO09, AC09, CA07, GP08, GKZ09, GH08, GMS09a, GMS09b, GLR09, Gun06, Pay02, SZM08, Tra09, Val02]. University [CP03, HY09, SdV^{+}09, Thu07, BD09a, Bor09b, Bra07, Bru05a, Bru05b, But07, CA07, CT01, EKHMW01, Etz02, EZ08, ERB^{+}08, HAPN09, KMM09, LWYP08, LE03, MMvPdlP08, MS01, PGB09, RP03, Rod06, SW00a, Sen05, ZP08, MdM09, MWC08]. University-industry [CP03, SdV^{+}09, Thu07]. University-owned [HY09]. Upping [BC06]. ups [LWYP08, MdM09]. Uruguay [Bor09b]. USA [Bel02, Fis07a]. use [Atl08, Cro03, HH01, Kin02, SS09]. user [DLR03]. Users [SR00, Von00]. Using [Swe07, BD09a, CA07, GG06]. Utopian [Tay02].

Valley [Lar04]. Value [Xie07, KKBdCB07, MB08]. values [BS05b, KKBdCB07, Phi02]. variation [KKA09]. variegated [Eva05]. varieties [Wri03b]. Varsavsky [Sch02]. venture [RP03]. ventures [CVI02]. versus [Bon07, BH06, Buf01]. vibrant [Gol08]. Vietnamese [Tra09]. view [Fis07b]. viewed [Pre06]. views [LLS06]. virtual [HJW03]. Virtues [VCDM^{+}08]. visible [Etz06, Vis04]. vision [ER09]. vital [Tur07]. Vitalizing [PGB09]. voice [Zit01]. volunteer [EW04]. vs [Don07a].

Wage [Val09]. want [Gum01]. war [AG02, Mon04]. waste [Dur06, Dur07]. Water [Ric05, Hea01, JGB09]. watershed [Bor09a]. wavering [Jas00]. waves [Ros02]. way [Euc06, Lin07]. ways [Har02]. weaknesses [Seg00]. Weale [vE01]. weapons [Suk08]. web [Zei04, Joh00c]. Welfare [Jen00]. West [Gün04]. Western [Dyk04, LCS02]. whatever [Lip02]. Where [Bru09d, Por03]. Which [Rub06]. Whistleblowing [Mar03]. Whither [Ned09]. Who [AG00c, Fil07, JRKC08, McC03]. whose [vHG03]. wide [Lan03a]. Will [PM09]. Winding [PPMP09]. Winnebagos^{T M} [Vog08b]. winner [ER09]. Winter [Bru02]. wireless [Gre05]. wisdom [Fos04]. within [EK08, Old06, PEP04, TZ05]. without [BH08, Bru09d]. Women [Mas01]. Work [Vis04, Bel00a, Gol03, LDDM08, Swe07, Zeh05]. working [Det06]. workshop [JGB09]. World [JN03, Bar07, Dra09, Fre01, KK01, LD07, Mil05]. worlds [Gan04]. would [Wil04]. writer [Iai04].

Year [Ree05, CA07, GAP05, San07]. years [Bru02, MWC08]. yes [Det06]. York [Gol08]. young [Ito04]. yourself [Bru06a].

Zealand [DW04]. Ziman [Rod07].

References

Abergel:2005:LPM

[Abe05] Elisabeth A. Abergel. Learning from past mistakes. Science and Public Policy, 32(1):91–92, February 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andras:2009:WTU

[AC09] Peter Andras and Bruce G. Charlton. Why are top universities losing their lead? An economics modelling-based approach. Science and Public Policy, 36(4):317–330, May 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:2005:WER

[ACM05] Erik Arnold, John Clark, and Alessandro Muscio. What the evaluation record tells us about European Union framework programme performance. Science and Public Policy, 32(5): 385–397, October 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adam:2006:HSC

[ACW06] Matthias Adam, Martin Carrier, and Torsten Wilholt. How to serve the customer and still be truthful: methodological characteristics of applied research. Science and Public Policy, 33(6):435–444, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abramo:2009:APR

[AD09] Giovanni Abramo and Andrea D’Angelo. The alignment of public research supply and industry demand for effective technology transfer: the case of Italy. Science and Public Policy, 36(1):2–14, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adeoti:2002:BTC

[Ade02] John Olatunji Adeoti. Building technological capability in the less developed countries: The role of a national system of innovation. Science and Public Policy, 29(2):95–104, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adamsone-Fiskovica:2009:MBT

[AFKTUO09] Anda Adamsone-Fiskovica, Janis Kristapsons, Erika Tjunina, and Inga Ulnicane-Ozolina. Moving beyond teaching and research: economic and social tasks of universities in Latvia. Science and Public Policy, 36(2):133–137, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aksnes:2008:SPD

[AFS08] Dag W. Aksnes, Nicoline Frølich, and Stig Slipersæter. Science policy and the driving forces behind the internationalisation of science: The case of Norway. Science and Public Policy, 35(6):445–457, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2000:FHP

[AG00a] Janet Atkinson-Grosjean. Friends in high places. Science and Public Policy, 27(3):223–228, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2000:GN

[AG00b] Janet Atkinson-Grosjean. Genetic nationalism. Science and Public Policy, 27(2):150–151, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2000:WKW

[AG00c] Janet Atkinson-Grosjean. Who knows what tomorrow brings? Science and Public Policy, 27 (6):463–464, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2001:MM

[AG01a] Janet Atkinson-Grosjean. Modal models. Science and Public Policy, 28(4):325–327, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2001:TYP

[AG01b] Janet Atkinson-Grosjean. Take your partners and bow. Science and Public Policy, 28(1):82–84, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson-Grosjean:2002:DMW

[AG02] Janet Atkinson-Grosjean. Don’t mention the war! Science and Public Policy, 29(1):77–79, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andersson:2003:CSJ

[AH03] Thomas Andersson and Magnus Henrekson. A critique of Staffan Jacobsson’s paper “Universities and industrial transformation”. Science and Public Policy, 30(6):455–459, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Amanatidou:2002:FPI

[Ama02] Effie Amanatidou. Foreign policy and international R&D collaboration policy in Greece. Science and Public Policy, 29 (6):439–450, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andree:2006:AEI

[And06] Peter Andrée. An analysis of efforts to improve genetically modified food regulation in Canada. Science and Public Policy, 33(5):377–389, June 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anon:2000:PO

[Ano00a] Anon. Plethora of objectives. Science and Public Policy, 27 (2):148–149, April 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anon:2000:SHA

[Ano00b] Anon. Southern hemisphere approach. Science and Public Policy, 27(3):227–228, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2000:IS

[Ano00c] Anonymous. Index, SPP 2000. Science and Public Policy, 27 (6):467–468, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anon:2001:HSA

[Ano01a] Anon. Historiography of science according to Needham. Science and Public Policy, 28 (2):149–150, April 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anon:2001:IIA

[Ano01b] Anon. Inspiration and innovation among the ivory towers. Science and Public Policy, 28 (4):327–328, August 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2001:IS

[Ano01c] Anonymous. Index, SPP 2001.

Science and Public Policy, 28 (6):487–488, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2002:IS

[Ano02] Anonymous. Index, SPP 2002.

Science and Public Policy, 29 (6):483–484, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2003:IS

[Ano03] Anonymous. Index, SPP 2003.

Science and Public Policy, 30 (6):471–472, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2004:IS

[Ano04] Anonymous. Index, SPP 2004.

Science and Public Policy, 31 (6):499–500, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2005:EC

[Ano05a] Anonymous. Editors and contributors. Science and Public Policy, 32(5):334, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2005:IS

[Ano05b] Anonymous. Index, SPP 2005.

Science and Public Policy, 32 (6):499–500, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2006:IS

[Ano06] Anonymous. Index, SPP 2006. Science and Public Policy, 33 (10):774–776, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2007:IS

[Ano07] Anonymous. Index, SPP 2007. Science and Public Policy, 34 (10):758–760, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2008:IS

[Ano08] Anonymous. Index, SPP 2008. Science and Public Policy, 35 (10):785–788, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2009:IS

[Ano09a] Anonymous. Index, SPP 2009. Science and Public Policy, 36 (10):816–820, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2009:LR

[Ano09b] Anonymous. List of referees 2008. Science and Public Policy, 36(1):79–80, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adeoti:2009:TIS

[AO09] John Olatunji Adeoti and Olayiwola Olubamiwa. Towards an innovation system in the traditional sector: the case of the Nigerian cocoa industry.

Science and Public Policy, 36 (1):15–31, February 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Appelbaum:2008:CBB

[AP08] Richard P. Appelbaum and Rachel A. Parker. China’s bid to become a global nanotech leader: Advancing nanotechnology through state-led programs and international collaborations. Science and Public Policy, 35(5):319–334, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Antonopoulos:2009:MIP

[APS^{+}09] Constantinos N. Antonopoulos, Vagelis G. Papadakis, Chrysostomos D. Stylios, Maria P. Efstathiou, and Petros P. Groumpos. Mainstreaming innovation policy in less favoured regions: the case of Patras Science Park, Greece. Science and Public Policy, 36 (7):511–521, August 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arundel:2002:ABI

[Aru02]  Anthony Arundel.  Agro-

biotechnology, innovation and employment. Science and Public Policy, 29(4):297–306, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alfranca:2004:CDB

[ARvT04] Oscar Alfranca, Ruth Rama, and Nicholas von Tunzelmann. Combining different brands of in-house knowledge: technological capabilities in food, biotechnology, chemicals and drugs in agri-food multinationals. Science and Public Policy, 31(3):227–244, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Astbury:2008:DBS

[Ast08] Dr. Alan Astbury. Doing big science: a sociological perspective. Science and Public Policy, 35(5):374–375, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:2001:CBR

[AT01] Erik Arnold and Ben Thuriaux. Contribution of basic research to the Irish national innovation system. Science and Public Policy, 28(2):86– 98, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atlas:2008:TGH

[Atl08] Ronald M. Atlas. Toward global harmonization for control of dual-use biothreat agents. Science and Public Policy, 35(1):21–27, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Axon:2000:URT

[Axo00] Colin Axon. Universities reach their third age. Science and Public Policy, 27(2):151–152, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Axon:2002:GKY

[Axo02] Colin Axon. Get to know your history. Science and Public Policy, 29(6):481–482, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ayele:2007:LGG

[Aye07] Seife Ayele. The legitimation of GMO governance in Africa. Science and Public Policy, 34 (4):239–249, May 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barrett:2000:BFE

[BA00] Katherine Barrett and Elisabeth Abergel. Breeding familiarity: Environmental risk assessment for genetically engineered crops in Canada. Science and Public Policy, 27(1): 2–12, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barrett:2002:DSG

[BA02] Katherine Barrett and Elisabeth Abergel. Defining a safe genetically modified organism: Boundaries of scientific risk assessment. Science and Public Policy, 29(1):47–58, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baldwin:2000:ISS

[Bal00] Sally Baldwin. Interactive social science in practice: New approaches to the production of knowledge and their implications. Science and Public Policy, 27(3):183–194, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ballenger:2006:EP

[Bal06] Jesse F. Ballenger. Engaging and provocative. Science and Public Policy, 33(8):619–620, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barre:2001:SNS

[Bar01] Rémi Barré. Sense and nonsense of S&T productivity indicators. Science and Public Policy, 28(4):259–266, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barben:2007:CRS

[Bar07]  Daniel Barben.  Changing

regimes of science and politics: Comparative and transnational perspectives for a world in transition. Science and Public Policy, 34(1):55–69, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baskaran:2000:DNI

[Bas00] A. Baskaran. Duality in national innovation systems: the case of India. Science and Public Policy, 27(5):367–374, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ben-Ari:2007:RFK

[BAV07] Guy Ben-Ari and Nicholas S.

Vonortas. Risk financing for knowledge-based enterprises: Mechanisms and policy options. Science and Public Policy, 34(7):475–488, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Burgess:2006:UAC

[BC06] Jacquelin Burgess and Jason Chilvers. Upping the ante: a conceptual framework for designing and evaluating participatory technology assessments. Science and Public Policy, 33 (10):713–728, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bonaccorsi:2009:CEU

[BD09a] Andrea Bonaccorsi and Cinzia Daraio. Characterizing the European university system: a preliminary classification using census microdata. Science and Public Policy, 36(10):763–775, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boussaguet:2009:TBF

[BD09b] Laurie Boussaguet and Renaud Dehousse. Too big to fly? A review of the first EU citizens’ conferences. Science and Public Policy, 36(10):777–789, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Beder:2002:EUS

[Bed02] Sharon Beder. Experts under scrutiny. Science and Public Policy, 29(2):146–147, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Beech:2005:GLF

[Bee05] Colin Beech. Gambling with life: futures, insurance, and catastrophe. Science and Public Policy, 32(4):326–327, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:2000:DDW

[Bel00a] David Bell. Drugs don’t work. Science and Public Policy, 27 (3):224–225, June 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:2000:MM

[Bel00b] David Bell. Machinic mysticisms. Science and Public Policy, 27(2):147–148, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:2001:ED

[Bel01]  David Bell.  Entertaining

drama. Science and Public Policy, 28(5):395–396, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:2002:CU

[Bel02] David Bell. Cyberspace USA.

Science and Public Policy, 29 (1):79–80, February 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:2006:CSP

[Bel06] Sarah Bell. Concerned scientists, pragmatic politics and Australia’s green drought. Science and Public Policy, 33(8): 561–570, October 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ben:2003:PTI

[Ben03] Esther Ruiz Ben. Policy trends in information and communication technologies. Science and Public Policy, 30(5):386– 388, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Berggren:2005:EHM

[Ber05] Dr J. L. Berggren. Exploration and the history of mathematics. Science and Public Policy, 32(2):167–168, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Beveridge:2007:SS

[Bev07] Ross Beveridge. Smart and snappy. Science and Public Policy, 34(8):603–604, October 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boaz:2009:AIR

[BFS09] Annette Boaz, Siobhan Fitzpatrick, and Ben Shaw. Assessing the impact of research on policy: a literature review. Science and Public Policy, 36(4): 255–270, May 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:2003:PAT

[BG03] Dietmar Braun and David H. Guston. Principal-agent theory and research policy: an introduction. Science and Public Policy, 30(5):302–308, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boschert:2005:GAB

[BG05] Karin Boschert and Bernhard Gill. Germany’s agribiotechnology policy: precaution for choice and alternatives. Science and Public Policy, 32(4):285–292, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bora:2006:PSG

[BH06] Dr Alfons Bora and Dr Heiko Hausendorf. Participatory science governance revisited: Normative expectations versus empirical evidence. Science and Public Policy, 33(7):478– 488, August 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowman:2008:GNG

[BH08] Diana M. Bowman and Graeme A.

Hodge. ‘governing’ nanotechnology without government? Science and Public Policy, 35 (7):475–487, August 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Biegelbauer:2007:LAA

[Bie07] Dr Peter Biegelbauer. Learning from abroad: The Austrian competence centre programme Kplus. Science and Public Policy, 34(9):606–618, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Birrer:2001:SSP

[Bir01] Frans A. J. Birrer. Social science political counseling: The good, the bad and the enticed. Science and Public Policy, 28 (6):438–440, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bedsworth:2002:SUE

[BK02] Louise Wells Bedsworth and William E. Kastenberg. Science and uncertainty in environmental regulation: Insights from the evaluation of California’s Smog Check program.

Science and Public Policy, 29 (1):13–24, February 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2009:PRE

[BK09] Susana Borrás and Brian Kahin. Patent reform in Europe and the US. Science and Public Policy, 36(8):631–640, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Biggiero:2003:CET

[BL03] Lucio Biggiero and Domenico Laise. Choosing and evaluating technology policy: a multicriteria approach. Science and Public Policy, 30(1):13–23, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bach:2007:IHE

[BL07] Laurent Bach and Patrick Llerena. Indicators of highereducation institutes and publicresearch organizations technology transfer activities: Insights from France. Science and Public Policy, 34(10):709–721, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bleischwitz:2005:BAI

[Ble05] Dr Raimund Bleischwitz. Bottom-up approaches to international environmental policy. Science and Public Policy, 32(3):251–252, June 2005.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bloch:2007:ARD

[Blo07] Carter Bloch. Assessing recent developments in innovation measurement: The third edition of the Oslo Manual. Science and Public Policy, 34 (1):23–34, February 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blue:2008:TN

[Blu08] Gwendolyn Blue. Thinking with networks. Science and Public Policy, 35(6):459–460, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barrios-Medina:2000:DSP

[BM00] Ariel Barrios-Medina. Disease as a social process. Science and Public Policy, 27(5):388– 389, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bogner:2006:SCK

[BM06] Alexander Bogner and Wolfgang Menz. Science crime. the Korean cloning scandal and the role of ethics. Science and Public Policy, 33(8):601–612, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boggio:2006:WCG

[Bog06] Andrea Boggio. When the constitution goes to the laboratory. Science and Public Policy, 33(2):161–162, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bonaccorsi:2007:EPP

[Bon07] Andrea Bonaccorsi. Explaining poor performance of European science: Institutions versus policies. Science and Public Policy, 34(5):303–316, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2004:ISI

[Bor04a] Susana Borrás. Introduction to special issue on a European system of innovation. Science and Public Policy, 31(6):422– 424, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2004:SIT

[Bor04b] Susana Borrás. System of innovation theory and the European Union. Science and Public Policy, 31(6):425–433, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2005:CSA

[Bor05] Susana Borrás. Careful systems analysis of ERA. Science and Public Policy, 32(3):253– 254, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2009:ISP

[Bor09a] Susana Borrás. The interface of science and politics: a watershed. Science and Public Policy, 36(10):814–815, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bortagaray:2009:BUS

[Bor09b]  Isabel Bortagaray.  Bridg-

ing university and society in Uruguay: perceptions and expectations. Science and Public Policy, 36(2):115–119, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ballesteros:2000:PCO

[BR00] Juan Acosta Ballesteros and Aurelia Modrego Rico. Promotion of co-operative research: a Spanish experience. Science and Public Policy, 27(5):337– 346, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:2003:LTR

[Bra03] Dietmar Braun. Lasting tensions in research policy-making — a delegation problem. Science and Public Policy, 30(5): 309–321, October 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:2005:SPT

[Bra05] Dietmar Braun. Snapshot of process of transformation. Science and Public Policy, 32(1):

89–91, February 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braunerhjelm:2007:AES

[Bra07] Pontus Braunerhjelm. Academic entrepreneurship: Social norms, university culture and policies. Science and Public Policy, 34(9):619–631, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:2008:LPC

[Bra08a] Dietmar Braun. Lessons on the political coordination of knowledge and innovation policies. Science and Public Policy, 35 (4):289–298, May 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:2008:OPC

[Bra08b] Dietmar Braun. Organising the political coordination of knowledge and innovation policies. Science and Public Policy, 35(4):227–239, May 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brezis:2007:FRO

[Bre07] Elise S. Brezis. Focal randomisation: an optimal mechanism for the evaluation of R&D projects. Science and Public Policy, 34(10):691–698, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brister:2006:CAS

[Bri06] Dr Evelyn Brister. Communicating about sustainability. Science and Public Policy, 33 (5):391–392, June 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:2001:BGM

[Bro01] Nik Brown. Blood and guts in the museum. Science and Public Policy, 28(5):396–398, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2002:INW

[Bru02] David Bruggeman. Influence of Nelson and Winter 20 years on. Science and Public Policy, 29 (4):311–312, August 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2005:CAU

[Bru05a] David Bruggeman. Context of American university commerce. Science and Public Policy, 32(5):407–408, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2005:SAR

[Bru05b] David Bruggeman. Snapshots of an American research university. Science and Public Policy, 32(2):170–171, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2006:EDI

[Bru06a] David Bruggeman. Encouraging do-it-yourself innovation. Science and Public Policy, 33 (4):306–308, May 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2006:RCK

[Bru06b] David Bruggeman. Reflections on the construction of knowledge. Science and Public Policy, 33(8):623–624, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2007:PDP

[Bru07] David Bruggeman. Part Dilbert, part Dale Carnegie — one for fun. Science and Public Policy, 34(1):77–78, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2009:FNI

[Bru09a] David Bruggeman. A friendly neighborhood innovation system. Science and Public Policy, 36(1):74–76, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2009:LTM

[Bru09b] David Bruggeman. Laboring through the models. Science and Public Policy, 36(10):813– 814, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2009:JPP

[Bru09c] David Bruggeman. Not just a pretty patent. Science and Public Policy, 36(5):414–416, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2009:WIC

[Bru09d] David Bruggeman. Where is the innovator without the capital? Science and Public Policy, 36(2):165–166, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Benner:2000:ICS

[BS00] Mats Benner and Ulf Sandström. Inertia and change in Scandinavian public-sector research systems: the case of biotechnology. Science and Public Policy, 27(6):443–454, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bousios:2005:AAS

[BS05a] Alexandros Bousios and Jacqueline Senker. Assessing the achievement of specific policy objectives: biotechnology in Greece. Science and Public Policy, 32(1):79–87, February 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bozeman:2005:PVP

[BS05b] Barry Bozeman and Daniel Sarewitz. Public values and public failure in US science policy. Science and Public Policy, 32(2):119–136, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bellavista:2009:STP

[BS09] Joan Bellavista and Luis Sanz. Science and technology parks: habitats of innovation: introduction to special section. Science and Public Policy, 36(7): 499–510, August 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bijman:2002:PPI

[BT02] Jos Bijman and Joyce Tait. Public policies influencing innovation in the agrochemical, biotechnology and seed industries. Science and Public Policy, 29(4):245–251, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Buccino:2003:STI

[Buc03] Alphonse Buccino. Science is too important to be left to scientists. Science and Public Policy, 30(1):63–64, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bufton:2001:CHD

[Buf01]  Mark W. Bufton.  Coro-

nary heart disease versus BSE: Characterising official British expert advisory committees. Science and Public Policy, 28 (5):381–388, October 2001.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bunting:2000:CFP

[Bun00a] Andrea Bunting. The changing face of power. Science and Public Policy, 27(1):78–79, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bunting:2000:GI

[Bun00b] Andrea Bunting. Greening of industry. Science and Public Policy, 27(6):465–466, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Butler:2007:AUR

[But07] Dr Linda Butler. Assessing university research: a plea for a balanced approach. Science and Public Policy, 34(8):565– 574, October 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bozeman:2001:TRS

[BW01] Barry Bozeman and Dennis Wittmer. Technical roles and success of US federal laboratory–industry partnerships. Science and Public Policy, 28(3):169–178, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Charlton:2007:EUU

[CA07] Dr Bruce G. Charlton and Dr Peter Andras. Evaluating universities using simple scientometric research-output metrics: Total citation counts per university for a retrospective seven-year rolling sample. Science and Public Policy, 34(8): 555–563, October 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carr:2002:ISE

[Car02] Susan Carr. Innovation strategies in European agricultural life sciences: Introduction. Science and Public Policy, 29(4): 242–244, August 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caswill:2003:PAC

[Cas03]  Chris Caswill.  Principals,

agents and contracts. Science and Public Policy, 30(5):337– 346, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caracostas:2004:ISI

[CB04] Paraskevas Caracostas and Marie-Christine Brichard. Introduction to special issue on innovation policies for biotechnology in Europe. Science and Public Policy, 31(5):342–343, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cozzens:2005:DES

[CBD^{+}05] Susan E. Cozzens, Kamau Bobb, Kendall Deas, Sonia Gatchair, Albert George, and Gonzalo Ordonez. Distributional effects of science and technology-based economic development strategies at state level in the United States. Science and Public Policy, 32(1): 29–38, February 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cruz-Castro:2005:BST

[CCSM05] Laura Cruz-Castro and Luis Sanz-Menéndez. Bringing science and technology human resources back in: the Spanish Ramón y Cajal programme. Science and Public Policy, 32 (1):39–53, February 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Costa:2008:FOS

[CF08] Ionara Costa and Sergey Filippov. Foreign-owned subsidiaries: a neglected nexus between foreign direct investment, industrial and innovation policies. Science and Public Policy, 35(6):379–390, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caron-Flinterman:2006:SPH

[CFBT^{+}06] J. Francisca Caron-Flinterman, Jacqueline E. W. Broerse, Julia Teerling, Melissa L. Y. van Alst, Simon Klaasen, L. Edwin Swart, and Joske F. G. Bunders. Stakeholder participation in health research agenda setting: the case of asthma and COPD research in The Netherlands. Science and Public Policy, 33(4):291–304, May 2006.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chart:2000:ACS

[Cha00] David Chart. Abandon current standards and accept new ones. Science and Public Policy, 27(6):461–462, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chai:2009:CE

[Cha09] Choon-Lee Chai. Culture as enabler. Science and Public Policy, 36(1):76–77, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chilvers:2001:TIH

[Chi01] C. A. J. Chilvers. Trapped in the ironies of history. Science and Public Policy, 28(2):151– 152, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Choudhary:2001:DTB

[Cho01] Bhagirath Choudhary. Development of transgenic Bt cotton technology in India and China: a policy perspective. Science and Public Policy, 28(3):219– 229, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Christoforou:2003:PPD

[Chr03] Theofanis Christoforou. The precautionary principle and democratizing expertise: a European legal perspective. Science and Public Policy, 30(3): 205–211, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chung:2002:CTI

[Chu02] Sungchul Chung. Catching up through international linkages: Science, technology and the Korean experience. Science and Public Policy, 29(6):431– 437, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Castanos-Lomnitz:2003:EMT

[CL03] Heriberta Castaños-Lomnitz.

Emigration of Mexican talent: What price development? Science and Public Policy, 30(1): 41–45, February 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Castanos-Lomnitz:2006:SSS

[CL06] Heriberta Castaños-Lomnitz.

Social sciences and science policies in Mexico. Science and Public Policy, 33(2):115– 123, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carson:2002:RSC

[CM02] Lyn Carson and Brian Martin.

Random selection of citizens for technological decision making. Science and Public Policy, 29(2):105–113, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Collinson:2000:IRS

[Col00] Dr Simon Collinson. Individual rights subordinate to authority. Science and Public Policy, 27(3):225–227, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cole:2007:HMJ

[Col07] Simon A. Cole. How much justice can technology afford? The impact of DNA technology on equal criminal justice. Science and Public Policy, 34(2): 95–107, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooke:2004:BRR

[Coo04] Philip Cooke. Biosciences and the rise of regional science policy. Science and Public Policy, 31(3):185–197, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooke:2007:EAC

[Coo07] Philip Cooke. European asymmetries: a comparative analysis of German and UK biotechnology clusters. Science and Public Policy, 34(7):454–474, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cox:2004:GEH

[Cox04] Susan M. Cox. Genetics, ethics and hubris. Science and Public Policy, 31(6):497–498, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cozzens:2007:DJS

[Coz07] Susan E. Cozzens. Distributive justice in science and technology policy. Science and Public Policy, 34(2):85–94, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Calvert:2003:UIR

[CP03] J. Calvert and P. Patel.

University-industry research collaborations in the UK: Bibliometric trends. Science and Public Policy, 30(2):85–96, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Collins:2006:ISE

[CP06] Patrick Collins and Dimitrios Pontikakis. Innovation systems in the European periphery: The policy approaches of Ireland and Greece. Science and Public Policy, 33(10):757– 769, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cinderby:2007:SCS

[CP07] Steve Cinderby and Laura Potts. Suspicious cartographers: Some realities of research into stakeholder understanding of the causes and possible prevention of breast cancer. Science and Public Policy, 34(5):345–354, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cremers:2007:SEP

[Cre07] Katrin Cremers. Small entity patenting. Science and Public Policy, 34(2):151–152, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cross:2003:DGC

[Cro03] Alan Cross. Drawing up guidelines for the collection and use of expert advice: The experience of the European Commission. Science and Public Policy, 30(3):189–192, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crompton:2007:MKP

[Cro07] Dr Helen Crompton. Mode 2 knowledge production: Evidence from orphan drug networks. Science and Public Policy, 34(3):199–211, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caswill:2000:IIS

[CS00a] Chris Caswill and Elizabeth Shove. Introducing interactive social science. Science and Public Policy, 27(3):154– 157, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caswill:2000:PSI

[CS00b] Chris Caswill and Elizabeth Shove. Postscript to special issue on interactive social science. Science and Public Policy, 27(3):220–222, June 2000.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Calvert:2004:BIS

[CS04] Jane Calvert and Jacqueline Senker. Biotechnology innovation systems in two small countries: a comparison of Portugal and Ireland. Science and Public Policy, 31(5):359–370, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cotic-Svetina:2008:DCL

[CSJP08] Anja Cotic-Svetina, Marko Jaklic, and Igor Prodan. Does collective learning in clusters contribute to innovation? Science and Public Policy, 35(5): 335–345, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Csonka:2006:USI

[Cso06] László Csonka. A unique source of information on collaboration. Science and Public Policy, 33(9):697–700, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crow:2001:ARU

[CT01] Michael M. Crow and Christopher Tucker. The American research university system as America’s de facto technology policy. Science and Public Policy, 28(1):2–10, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cuntz:2008:IPR

[Cun08] Alexander Cuntz. Intellectual property regulation and international trade: National and global economic perspectives. Science and Public Policy, 35 (2):139–143, March 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caloghirou:2002:STP

[CVI02] Yannis Caloghirou, Nicholas S. Vonortas, and Stavros Ioannides. Science and technology policies towards research joint ventures. Science and Public Policy, 29(2):82–94, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dalrymple:2006:SAS

[Dal06] Dana G. Dalrymple. Setting the agenda for science and technology in the public sector: the case of international agricultural research. Science and Public Policy, 33(4):277– 290, May 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dawson:2002:CDM

[Daw02] Sandra Dawson. A century defined by its medicine. Science and Public Policy, 29(3):234– 240, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeMiranda:2009:CPE

[DAZ09] Paulo C. De Miranda, José Alberto S. Aranha, and Julia Zardo. Creativity: people, environment and culture, the key elements in its understanding and interpretation. Science and Public Policy, 36(7):523– 535, August 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dinges:2007:MSS

[DBFK07] Michael Dinges, Martin Berger, Rainer Frietsch, and Aris Kaloudis. Monitoring sector specialisation of public and private funded business research and development. Science and Public Policy, 34(6):431–443, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dowdeswell:2005:GGG

[DDS05] Elizabeth Dowdeswell, Dr Abdallah S. Daar, and Peter A. Singer. Getting governance into genomics. Science and Public Policy, 32(6):497–498, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeMarchi:2003:PPR

[De 03] Bruna De Marchi. Public participation and risk governance. Science and Public Policy, 30 (3):171–176, June 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Demortain:2008:STC

[Dem08] David Demortain. Standardising through concepts: The power of scientific experts in international standard-setting. Science and Public Policy, 35 (6):391–402, July 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Desai:2007:TKI

[Des07]  Pranav N. Desai.  Tradi-

tional knowledge and intellectual property protection: Past and future. Science and Public Policy, 34(3):185–197, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Deterding:2006:NPP

[Det06] Sebastian Deterding. A new policy paradigm, yes — but a working one? Science and Public Policy, 33(10):771–773, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Djellal:2003:RDR

[DFG^{+}03] Faridah Djellal, Dominique Francoz, Camal Gallouj, Faïz Gallouj, and Yves Jacquin. Revising the definition of research and development in the light of the specificities of services. Science and Public Policy, 30(6): 415–429, December 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Djellal:2001:PIO

[DG01] Faridah Djellal and Faïz Gallouj. Patterns of innovation organisation in service firms: postal survey results and theoretical models. Science and Public Policy, 28(1): 57–67, February 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dietz:2000:BSC

[Die00] James S. Dietz. Building a social capital model of research development: The case of the Experimental Program to Stimulate Competitive Research. Science and Public Policy, 27(2):137–145, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dinges:2007:TKB

[Din07] Michael Dinges. Transformation to a knowledge-based economy. Science and Public Policy, 34(7):529–531, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deJong:2001:PSO

[dJM01] Martin de Jong and Maarten Mentzel. Policy and science: Options for democratisation in European countries. Science and Public Policy, 28(6):403– 412, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schovsbo:2009:IAP

[dJS09] dr.jur. Jens Schovsbo. Increasing access to patented inventions by post-grant measures. Science and Public Policy, 36(8):609–618, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davenport:2005:AAM

[DL05] Sally Davenport and Shirley Leitch. Agoras, ancient and modern, and a framework for science-society debate. Science and Public Policy, 32(2):137– 153, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davenport:2003:URF

[DLR03] Sally Davenport, Shirley Leitch, and Arie Rip. The ‘user’ in research funding negotiation processes. Science and Public Policy, 30(4):239–250, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dethlefs:2006:JTP

[DM06] Noriko Dethlefs and Brian Martin. Japanese technology policy for aged care. Science and Public Policy, 33(1):47–57, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Delanghe:2007:EAI

[DM07] Henri Delanghe and Ugur Muldur. Ex-ante impact assessment of research programmes: The experience of the European Union’s 7th Framework Programme. Science and Public Policy, 34(3):169–183, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Donovan:2007:IFP

[Don07a] Claire Donovan. Introduction: Future pathways for science policy and research assessment: Metrics vs peer review, quality vs impact. Science and Public Policy, 34(8): 538–542, October 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Donovan:2007:QFR

[Don07b] Claire Donovan. The qualitative future of research evaluation. Science and Public Policy, 34(8):585–597, October 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Doubleday:2004:ING

[Dou04] Robert Doubleday. Institutionalising non-governmental organisation dialogue at unilever: framing the public as ‘consumercitizens’. Science and Public Policy, 31(2):117–126, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dowty:2005:TPS

[Dow05] Rachel Dowty. Transnational privacy standards, equity in practice. Science and Public Policy, 32(4):325–326, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Doern:2001:SSR

[DR01] G. Bruce Doern and Ted Reed.

Science and scientists in regulatory governance: a mezzolevel framework for analysis. Science and Public Policy, 28 (3):195–204, June 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Drahos:2009:CTR

[Dra09] Peter Drahos. Cooperation, trust and risk in the world’s major patent offices. Science and Public Policy, 36(8):641– 647, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dresner:2002:TTM

[Dre02] Simon Dresner. A tale of two ministers: Attempts at reform of research systems in The Netherlands and the United Kingdom. Science and Public Policy, 29(3):169–180, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:2002:TRF

[Duf02] Paul Dufour. Taking the (right?) fork in the road: Canada’s two-track approach to domestic and international science and technology. Science and Public Policy, 29(6): 419–429, December 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:2009:PM

[Duf09] Paul Dufour. Principles matter. Science and Public Policy, 36(9):737–738, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Durant:2006:MEP

[Dur06] Darrin Durant. Managing expertise: performers, principals, and problems in Canadian nuclear waste management. Science and Public Policy, 33(3): 191–204, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Durant:2007:BGA

[Dur07] Darrin Durant. Burying globally, acting locally: Control and co-option in nuclear waste management. Science and Public Policy, 34(7):515–528, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Devine:2004:GTA

[DW04] Sean Devine and Colin Webb.

A game theory analysis of how research organisations adapt their behaviour in the New Zealand competitively funded science system. Science and Public Policy, 31(5):407–414, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dyker:2000:EPT

[Dyk00] David A. Dyker. Economic performance in the transition economies: a comparative perspective. Science and Public Policy, 27(4):275–283, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dyker:2004:CPG

[Dyk04] David A. Dyker. Closing the productivity gap between Eastern and Western Europe: The role of foreign direct investment. Science and Public Policy, 31(4):279–287, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edler:2001:BNP

[EB01] Jakob Edler and Dr Patries Boekholt. Benchmarking national public policies to exploit international science and industrial research: a synopsis of current developments. Science and Public Policy, 28(4):313– 321, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edler:2004:BBD

[Edl04] Jakob Edler. Blurred borders — distinct concepts. Science and Public Policy, 31(3):245– 247, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edquist:2004:RSI

[Edq04] Charles Edquist. Reflections on the systems of innovation approach. Science and Public Policy, 31(6):485–489, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Engels:2006:WSM

[EHvM06] Dr Anita Engels, Dr Matthijs Hisschemöller, and Konrad von Moltke. When supply meets demand, yet no market emerges: The contribution of integrated environmental assessment to the rationalisation of EU environmental policymaking. Science and Public Policy, 33(7):519–528, August 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edler:2008:CWF

[EK08] Jakob Edler and Stefan Kuhlmann. Coordination within fragmentation: Governance in knowledge policy in the German federal system. Science and Public Policy, 35 (4):265–276, May 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Erno-Kjolhede:2001:MUR

[EKHMW01] Erik Ernø-Kjølhede, Kenneth Husted, Mette Mønsted, and Søren Barlebo Wenneberg. Managing university research in the Triple Helix. Science and Public Policy, 28(1):49–55, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ekins:2005:EAH

[Eki05] Paul Ekins. ‘evolutionary approach’ to human nature. Science and Public Policy, 32(3): 254–255, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elsmore:2009:GPPb

[Els09a] Matthew J. Elsmore. Getting patent policy right: a conclusion to a special issue on the European patent system. Science and Public Policy, 36(8): 649–652, October 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elsmore:2009:GPPa

[Els09b] Matthew J. Elsmore. Getting patent policy right: an introduction to a special issue on the European patent system. Science and Public Policy, 36(8):583–585, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elsmore:2009:PIT

[Els09c] Matthew J. Elsmore. Prevailing impact trends in patenting. Science and Public Policy, 36(8):587–594, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Engels:2008:USG

[ER08] Anita Engels and Tina Ruschenburg. The uneven spread of global science: Patterns of international collaboration in global environmental change research. Science and Public Policy, 35(5):347–360, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2009:TKV

[ER09] Henry Etzkowitz and Marina Ranga. A trans-Keynesian vision of innovation for the contemporary economic crisis: ‘picking winners’ revisited. Science and Public Policy, 36(10): 799–808, December 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2008:PEU

[ERB^{+}08] Henry Etzkowitz, Marina Ranga, Mats Benner, Lucia Guaranys, Anne Marie Maculan, and Robert Kneller. Pathways to the entrepreneurial university: towards a global convergence. Science and Public Policy, 35(9):681–695, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2000:TTE

[Etz00] Henry Etzkowitz. Technology transfer and the East European transition. Science and Public Policy, 27(4):230–234, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2002:III

[Etz02]  Henry Etzkowitz.  Incu-

bation of incubators: innovation as a Triple Helix of university–industry– government networks. Science and Public Policy, 29(2):115– 128, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2006:NVH

[Etz06] Henry Etzkowitz. The new visible hand: an assisted linear model of science and innovation policy. Science and Public Policy, 33(5):310–320, June 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eubanks:2007:PTE

[Eub07] Virginia Eubanks. Popular technology: Exploring inequality in the information economy. Science and Public Policy, 34 (2):127–138, March 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Euchler:2006:PHW

[Euc06] Gordon Euchler. A pre-hopeful way! Science and Public Policy, 33(3):231–232, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Evangelista:2005:SNL

[Eva05] Rinaldo Evangelista. Shedding new light on a variegated universe. Science and Public Policy, 32(2):162–163, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Evangelista:2006:IES

[Eva06] Rinaldo Evangelista. Innovation in the European service industries. Science and Public Policy, 33(9):653–668, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Enzing:2004:CBD

[EvdGK04] Christien Enzing, Annelieke van der Giessen, and Sander Kern. Commercialisation of biotechnology: Do dedicated public policies matter? Science and Public Policy, 31(5): 371–383, October 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ellis:2004:ECM

[EW04] Rebecca Ellis and Claire Waterton. Environmental citizenship in the making: the participation of volunteer naturalists in UK biological recording and biodiversity policy. Science and Public Policy, 31(2): 95–105, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2006:THT

[EZ06] Henry Etzkowitz and Chunyan Zhou. Triple Helix twins: innovation and sustainability. Science and Public Policy, 33(1): 77–83, February 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:2008:ISI

[EZ08] Henry Etzkowitz and Chunyan Zhou. Introduction to special issue building the entrepreneurial university: a global perspective. Science and Public Policy, 35(9):627–635, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Faulkner:2001:IME

[Fau01] Alex Faulkner. Innovation and the multiplicity of expertise. Science and Public Policy, 28 (1):80–82, February 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandez-Carro:2007:PAM

[FC07] Remo Fernández-Carro. A principal-agent model of public research with a retrospective payoff rule. Science and Public Policy, 34(5):317–328, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Feeney:2003:QD

[Fee03] Mary K. Feeney. Questions of democracy. Science and Public Policy, 30(6):467–468, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Feller:2007:MFE

[Fel07] Irwin Feller. Mapping the frontiers of evaluation of publicsector R&D programs. Science and Public Policy, 34(10):681– 690, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernos:2007:BP

[Fer07] Rodrigo Fernós. Broadband penetration. Science and Public Policy, 34(7):534–535, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Furtado:2003:AEI

[FF03] André Tosi Furtado and Edmilson Jesus Costa Filho. Assessing the economic impacts of the China–Brazil resources satellite program. Science and Public Policy, 30(1):25–39, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Felt:2008:BMC

[FF08] Ulrike Felt and Maximilian Fochler. The bottom-up meanings of the concept of public participation in science and technology. Science and Public Policy, 35(7):489–499, August 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Filas:2007:WIB

[Fil07] Michael Filas. Who invents biomedicine: Scientists, industry, or literature? Science and Public Policy, 34(1):74–75, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Finney:2000:ICB

[Fin00] Colin Finney. Implementing a citizen-based deliberative process on the Internet: The Buckinghamshire Health Authority Electronic Citizens’ Jury in the UK. Science and Public Policy, 27(1):45–64, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fisher:2007:GHS

[Fis07a] Jill A. Fisher. Governing human subjects research in the USA: Individualized ethics and structural inequalities. Science and Public Policy, 34(2):117– 126, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fisher:2007:SAD

[Fis07b] Saul Fisher. Shifts in academic disciplines from a constructivist view. Science and Public Policy, 34(6):447–449, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Flanagan:2000:CC

[Fla00] Kieron Flanagan. Comprehensive coverage. Science and Public Policy, 27(6):462–463, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fischer:2004:SSE

[FLS04] Corinna Fischer, Loet Leydesdorff, and Malte Schophaus. Science shops in Europe: the public as stakeholder. Science and Public Policy, 31(3):199– 211, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fujimoto:2000:IIG

[FM00] Dr Masayo Fujimoto and Dr Kumiko Miyazaki. Industrial innovation, government and society: telemedicine and healthcare systems in Japan. Science and Public Policy, 27 (5):347–366, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fisher:2006:CIU

[FM06] Erik Fisher and Roop L. Mahajan. Contradictory intent? US federal legislation on integrating societal concerns into nanotechnology research and development. Science and Public Policy, 33(1):5–16, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fagerberg:2009:ENN

[FMV09] Jan Fagerberg, David C. Mowery, and Bart Verspagen. The evolution of Norway’s national innovation system. Science and Public Policy, 36(6):431– 444, July 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fontes:2007:SMP

[Fon07] Margarida Fontes. Scientific mobility policies: How Portuguese scientists envisage the return home. Science and Public Policy, 34(4):284–298, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Foray:2004:PSD

[For04] Dominique Foray. The patent system and the dynamics of innovation in Europe. Science and Public Policy, 31(6):449– 456, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Foster:2004:AWE

[Fos04] Lynsey Foster. Applying wisdom to explosive issues. Science and Public Policy, 31(1): 78–80, February 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fourniau:2001:IAD

[Fou01] Jean-Michel Fourniau. Information, access to decisionmaking and public debate in France: The growing demand for deliberative democracy. Science and Public Policy, 28 (6):441–451, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ferretti:2009:WDC

[FP09] Maria Paola Ferretti and Vincenzo Pavone. What do civil society organisations expect from participation in science? Lessons from Germany and Spain on the issue of GMOs. Science and Public Policy, 36 (4):287–299, May 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandez-Ribas:2009:PSP

[FR09] Andrea Fernández-Ribas. Public support to private innovation in multi-level governance systems: an empirical investigation. Science and Public Policy, 36(6):457–467, July 2009.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freeman:2001:IRW

[Fre01] Christopher Freeman. If I ruled the world. Science and Public Policy, 28(6):477–479, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freitas:2007:NII

[Fre07] Isabel Maria Bodas Freitas.

New instruments in innovation policy: The case of the Department of Trade and Industry in the UK. Science and Public Policy, 34(9):644–656, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frietsch:2008:IIE

[Fri08] Rainer Frietsch. Inventions and innovations in the electrical and communications sectors: a historical perspective. Science and Public Policy, 35(8):618–620, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frost:2003:FGT

[Fro03] Robert Frost. Future of genetic technologies. Science and Public Policy, 30(5):385–386, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frewer:2002:PAS

[FS02] Lynn Frewer and Brian Salter. Public attitudes, scientific advice and the politics of regulatory policy: The case of BSE. Science and Public Policy, 29 (2):137–145, April 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Funtowicz:2000:SGE

[FSWR00] Silvio Funtowicz, Iain Shepherd, David Wilkinson, and Jerry Ravetz. Science and governance in the European Union: a contribution to the debate. Science and Public Policy, 27(5):327–336, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fukugawa:2008:ESL

[Fuk08] Nobuya Fukugawa. Evaluating the strategy of local public technology centers in regional innovation systems: Evidence from Japan. Science and Public Policy, 35(3):159– 170, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forbes:2004:WRW

[FW04] Naushad Forbes and David Wield. What is R&D? why does it matter? Science and Public Policy, 31(4):267–277, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Galles:2001:OSR

[Gal01] Carlos D. Galles. Origin of scientific revolutions. Science and Public Policy, 28(2):148– 149, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ganchoff:2004:IWC

[Gan04] Ma Chris Ganchoff. Imposing worlds of contemporary biotechnology. Science and Public Policy, 31(3):251–252, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Guy:2005:FPF

[GAP05] Ken Guy, Effie Amanatidou, and Foteini Psarra. Framework Programme 5 (FP5) impact assessment: a survey conducted as part of the five-year assessment of European Union research activities (1999–2003). Science and Public Policy, 32 (5):349–366, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Griessen:2008:PCK

[GB08] Thomas Griessen and Dietmar Braun. The political coordination of knowledge and innovation policies in Switzerland. Science and Public Policy, 35 (4):277–288, May 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geels:2006:LGD

[GD06] Frank Geels and J. Jasper Deuten. Local and global dynamics in technological development: a socio-cognitive perspective on knowledge flows and lessons from reinforced concrete. Science and Public Policy, 33(4):265–275, May 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goncalves:2009:PRC

[GD09] Maria Eduarda Gonçalves and Ana Delicado. The politics of risk in contemporary Portugal: Tensions in the consolidation of science-policy relations. Science and Public Policy, 36(3): 229–239, April 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gelfert:2001:BF

[Gel01] Axel Gelfert. Back to the future. Science and Public Policy, 28(3):230–232, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Georgieff:2004:MPM

[Geo04]  Peter Georgieff.  Mobile

phones/masts and health risks. Science and Public Policy, 31 (5):415–417, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Georghiou:2007:WLB

[Geo07] Luke Georghiou. What lies beneath: Avoiding the risk of under-evaluation. Science and Public Policy, 34(10):743–752, December 2007. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Georghiou:2001:BPS

[GFH01] Luke Georghiou, Dr Kieron Flanagan, and Peter Halfpenny. Benchmarking the provision of scientific equipment. Science and Public Policy, 28 (4):303–311, August 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Godin:2000:ICR

[GG00] Benoît Godin and Yves Gingras. Impact of collaborative research on academic science. Science and Public Policy, 27 (1):65–73, February 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ghazinoory:2006:DIG

[GG06] Sepehr Ghazinoory and Soroush Ghazinoori. Developing Iran’s government strategies for strengthening the national system of innovation using SWOT analysis. Science and Public Policy, 33(7):529–540, August 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gravalos:2002:PII

[GGB02] Esther Grávalos, Alejandro García, and Nick Barnes. Policy influences on innovation strategies of small and medium enterprises in the agrochemical, seed and plant biotechnology sectors. Science and Public Policy, 29(4):277–285, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goktepe-Hulten:2008:AIR

[GH08] Devrim Göktepe-Hultén. Academic inventors and research groups: entrepreneurial cultures at universities. Science and Public Policy, 35(9):657– 667, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ghazinoory:2009:DLI

[Gha09] Sepehr Ghazinoory. A day in the life of an Iranian S&T policy researcher. Science and Public Policy, 36(10):809–811, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grupp:2009:DRR

[GHB09] Hariolf Grupp, Sybille Hinze, and Barbara Breitschopf. Defining regional research priorities: a new approach. Science and Public Policy, 36(7): 549–559, August 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gibbons:2000:MSE

[Gib00] Michael Gibbons. Mode 2 society and the emergence of context-sensitive science. Science and Public Policy, 27(3): 159–163, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Granberg:2006:MRS

[GJ06] Anders Granberg and Staffan Jacobsson. Myths or reality — a scrutiny of dominant beliefs in the Swedish science policy debate. Science and Public Policy, 33(5):321–340, June 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gokhberg:2009:TNR

[GKZ09] Leonid Gokhberg, Tatiana Kuznetsova, and Dr Stanislav Zaichenko. Towards a new role of universities in Russia: prospects and limitations. Science and Public Policy, 36(2): 121–126, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gregersen:2009:LBD

[GLR09] Birgitte Gregersen, Lisbeth Tved Linde, and Jorgen Gulddahl Rasmussen. Linking between Danish universities and society. Science and Public Policy, 36(2):151–156, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Glynn:2001:PNS

[Gly01] Steven Glynn. Problematic nature of scientific advances. Science and Public Policy, 28(5): 398–399, October 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ghazinoory:2009:MNP

[GMG09] Sepehr Ghazinoory, Saber Mirzaei, and Soroush Ghazinoori. A model for national planning under new roles for government: Case study of the National Iranian Nanotechnology Initiative. Science and Public Policy, 36(3):241–249, April 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goransson:2009:INC

[GMS09a] Bo Göransson, Rasigan Maharajh, and Ulrich Schmoch. Introduction: New challenges for universities beyond education and research. Science and Public Policy, 36(2):83–84, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goransson:2009:NAU

[GMS09b] Bo Göransson, Rasigan Maharajh, and Ulrich Schmoch. New activities of universities in transfer and extension: multiple requirements and manifold solutions. Science and Public Policy, 36(2):157–164, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Godin:2003:MCC

[God03] Benoît Godin. A model to clarify concepts in evaluating S&T. Science and Public Policy, 30 (2):136–137, April 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Godin:2006:RDH

[God06] Benoît Godin. Research and development: how the ‘D’ got into R&D. Science and Public Policy, 33(1):59–76, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldhaber:2003:DPD

[Gol03] Michael H. Goldhaber. A device (paradigm) that does not work. Science and Public Policy, 30(1):65–67, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldhaber:2008:WMN

[Gol08] Michael H. Goldhaber. What makes new York vibrant? Science and Public Policy, 35(3): 223–224, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goncalves:2004:RSG

[Gon04] Maria Eduarda Gonçalves.

Risk society and the governance of innovation in Europe: opening the black box? Science and Public Policy, 31(6): 457–464, December 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gow:2003:MSM

[Gow03] Gordon A. Gow. Making space for McLuhan. Science and Public Policy, 30(2):143–144, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geoghegan:2008:ITF

[GP08] Will Geoghegan and Dimitrios Pontikakis. From ivory tower to factory floor? How universities are changing to meet the needs of industry. Science and Public Policy, 35(7):462–474, August 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gretchen:2005:DWE

[Gre05] Gano Gretchen. Developing wireless emergency services: a constructivist approach. Science and Public Policy, 32(6): 495–496, December 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grunwald:2003:TAG

[Gru03] Dr Armin Grunwald. Technology assessment at the German Bundestag: ‘expertising’ democracy for ‘democratising’ expertise. Science and Public Policy, 30(3):193–198, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grunwald:2006:SIC

[Gru06] Dr Armin Grunwald. Scientific independence as a constitutive part of parliamentary technology assessment. Science and Public Policy, 33(2):103– 113, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grundmann:2003:SCK

[GS03] Reiner Grundmann and Nico Stehr. Social control and knowledge in democratic societies. Science and Public Policy, 30(3):183–188, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Guggenheim:2006:URP

[Gug06] Michael Guggenheim. Undisciplined research: the proceduralisation of quality control in transdisciplinary projects. Science and Public Policy, 33(6): 411–421, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gulbrandsen:2004:MIN

[Gul04] Magnus Gulbrandsen. Multinational, international, national? Science and Public Policy, 31(6):493–496, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gulbrandsen:2005:TRC

[Gul05] Magnus Gulbrandsen. Tensions in the research council– research community relationship. Science and Public Policy, 32(3):199–209, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:2001:DSP

[Gum01] Philip Gummett. Delivering science that people want. Science and Public Policy, 28(6):

483–484, December 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gunther:2004:ICE

[Gün04] Dr Jutta Günther. Innovation cooperation: experiences from East and West Germany. Science and Public Policy, 31(2): 151–158, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gunther:2005:STT

[Gün05] Jutta Günther. Science and technology in transition. Science and Public Policy, 32(4): 328–329, August 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gunasekara:2006:GDR

[Gun06] Chrys Gunasekara. The generative and developmental roles of universities in regional innovation systems. Science and Public Policy, 33(2):137–150, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Guston:2003:PAT

[Gus03] David H. Guston. Principalagent theory and the structure of science policy, revisited: ‘science in policy’ and the US report on carcinogens. Science and Public Policy, 30(5):347– 357, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gusterson:2008:NFA

[Gus08] Hugh Gusterson. Nuclear futures: anticipatory knowledge, expert judgment, and the lack that cannot be filled. Science and Public Policy, 35(8):551– 560, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Halffman:2005:SPB

[Hal05] Willem Halffman. Sciencepolicy boundaries: national styles? Science and Public Policy, 32(6):457–467, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Halffman:2006:RE

[Hal06] Willem Halffman. Review essay. Science and Public Policy, 33(6):469–472, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hamill:2006:MJ

[Ham06] Lynne Hamill. Mobiles in Japan. Science and Public Policy, 33(7):543–544, August 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hansen:2006:OPP

[Han06] Janus Hansen. Operationalising the public in participatory technology assessment: a framework for comparison applied to three cases. Science and Public Policy, 33(8):571– 584, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Himanen:2009:IRF

[HAPN09] Laura Himanen, Otto Auranen, Hanna-Mari Puuska, and Mika Nieminen. Influence of research funding and science policy on university research performance: a comparison of five countries. Science and Public Policy, 36(6):419–430, July 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hart:2002:PTC

[Har02] David M. Hart. Private technological capabilities as products of national innovation systems: Four ways of looking at the state. Science and Public Policy, 29(3):181–188, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hart:2007:UIN

[Har07] David M. Hart. Understanding immigration in a national systems of innovation framework. Science and Public Policy, 34 (1):45–53, February 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawkins:2008:JSL

[Haw08a]  Richard Hawkins.  Joseph

Schumpeter’s life and legacy. Science and Public Policy, 35 (8):621–623, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawkins:2008:RIP

[Haw08b] Richard Hawkins. Rethinking intellectual property rights. Science and Public Policy, 35 (10):783–784, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hicks:2000:REP

[HBHN00] Diana Hicks, Sr Anthony Breitzman, Kimberly Hamilton, and Francis Narin. Research excellence and patented innovation. Science and Public Policy, 27(5):310–320, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Healy:2001:PPF

[Hea01] Stephen Healy. Privileging process over ‘fact’: The Sydney water scare as ‘organised irresponsibility’. Science and Public Policy, 28(2):123–129, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heijs:2005:IFS

[Hei05a] Joost Heijs. Identification of firms supported by technology policies: the case of Spanish low interest credits. Science and Public Policy, 32(3):219– 230, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heinze:2005:PAD

[Hei05b] Thomas Heinze. Pattern of accidental discoveries in science. Science and Public Policy, 32(5):410–413, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heinze:2008:HSG

[Hei08a]  Thomas Heinze.  How to

sponsor ground-breaking research: a comparison of funding schemes. Science and Public Policy, 35(5):302–318, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heitor:2008:SAT

[Hei08b] Manuel Heitor. A system approach to tertiary education institutions: towards knowledge networks and enhanced societal trust. Science and Public Policy, 35(8):607–617, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hemphill:2003:RCP

[Hem03] Thomas A. Hemphill. Role of competition policy in the US innovation system. Science and Public Policy, 30(4):285–294, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hemphill:2005:NTE

[Hem05] Thomas A. Hemphill. National technology entrepreneurship policy: foundation of a network economy. Science and Public Policy, 32(6):469–478, December 2005. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Heraud:2005:TKB

[Hér05] Jean-Alain Héraud. Transition to a knowledge-based economy. Science and Public Policy, 32 (2):164–167, April 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:2001:CUO

[HH01] J. Adam Holbrook and Lindsay P. Hughes. Comments on the use of the Organisation for Economic Cooperation and Development’s Oslo Manual in non-manufacturing based economies. Science and Public Policy, 28(2):139–144, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horta:2008:DCR

[HHH08] Hugo Horta, Jeroen Huisman, and Manuel Heitor. Does competitive research funding encourage diversity in higher education? Science and Public Policy, 35(3):146–158, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hicks:2009:III

[Hic09] Diana Hicks. Informing and improving innovation policy. Science and Public Policy, 36 (3):250–251, April 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellstrom:2003:BOS

[HJ03] Tomas Hellström and Merle Jacob. Boundary organisations in science: From discourse to construction. Science and Public Policy, 30(4):235–238, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hage:2007:TBI

[HJM07] Dr Jerald Hage, Dr Gretchen Jordan, and Dr Jonathan Mote. A theory-based innovation systems framework for evaluating diverse portfolios of research, part two: Macro indicators and policy interventions. Science and Public Policy, 34(10):731–741, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellstrom:2003:DPA

[HJW03] Tomas Hellström, Merle Jacob, and Søren Barlebo Wenneberg. The ‘discipline’ of post-academic science: reconstructing the paradigmatic foundations of a virtual research institute. Science and Public Policy, 30(4):251–260, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:2006:BR

[Hol06] Professor Adam Holbrook. Book reviews. Science and Public Policy, 33(6):474–475, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:2008:SER

[Hol08]  Adam Holbrook.  Some-

thing for everyone: a real smorgåsbord for science and technology studies. Science and Public Policy, 35(8):620, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horst:2008:LPD

[Hor08] Maja Horst. The laboratory of public debate: Understanding the acceptability of stem cell research. Science and Public Policy, 35(3):197–205, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

House:2003:MAE

[Hou03a] Dawn House. Multilevel analysis effective for mapping a complex entity. Science and Public Policy, 30(2):139–140, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

House:2003:SPP

[Hou03b] Dawn House. Shaping public policy and opinion. Science and Public Policy, 30(4):298– 299, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harsh:2007:TGP

[HS07] Matthew Harsh and James Smith. Technology, governance and place: Situating biotechnology in Kenya. Science and Public Policy, 34(4):251–260, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Huang:2008:PFG

[HS08] Can Huang and Luc Soete. Policy forum: The global challenges of the knowledge economy: China and the European Union. Science and Public Policy, 35(10):771–781, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holmes:2009:PMC

[HS09] John Holmes and Jennie Savgård. The planning, management and communication of research to inform environmental policy making and regulation: an empirical study of current practices in Europe. Science and Public Policy, 36(9): 709–721, November 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harty:2001:RI

[HSA01] Chris Harty, Elizabeth Shove, and Luis Araujo. Riddled with imprecision. Science and Public Policy, 28(3):233–235, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Howells:2006:ITR

[HTCR06] Jeremy Howells, Bruce Tether, Deborah Cox, and John Rigby.

Information technology research in the UK: perspectives on services research and development, and systems of innovation. Science and Public Policy, 33(1):17–31, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hudson:2000:PT

[Hud00] Robert G. Hudson. Philosophy of technology. Science and Public Policy, 27(5):387–388, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Huff:2004:RES

[Huf04] Toby E. Huff. Review essay:

Scientific enterprise in Islam. Science and Public Policy, 31 (1):69–75, February 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hessels:2009:SRC

[HvLS09] Laurens K. Hessels, Harro van Lente, and Ruud Smits. In search of relevance: The changing contract between science and society. Science and Public Policy, 36(5):387–401, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Howard:2001:PD

[HW01] Jeff Howard and Edward J.

Woodhouse. Provocative and disturbing. Science and Public Policy, 28(1):79–80, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:2005:ISR

[HW05] J. Adam Holbrook and David A. Wolfe. The innovation systems research network: a Canadian experiment in knowledge management. Science and Public Policy, 32(2):109–118, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haiyan:2009:UOE

[HY09] Wang Haiyan and Zhou Yuan. University-owned enterprises as entry point to the knowledge economy in China. Science and Public Policy, 36(2): 103–108, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Iain:2004:AMS

[Iai04] Taylor Iain. An absolute ‘must’ for any science writer. Science and Public Policy, 31(2):170– 171, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Iles:2004:MSS

[Ile04] Alastair Iles. Making seafood sustainable: merging consumption and citizenship in the United States. Science and Public Policy, 31(2):127– 138, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Islam:2009:NEN

[IM09] Nazrul Islam and Kumiko Miyazaki. NanoI: Exploring nanotechnology research conflation and nano-innovation dynamism in the case of Japan. Science and Public Policy, 36 (3):170–182, April 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Inzelt:2000:FDI

[Inz00] Dr Annamária Inzelt. Foreign direct investment in R&D: Skin-deep and soul-deep cooperation. Science and Public Policy, 27(4):241–251, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Inzelt:2002:ASI

[Inz02] Annamária Inzelt. Attempts to survey innovation in the Hungarian service sector. Science and Public Policy, 29(5):367– 383, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Inzelt:2008:PSI

[Inz08] Dr Annamária Inzelt. Private sector involvement in science and innovation policy-making in Hungary. Science and Public Policy, 35(2):81–94, March 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Iskandarani:2008:PMG

[IR08] Maria Iskandarani and Francisco J. B. Reifschneider. Performance measurement in a global program: motivation, new concepts and early lessons from a new system. Science and Public Policy, 35(10):745– 755, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ito:2004:JRG

[Ito04] Dr Yuko Ito. Japanese research grants for young researchers. Science and Public Policy, 31 (3):175–184, June 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jacobsson:2002:UIT

[Jac02] Staffan Jacobsson. Universities and industrial transformation: an interpretative and selective literature study with special emphasis on Sweden. Science and Public Policy, 29(5):345– 365, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jacobsson:2003:RCA

[Jac03] Staffan Jacobsson. Response to the critique by Andersson and Henrekson. Science and Public Policy, 30(6):459–461, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jakobs:2002:DPD

[Jak02] Kai Jakobs. Data protection dangers. Science and Public Policy, 29(2):149–150, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

James:2004:MM

[Jam04] Tansey James. Missing the mark. Science and Public Policy, 31(2):169–170, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jappe:2009:ICG

[Jap09] Arlette Jappe. International collaboration in global environmental research: a comparison of the International Geosphere Biosphere Program and the International Hydrological Program. Science and Public Policy, 36(3):183–197, April 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:2000:TTP

[Jas00] Andrzej H. Jasinski. Technology transfer in Poland: a poor state of affairs and a wavering policy. Science and Public Policy, 27(4):235–240, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasanoff:2003:NAE

[Jas03a] Sheila Jasanoff. (no?) accounting for expertise. Science and Public Policy, 30(3):157– 162, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:2003:IPI

[Jas03b] Andrzej H. Jasinski. Has innovation policy an influence on innovation? The case of a country in transition. Science and Public Policy, 30(6):431– 440, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasanoff:2004:SCN

[Jas04] Sheila Jasanoff. Science and citizenship: a new synergy. Science and Public Policy, 31 (2):90–94, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

James:2005:TBP

[JCR05] Andrew D. James, Deborah Cox, and John Rigby. Testing the boundaries of public private partnership: the privatisation of the UK defence evaluation and research agency. Science and Public Policy, 32(2): 155–161, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jennings:2000:WRJ

[Jen00] Richard Jennings. Welfare, rights and justice. Science and Public Policy, 27(4):307–308, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:2009:MIR

[JG09] Dr Mavis Jones and Dr Janice E. Graham. Multiple institutional rationalities in the regulation of health technologies: an ethnographic examination. Science and Public Policy, 36(6):445–455, July 2009.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jacobs:2009:SPI

[JGB09] Katharine Jacobs, Dr Gregg Garfin, and James Buizer. The science-policy interface: experience of a workshop for climate change researchers and water managers. Science and Public Policy, 36(10):791–798, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joly:2002:TTA

[JL02] Pierre-Benoit Joly and Stéphane Lemarié. The technological trajectories of the agrochemical industry: Change and continuity. Science and Public Policy, 29(4):259–266, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jablecka:2009:BHH

[JL09] Julita Jablecka and Benedetto Lepori. Between historical heritage and policy learning: the reform of public research funding systems in Poland, 1989– 2007. Science and Public Policy, 36(9):697–708, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joerges:2003:PRM

[JN03] Christian Joerges and Dr Jürgen Neyer. Politics, risk management, world trade organisation governance and the limits of legalisation. Science and Public Policy, 30(3):219–225, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:2000:CEC

[Joh00a] Sean F. Johnston. Creating energy in 19th century Scotland. Science and Public Policy, 27 (1):79–80, February 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:2000:MM

[Joh00b] Sean F. Johnston. Mixing metaphors. Science and Public Policy, 27(4):303–304, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:2000:WI

[Joh00c] Sean F. Johnston. Web of influence. Science and Public Policy, 27(5):389–390, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:2001:SPP

[Joh01] Sean Johnston. Science as a passionate pursuit. Science and Public Policy, 28(2):147– 148, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:2002:PMP

[Joh02] Sean Johnston. Practice makes perfect. Science and Public Policy, 29(6):478–479, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnson:2009:UNR

[Joh09] Rebecca Johnson. Underscoring the need for rigorous energy technology policy analysis. Science and Public Policy, 36(1):73–74, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jensen:2008:SWE

[JRKC08] Pablo Jensen, Jean-Baptiste Rouquier, Pablo Kreimer, and Yves Croissant. Scientists who engage with society perform better academically. Science and Public Policy, 35(7):527– 541, August 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:2006:SBC

[JWHJ06] Mavis Jones, John Walls, and Tom Horlick-Jones. Separated at birth? Consensus and contention in the UK agriculture and human biotechnology commissions. Science and Public Policy, 33(10):729–744, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kahn:2006:SAN

[Kah06] Michael Kahn. The South African national system of innovation: From constructed crisis to constructed advantage? Science and Public Policy, 33(2):125–136, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kariya:2002:APS

[Kar02] Dr Paul Kariya. Add people to science and it’ risky business. Science and Public Policy, 29 (4):307–308, August 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kash:2000:EB

[Kas00a] Don E. Kash. Empiricism at its best. Science and Public Policy, 27(1):76–77, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kass:2000:PDS

[Kas00b] Gary Kass. Public debate on science and technology: issues for legislators. Science and Public Policy, 27(5):321–326, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kastrinos:2001:CSE

[Kas01] Nikos Kastrinos. Contribution of socio-economic research to the benchmarking of RTD policies in Europe. Science and Public Policy, 28(4):238–246, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kass:2003:GE

[Kas03] Gary Kass. Green experimentation. Science and Public Policy, 30(3):231–232, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Katz:2000:SII

[Kat00]  J. Sylvan Katz.  Scale-

independent indicators and research evaluation. Science and Public Policy, 27(1):23–36, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keenan:2001:CC

[Kee01] Michael Keenan. Concise and cogent. Science and Public Policy, 28(3):235–236, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keelan:2003:CCG

[Kee03] Jennifer Keelan. Cyborg or cyber-goddess? Science and Public Policy, 30(1):64–65, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keune:2003:NSC

[Keu03] Hans Keune. Need for system change. Science and Public Policy, 30(3):228–229, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keune:2004:TLB

[Keu04] Hans Keune. Thin lines between science and society. Science and Public Policy, 31(1): 87–88, February 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kurek:2007:RES

[KGR07] Kasia Kurek, Peter A. T. M. Geurts, and Hans E. Roosendaal. The research entrepreneur: Strategic positioning of the researcher in his societal environment. Science and Public Policy, 34(7):501–513, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kahn:2008:RDS

[KH08] Michael Kahn and Lidwine Hounwanou. Research and development in the services sector of an emerging economy: The case of South Africa. Science and Public Policy, 35(7): 515–526, August 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kakko:2009:HCC

[KI09] Ilkka Kakko and Dr Sam Inkinen. Homo creativus: creativity and serendipity management in third generation science and technology parks. Science and Public Policy, 36 (7):537–548, August 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kidd:2004:GPP

[Kid04] Kerry Kidd. Genetics, policy and public education. Science and Public Policy, 31(1):76–77, February 2004. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2000:GIR

[Kim00] Sun G. Kim. Is government investment in R&D and market environment needed for indigenous private R&D in less developed countries?: Evidence from Korea. Science and Public Policy, 27(1):13–22, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kimpeler:2006:SCT

[Kim06] Dr Simone Kimpeler. Senior citizens and technology. Science and Public Policy, 33(1): 86–88, February 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kinder:2002:GPB

[Kin02] Tony Kinder. Good practice in best practice: The use of best practice case studies in service innovation by local public administrations. Science and Public Policy, 29(3):221– 233, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kingsley:2006:FFP

[Kin06] Danny Kingsley. Fairer and faster publishing. Science and Public Policy, 33(8):620–621, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kash:2001:PWC

[KK01] Don E. Kash and William Kingston. Patents in a world of complex technologies. Science and Public Policy, 28(1):11–22, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kaufmann:2009:DPD

[KK09] Alexander Kaufmann and Andrea Kasztler. Differences in publication and dissemination practices between disciplinary and transdisciplinary science and the consequences for research evaluation. Science and Public Policy, 36(3):215–227, April 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kleinman:2009:LVG

[KKA09] Daniel Lee Kleinman, Abby J. Kinchy, and Robyn Autry. Local variation or global convergence in agricultural biotechnology policy? A comparative analysis. Science and Public Policy, 36(5):361–371, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kupper:2007:VLE

[KKBdCB07] Frank Kupper, Linda Krijgsman, Henriette Bout, and Tjard de Cock Buning. The value lab: Exploring moral frameworks in the deliberation of values in the animal biotechnology debate. Science and Public Policy, 34(9):657–670, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kinder:2001:BAC

[KL01] Tony Kinder and Nick Lancaster. Building absorptive capacity in a learning region: a socio-technical model. Science and Public Policy, 28(1):23–40, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Klerkx:2008:DAR

[KL08a] Laurens Klerkx and Cees Leeuwis. Delegation of authority in research funding to networks: Experiences with a multiple goal boundary organization. Science and Public Policy, 35(3):183–196, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Knie:2008:TES

[KL08b] Andreas Knie and Martin Lengwiler. Token endeavors: The significance of academic spin-offs in technology transfer and research policy in Germany. Science and Public Policy, 35(3):171–182, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krucken:2009:LCS

[KMM09] Georg Krücken, Dr Frank Meier, and Andre Müller. Linkages to the civil society as ‘leisure time activities’? Experiences at a German university.

Science and Public Policy, 36 (2):139–144, March 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kneucker:2008:TPA

[Kne08] Raoul Kneucker. Transition policies: Austrian research and technology 1945–2005. Science and Public Policy, 35(3):221– 222, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kobayashi:2004:DAK

[KO04] Shinichi Kobayashi and Yoshiko Okubo. Demand articulation, a key factor in the reconfiguration of the present Japanese science and technology system. Science and Public Policy, 31 (1):55–67, February 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koch:2004:PUC

[Koc04] Tom Koch. Puzzle unfolding: complexity of interdisciplinary research. Science and Public Policy, 31(5):417–419, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koch:2008:SAI

[Koc08] Christian Koch. The superministry approach: Integrated governance of science, technology and innovation with contracted autonomy. Science and Public Policy, 35(4):253–264, May 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krishna:2001:CPC

[Kri01] V. V. Krishna. Changing policy cultures, phases and trends in science and technology in India. Science and Public Policy, 28(3):179–194, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kulawik:2009:SPP

[Kul09] Teresa Kulawik. Science policy and public accountability in Poland: The case of embryonic stem-cell research. Science and Public Policy, 36(6):469– 482, July 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kwa:2006:PIR

[Kwa06] Chunglin Kwa. The programming of interdisciplinary research through informal science-policy interactions. Science and Public Policy, 33 (6):457–467, July 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kale:2008:DKT

[KWC08] Dr Dinar Kale, David Wield, and Joanna Chataway. Diffusion of knowledge through migration of scientific labour in India. Science and Public Policy, 35(6):417–430, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kubielas:2000:SAT

[KY00] Stanislaw Kubielas and Igor Yegorov. Strategic alliances and technology transfer in Central and Eastern Europe. Science and Public Policy, 27 (4):265–273, August 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2002:AE

[Lan02] Cooper H. Langford. Actors’ evaluations. Science and Public Policy, 29(1):76–77, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2003:DWA

[Lan03a] Cooper H. Langford. Deserves a wide audience. Science and Public Policy, 30(3):227–228, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2003:EPI

[Lan03b] Cooper H. Langford. Economics paradigm inadequate. Science and Public Policy, 30 (6):468–469, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2004:FER

[Lan04] Cooper H. Langford. Fascinating but not an easy read. Science and Public Policy, 31(2): 166–167, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2005:CAB

[Lan05] Cooper H. Langford. Conversations across boundaries. Science and Public Policy, 32(5): 414–415, October 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2006:LR

[Lan06a] Cooper H. Langford. A lively read. Science and Public Policy, 33(10):770–771, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2006:PSS

[Lan06b] Cooper H. Langford. A popularization of science studies. Science and Public Policy, 33 (6):473–474, July 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lanouette:2006:SPL

[Lan06c] William Lanouette. The science and politics of Leo Szilard, 1898–1964: Evolution, revolution, or subversion? Science and Public Policy, 33(8): 613–617, October 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2007:CYG

[Lan07a] Cooper H. Langford. Can you grow it? Science and Public Policy, 34(10):753–754, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2007:TCR

[Lan07b] Cooper H. Langford. A toolbox for complexity research. Science and Public Policy, 34(4): 299–300, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laredo:2001:BRP

[Lar01] Philippe Larédo. Benchmarking of RTD policies in Europe: ‘research collectives’ as an entry point for renewed comparative analyses. Science and Public Policy, 28(4):285–294, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laredo:2003:SMC

[Lar03] Philippe Larédo. Six major challenges facing public intervention in higher education, science, technology and innovation. Science and Public Policy, 30(1):4–12, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laranja:2004:ISR

[Lar04]  Manuel Laranja.  Innova-

tion systems as regional policy frameworks: The case of Lisbon and Tagus Valley. Science and Public Policy, 31(4):313– 327, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Larsen:2008:PIS

[Lar08] Lars Thorup Larsen. The political impact of science: is tobacco control science- or policy-driven? Science and Public Policy, 35(10):757–769, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laudel:2006:AGF

[Lau06] Grit Laudel. The art of getting funded: How scientists adapt to their funding conditions. Science and Public Policy, 33(7):489–504, August 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2005:PEE

[LC05] Les Levidow and Susan Carr.

Precautionary expertise for European Union agbiotech regulation. Science and Public Policy, 32(4):258–259, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2002:APS

[LCS02] Les Levidow, Susan Carr, and Villy Søgaard. Agricultural public-sector research establishments in Western Europe: Research priorities in conflict. Science and Public Policy, 29 (4):287–295, August 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2005:EUR

[LCW05] Les Levidow, Susan Carr, and David Wield. European Union regulation of agribiotechnology: precautionary links between science, expertise and policy. Science and Public Policy, 32(4):261–276, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lemos:2007:EFC

[LD07] Maria Carmen Lemos and Lisa Dilling. Equity in forecasting climate: Can science save the world’s poor? Science and Public Policy, 34(2):109– 116, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lehoux:2008:SPM

[LDDM08] Pascale Lehoux, Geneviève Daudelin, Jean-Louis Denis, and Fiona Miller. Scientists and policy-makers at work: Listening to epistemic conversations in a genetics science network. Science and Public Policy, 35(3):207–220, April 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leydesdorff:2003:CPC

[LE03] Loet Leydesdorff and Henry Etzkowitz. Can ‘the public’ be considered as a fourth helix in university–industry– government relations? Report on the Fourth Triple Helix Conference, 2002. Science and Public Policy, 30(1):55–61, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lee:2003:SS

[Lee03] Chai Choon Lee. Supremacy of science. Science and Public Policy, 30(6):469–470, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lehr:2003:TEC

[Leh03] Jane L. Lehr. Tackling environmental concerns productively. Science and Public Policy, 30(2):137–139, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lehr:2004:BPR

[Leh04] Jane L. Lehr. Beyond the public: revising democracy as we know it. Science and Public Policy, 31(2):167–169, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leinhos:2005:UNB

[Lei05] Mary Leinhos. The US national bioethics advisory commission as a boundary organization. Science and Public Policy, 32(6):423–433, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lengwiler:2006:BCH

[Len06] Dr Martin Lengwiler. Between charisma and heuristics: four styles of interdisciplinarity. Science and Public Policy, 33(6):423–434, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lentzos:2008:CML

[Len08] Filippa Lentzos. Countering misuse of life sciences through regulatory multiplicity. Science and Public Policy, 35(1): 55–64, February 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lepori:2006:PRF

[Lep06] Benedetto Lepori. Public research funding and research policy: a long-term analysis for the Swiss case. Science and Public Policy, 33(3):205– 216, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lepori:2007:CMQ

[Lep07a] Benedetto Lepori. From conceptual modeling to quantitative measure. Science and Public Policy, 34(3):226–227, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lepori:2007:PII

[Lep07b] Benedetto Lepori. The power of indicators: Introduction to special issue on public project funding of research. Science and Public Policy, 34(6):370– 371, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lewis:2002:CRS

[Lew02] Jeffrey W. Lewis. Creating a responsible system of oversight. Science and Public Policy, 29(4):310–311, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leydesdorff:2001:RTR

[Ley01] Loet Leydesdorff. Re-thinking relations between texts and contexts in science. Science and Public Policy, 28(6):484– 486, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Liberatore:2003:DEE

[LF03] Angela Liberatore and Silvio Funtowicz. ‘democratising’ expertise, ‘expertising’ democracy: What does this mean, and why bother? Science and Public Policy, 30(3):146– 150, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lopez:2008:LPN

[LHJOS08] Dr Ana Prades López, Dr Tom Horlick-Jones, Dr Christian Oltra, and Dr Rosario Solá. Lay perceptions of nuclear fusion: multiple modes of understanding. Science and Public Policy, 35(2):95–105, March 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Liston-Heyes:2004:ICP

[LHP04] Catherine Liston-Heyes and Alan Pilkington. Inventive concentration in the production of green technology: a comparative analysis of fuel cell patents. Science and Public Policy, 31(1):15–25, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lin:2007:KHT

[Lin07] Yu-Wei Lin. Katherine Hayles’ third way towards posthumanity. Science and Public Policy, 34(3):227–228, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lipsett:2002:SCW

[Lip02] Morley Lipsett. Ships, chips and whatever is next. Science and Public Policy, 29(5):397– 398, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lipsey:2004:CO

[Lip04] Richard G. Lipsey. Conflicting opinions. Science and Public Policy, 31(4):331–340, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Littlewood:2008:MBD

[Lit08] Jez Littlewood. Managing biological disarmament: The UK experience. Science and Public Policy, 35(1):13–20, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lenhard:2006:EKM

[LLS06] Johannes Lenhard, Holger Lücking, and Holger Schwechheimer. Expert knowledge, Mode-2 and scientific disciplines: Two contrasting views. Science and Public Policy, 33 (5):341–350, June 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lissoni:2009:APP

[LLST09] Francesco Lissoni, Peter Lotz, Jens Schovsbo, and Adele Treccani. Academic patenting and the professor’s privilege: evidence on Denmark from the KEINS database. Science and Public Policy, 36(8):595–607, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2001:SGE

[LM01] Les Levidow and Claire Marris. Science and governance in Europe: Lessons from the case of agricultural biotechnology. Science and Public Policy, 28(5):345–360, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lepori:2009:COP

[LMJ^{+}09] Benedetto Lepori, Jaan Masso, Julita Jablecka, Karel Sima, and Kadri Ukrainski. Comparing the organization of public research funding in Central and Eastern European countries. Science and Public Policy, 36(9):667–681, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Luukkonen:2006:UDN

[LNB06] Terttu Luukkonen, Maria Nedeva, and Rémi Barré. Understanding the dynamics of networks of excellence. Science and Public Policy, 33(4):239– 252, May 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2002:IRE

[LOAJ02] Les Levidow, Sue Oreszczyn, Gérald Assouline, and Pierre-Benoit Joly. Industry responses to the European controversy over agricultural biotechnology. Science and Public Policy, 29(4):267–275, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lopez:2004:BRH

[Lóp04] José López. Bioethics and regulation of human genetic engineering. Science and Public Policy, 31(2):165–166, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lorentzen:2009:LFB

[Lor09] Jo Lorentzen. Learning by firms: the black box of South Africa’s innovation system. Science and Public Policy, 36 (1):33–45, February 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lovell:2007:GIS

[Lov07] Dr Heather Lovell. The governance of innovation in sociotechnical systems: The difficulties of strategic niche management in practice. Science and Public Policy, 34(1):35–44, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Longo:2000:RER

[LRdCT00] Waldimir Pirróe Longo, Ivan Rocha, and Maria Hortência da Costa Telles. ‘reengineering’ engineering research and education in Brazil: Cooperative networks and coalitions. Science and Public Policy, 27 (1):37–44, February 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lacasa:2004:TGB

[LRS04] Iciar Dominguez Lacasa, Thomas Reiss, and Jacqueline Senker. Trends and gaps in biotechnology policies in European Member States since 1994. Science and Public Policy, 31(5):385– 395, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lyall:2004:FML

[LT04] Catherine Lyall and Joyce Tait. Foresight in a multi-level governance structure: policy integration and communication. Science and Public Policy, 31(1):27–37, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lundin:2007:IIM

[Lun07] Nannan Lundin. Inspiring insights into MNEs activities in Asia. Science and Public Policy, 34(5):364–367, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Luukkonen:2001:ONS

[Luu01] Terttu Luukkonen. Old and new strategic roles for the European Union framework programme. Science and Public Policy, 28(3):205–218, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Luukkonen:2007:NES

[Luu07] Terttu Luukkonen. A new European S&T governance. Science and Public Policy, 34(8): 599–601, October 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lepori:2007:CEN

[LvdBD^{+}07] Benedetto Lepori, Dr Peter van den Besselaar, Michael Dinges, Bianca Potì, Emanuela Reale, Stig Slipersæter, Jean Thèves, and Barend van der Meulen. Comparing the evolution of national research policies: What patterns of change? Science and Public Policy, 34 (6):372–388, July 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leong:2008:EFN

[LWYP08] Bernard Leong, Alison Kim Shan Wee, and Ho Yuen-Ping. Is an enterprise framework necessary for an entrepreneurial university? A comparison of technology start-ups in Singapore and Sweden. Science and Public Policy, 35(9):647–656, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lyall:2007:CBR

[Lya07] Dr Catherine Lyall. Changing boundaries: The role of policy networks in the multi-level governance of science and innovation in Scotland. Science and Public Policy, 34(1):3–14, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maarten:2003:SEK

[Maa03] Mentzel Maarten. Scientific ethos in a knowledge society. Science and Public Policy, 30 (5):388, October 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mabawonku:2003:CFD

[Mab03] Azeez Olugbenga Mabawonku.

Cultural framework for the development of science and technology in Africa. Science and Public Policy, 30(2):117–125, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahroum:2007:AHR

[Mah07] Dr Sami Mahroum. Assessing human resources for science and technology: The 3Ds framework. Science and Public Policy, 34(7):489–499, August 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahajan:2008:DEM

[Mah08] Manjari Mahajan. Designing epidemics: models, policymaking, and global foreknowledge in India’s AIDS epidemic. Science and Public Policy, 35(8):585–596, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mangematin:2004:SHP

[Man04] Vincent Mangematin. From sectoral to horizontal public policies: The evolution of support for biotechnology in Europe, 1994–2001. Science and Public Policy, 31(5):397–406, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:2003:WA

[Mar03] Brian Martin. Whistleblowing, anybody? Science and Public Policy, 30(3):229–231, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martello:2004:GCS

[Mar04] Marybeth Long Martello. Global change science and the Arctic citizen. Science and Public Policy, 31(2):107–115, April 2004. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Mason:2001:WSO

[Mas01] Dr Joan Mason. Women in science: Outsiders and insiders. Science and Public Policy, 28 (2):145–146, April 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Matos:2007:GEA

[Mat07] Stelvia Matos. Global environmental assessments: How to get it right. Science and Public Policy, 34(6):445–447, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayer:2003:SSP

[May03] Sue Mayer. Science out of step with the public: The need for public accountability of science in the UK. Science and Public Policy, 30(3):177–181, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maze:2002:HWG

[Maz02] Jack Maze. Here we go again, creationists and evolutionists in battle. Science and Public Policy, 29(5):399–400, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Miller:2008:TLT

[MB08] Clark A. Miller and Ira Bennett. Thinking longer term about technology: is there value in science fiction-inspired approaches to constructing futures? Science and Public Policy, 35(8):597–606, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McCormick:2002:TSR

[McC02] Brad McCormick. Towards a self-reflexive society. Science and Public Policy, 29(3):237– 238, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McCormick:2003:UWW

[McC03] Brad McCormick. Understanding who we conceive ourselves to be. Science and Public Policy, 30(5):382–383, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McCormick:2007:ALQ

[McC07]  Brad McCormick.  Asking

large questions about technology. Science and Public Policy, 34(1):75–76, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McNay:2003:AAA

[McN03] Ian McNay. Assessing the assessment: an analysis of the UK Research Assessment Exercise, 2001, and its outcomes, with special reference to research in education. Science and Public Policy, 30(1):47–54, February 2003. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Melin:2006:TEP

[MD06] Göran Melin and Rickard Danell. The top eight percent: Development of approved and rejected applicants for a prestigious grant in Sweden. Science and Public Policy, 33(10):702– 712, December 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mwamila:2009:USE

[MD09] Burton L. M. Mwamila and Bitrina D. Diyamett. Universities and socio-economic development in Tanzania: public perceptions and realities on the ground. Science and Public Policy, 36(2):85–90, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maculan:2009:USU

[MdM09] Anne-Marie Maculan and José Manoel Carvalho de Mello. University start-ups for breaking lock-ins of the Brazilian economy. Science and Public Policy, 36(2):109–114, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Muller:2000:SHR

[ME00] Associate Professor Karel Müller and Henry Etzkowitz. S&T human resources: The comparative advantage of the post-socialist countries. Science and Public Policy, 27(4): 285–291, August 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mehta:2005:BBS

[Meh05] Michael D. Mehta. Balancing the books on social cohesion. Science and Public Policy, 32 (1):88, February 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mejlgaard:2009:TSC

[Mej09] Niels Mejlgaard. The trajectory of scientific citizenship in Denmark: Changing balances between public competence and public participation. Science and Public Policy, 36 (6):483–496, July 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Menrad:2007:TCA

[Men07] Klaus Menrad. Transatlantic conflict over agro-food biotechnology. Science and Public Policy, 34(1):70–73, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meske:2000:CIS

[Mes00] Dr Werner Meske. Changes in the innovation system in economies in transition: Basic patterns, sectoral and national particularities. Science and Public Policy, 27(4):253– 264, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer:2008:DSS

[Mey08] Morgan B. Meyer. The dynamics of science in a small country: The case of Luxembourg. Science and Public Policy, 35 (5):361–371, June 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McLeish:2008:BSR

[MF08] Caitríona McLeish and Daniel Feakes. Biosecurity and stakeholders: The rise of networks and non-state actors. Science and Public Policy, 35(1):5–12, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Musyck:2005:TKB

[MH05] Bernard Musyck and Athanasios Hadjimanolis. Towards a knowledge-based economy: does the Cyprus R&D capability meet the challenge? Science and Public Policy, 32(1): 65–77, February 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Michelson:2006:ARD

[Mic06]  Evan S. Michelson.  Ap-

proaches to research and development performance assessment in the United States: an analysis of recent evaluation trends. Science and Public Policy, 33(8):546–560, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Middendorf:2002:OEM

[Mid02] Gerad Middendorf. One Europe or many? Science and Public Policy, 29(6):479–481, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Miller:2005:SDG

[Mil05]  Clark A. Miller.  Science

and democracy in a globalizing world: challenges for American foreign policy. Science and Public Policy, 32(3):174– 186, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mirskaya:2004:RCK

[Mir04] Elena Z. Mirskaya. Reply to the comments of Kostadinka Simeonova. Science and Public Policy, 31(6):490–491, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mittra:2006:GEP

[Mit06] James Mittra. ‘genetic exceptionalism’ and precautionary politics: Regulating for uncertainty in Britain’s genetics and insurance policy process. Science and Public Policy, 33(8): 585–600, October 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Macdonald:2001:TTI

[MJ01] Stuart Macdonald and Richard Joseph. Technology transfer or incubation? Technology business incubators and science and technology parks in the Philippines. Science and Public Policy, 28(5):330–344, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marris:2005:HFG

[MJRB05] Claire Marris, Pierre-Benoit Joly, Stéphanie Ronda, and Christophe Bonneuil. How the French GM controversy led to the reciprocal emancipation of scientific expertise and policy making. Science and Public Policy, 32(4):301–308, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maasen:2006:TNM

[ML06a] Professor Dr Sabine Maasen and Olivier Lieven. Transdisciplinarity: a new mode of governing science? Science and Public Policy, 33(6):399– 410, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Millo:2006:REG

[ML06b] Dr Yuval Millo and Javier Lezaun. Regulatory experiments: genetically modified crops and financial derivatives on trial. Science and Public Policy, 33(3):179–190, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maasen:2006:PTR

[MLG06] Dr Sabine Maasen, Dr Martin Lengwiler, and Dr Michael Guggenheim. Practices of transdisciplinary research: close(r) encounters of science and society. Science and Public Policy, 33(6):394–398, July 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mathieu:2008:TSB

[MMvPdlP08] Azèle Mathieu, Martin Meyer, and Bruno van Pottelsberghe de la Potterie. Turning science into business: a case study of a major European research university. Science and Public Policy, 35(9):669–679, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moed:2007:FRE

[Moe07] Dr Henk F. Moed. The future of research evaluation rests with an intelligent combination of advanced metrics and transparent peer review. Science and Public Policy, 34(8):575– 583, October 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mohr:2002:BSD

[Moh02] Alison Mohr. Of being seen to do the right thing: Provisional findings from the first Australian Consensus Conference on Gene Technology in the Food Chain. Science and Public Policy, 29(1):2–12, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Montgomery:2004:BMI

[Mon04] Scott L. Montgomery. Battles for the mind, images at war. Science and Public Policy, 31 (3):248–250, June 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morris:2003:ARA

[Mor03] Norma Morris. Academic researchers as ‘agents’ of science policy. Science and Public Policy, 30(5):359–370, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marschan-Piekkari:2001:BSF

[MPMA01] Dr Rebecca Marschan-Piekkari, Stuart Macdonald, and Dimitris Assimakopoulos. In bed with a stranger: finding partners for collaboration in the European information technology programme. Science and Public Policy, 28(1):68–78, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mangematin:2003:TFP

[MR03] V. Mangematin and S. Robin.

The two faces of PhD students: Management of early careers of French PhDs in life sciences. Science and Public Policy, 30 (6):405–414, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mirskaya:2004:RAS

[MR04] Elena Z. Mirskaya and Yakov M. Rabkin. Russian academic scientists in the first post-Soviet decade: Empirical study. Science and Public Policy, 31(1): 2–14, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morris:2006:SCS

[MR06] Norma Morris and Arie Rip. Scientists’ coping strategies in an evolving research system: the case of life scientists in the UK. Science and Public Policy, 33(4):253–263, May 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Menrad:2008:AEE

[MR08] Klaus Menrad and Daniela Reitmeier. Assessing economic effects: Co-existence of genetically modified maize in agriculture in France and Germany. Science and Public Policy, 35 (2):107–119, March 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morgan:2001:UUR

[MS01] Robert P. Morgan and Donald E. Strickland. US university research contributions to industry: Findings and conjectures. Science and Public Policy, 28(2):113–121, April 2001.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Masso:2009:CPP

[MU09] Jaan Masso and Kadri Ukrainski. Competition for public project funding in a small research system: the case of Estonia. Science and Public Policy, 36(9):683–695, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Muchie:2000:PNC

[Muc00] Mammo Muchie. Are the UN programmes of national cleaner production centres necessary for introducing cleaner production practices in industry? Science and Public Policy, 27(5):375–385, October 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mundell:2000:LBD

[Mun00] Ian Mundell. A little biodiversity is a dangerous thing. Science and Public Policy, 27 (1):77–78, February 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munn:2008:PPF

[Mun08] Dr R. E. (Ted) Munn. Past, present and future of the transport sector. Science and Public Policy, 35(9):699–700, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Millstone:2001:PEA

[MvZ01] Dr Erik Millstone and Dr Patrick van Zwanenberg. Politics of expert advice: Lessons from the early history of the BSE saga. Science and Public Policy, 28 (2):99–112, April 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meske:2001:EUE

[MW01] Werner Meske and K. Matthias Weber. European Union enlargement: economic restructuring in candidate countries and the roles of technological change and education. Science and Public Policy, 28(3):154– 168, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mustar:2008:USF

[MWC08] Philippe Mustar, Mike Wright, and Bart Clarysse. University spin-off firms: Lessons from ten years of experience in Europe. Science and Public Policy, 35(2):67–80, March 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mohan:2007:GDT

[MY07a] Dr Giles Mohan and Dr Helen Yanacopulos. Governing and democratising technology for development: Bridging theory and practice. Science and Public Policy, 34(4):233–238, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mohan:2007:ISI

[MY07b] Dr Giles Mohan and Dr Helen Yanacopulos. Introduction to special issue on governing technology for development. Science and Public Policy, 34(4): 231–232, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Navratil:2007:EMC

[Nav07]  David Navratil.  Explor-

ing masculinity construction through reproduction. Science and Public Policy, 34(2):152– 153, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nedeva:2009:WST

[Ned09] Dr Maria Nedeva. Whither science and technology studies? Science and Public Policy, 36 (8):655, October 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Newell:2003:CM

[New03] Dianne Newell. Collective memory. Science and Public Policy, 30(4):299–300, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelson:2008:IAS

[NGH08] Nicole Nelson, Anna Geltzer, and Stephen Hilgartner. Introduction: the anticipatory state: making policy-relevant knowledge about the future. Science and Public Policy, 35 (8):546–550, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Negro:2008:SRE

[NHS08] Simona O. Negro, Marko P. Hekkert, and Ruud E. H. M. Smits. Stimulating renewable energy technologies by innovation policy. Science and Public Policy, 35(6):403–416, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Niederberger:2005:SCC

[Nie05] Dr Anne Arquit Niederberger. Science for climate change policy-making: applying theory to practice to enhance effectiveness. Science and Public Policy, 32(1):2–16, February 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nishizawa:2005:CDS

[Nis05] Mariko Nishizawa. Citizen deliberations on science and technology and their social environments: case study on the Japanese consensus conference on GM crops. Science and Public Policy, 32(6):479–489, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ndabeni:2009:RLB

[NM09] Lindile L. Ndabeni and Rasigan Maharajh. Rethinking the linkages between teaching and extension in South Africa. Science and Public Policy, 36(2): 127–132, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nowotny:2003:DES

[Now03] Dr Helga Nowotny. Democratising expertise and socially robust knowledge. Science and Public Policy, 30(3):151–156, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nerdrum:2006:MFR

[NS06] Dr Lars Nerdrum and Bo Sarpebakken. Mobility of foreign researchers in Norway. Science and Public Policy, 33(3):217– 229, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nightingale:2007:PRR

[NS07] Dr Paul Nightingale and Dr Alister Scott. Peer review and the relevance gap: Ten suggestions for policy-makers. Science and Public Policy, 34(8): 543–553, October 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nuttall:2000:SPP

[Nut00] William J. Nuttall. Science — pure politics? Science and Public Policy, 27(4):304–306, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ones:2009:HES

[OJ09] Isarelis Pérez Ones and Dr Jorge Núñez Jover. Higher education and socio-economic development in Cuba: high rewards of a risky high-tech strategy. Science and Public Policy, 36(2):97– 101, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oldham:2006:STA

[Old06] Geoffrey Oldham. Science and technology advice within the United Nations: Some lessons from past experience. Science and Public Policy, 33(9):647– 651, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oldham:2008:MGA

[Old08]  Geoffrey Oldham.  Mrs

Gandhi’s approach to science and technology. Science and Public Policy, 35(10):782–783, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

OReilly:2009:DEG

[O’R09] Jessica O’Reilly. Defining and exploring governance. Science and Public Policy, 36(5):413– 414, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oravec:2004:EEC

[Ora04] Jo Ann Oravec. Examining the examined career: Diana Forsythe as ethnographer and participant in computing research. Science and Public Policy, 31(2):159–163, April 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oreszczyn:2005:GCU

[Ore05] Sue Oreszczyn. GM crops in the United Kingdom: precaution as process. Science and Public Policy, 32(4):317–324, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Orme:2000:ISS

[Orm00] Joan Orme. Interactive social sciences: Patronage or partnership? Science and Public Policy, 27(3):211–219, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ossenkopf:2004:SFI

[Oss04] Birgit Ossenkopf. Small firms as innovators. Science and Public Policy, 31(6):496–497, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Otsuki:2006:BE

[Ots06] Grant Otsuki. Biography of the electron. Science and Public Policy, 33(9):695–696, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ormala:2005:EEU

[OV05] Erkki Ormala and Nicholas S.

Vonortas. Evaluating the European Union’s Research Framework Programmes: 1999– 2003. Science and Public Policy, 32(5):399–406, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Okubo:2004:SRI

[OZ04] Yoshiko Okubo and Michel Zitt. Searching for research integration across Europe: a closer look at international and inter-regional collaboration in France. Science and Public Policy, 31(3):213–226, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Paasi:2005:CBP

[Paa05] Dr Marianne Paasi. Collective benchmarking of policies: an instrument for policy learning in adaptive research and innovation policy. Science and Public Policy, 32(1):17–27, February 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Papaioannou:2006:ECA

[Pap06] Theo Papaioannou. European competitive advantage: Quality products and innovation. Science and Public Policy, 33 (7):541–543, August 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Parthasarathi:2002:TBD

[Par02] Ashok Parthasarathi. Tackling the brain drain from India’s information and communication technology sector: The need for a new industrial, and science and technology strategy. Science and Public Policy, 29 (2):129–136, April 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pavitt:2000:WEU

[Pav00] Professor Keith Pavitt. Why European Union funding of academic research should be increased: a radical proposal. Science and Public Policy, 27 (6):455–460, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Payne:2002:DUC

[Pay02] A. Abigail Payne. Do US Congressional earmarks increase research output at universities? Science and Public Policy, 29(5):314–330, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Philbrick:2009:NCT

[PB09] Mark Philbrick and Javiera Barandiaran. The national citizens’ technology forum: lessons for the future. Science and Public Policy, 36(5):335– 347, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pelkonen:2006:PII

[Pel06] Antti Pelkonen. The problem of integrated innovation policy: Analyzing the governing role of the Science and Technology Policy Council of Finland. Science and Public Policy, 33(9): 669–680, November 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pelaez:2009:SER

[Pel09] Victor Pelaez. State of exception in the regulation of genetically modified organisms in Brazil. Science and Public Policy, 36(1):61–71, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Perez-Escolano:2004:IRS

[PEP04] Isabel Pérez-Escolano and Gonzalo París. Industry response to the Spanish governmental Plan for the Promotion of R&D within the Pharmaceutical Industry (1986–1996). Science and Public Policy, 31 (4):301–312, August 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pereira:2002:IDR

[Per02] Tiago Santos Pereira. International dimension of research in Portugal: The European Research Area and beyond. Science and Public Policy, 29(6): 451–461, December 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pereira:2004:PTA

[Per04] Tiago Santos Pereira. Parttimers in academia. Science and Public Policy, 31(3):250– 251, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Prainsack:2006:SSB

[PF06] Barbara Prainsack and Ofer Firestine. ‘science for survival’: biotechnology regulation in Israel. Science and Public Policy, 33(1):33–46, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Palsson:2009:VSU

[PGB09] Carl Magnus Pâlsson, Bo Göransson, and Claes Brundenius. Vitalizing the Swedish university system: implementation of the ‘third mission’. Science and Public Policy, 36(2):145–150, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Polt:2001:BIS

[PGS^{+}01] Wolfgang Polt, Helmut Gassler, Andreas Schibany, Christian Rammer, and Doris Schartinger. Benchmarking industry-science relations: the role of framework conditions. Science and Public Policy, 28 (4):247–258, August 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Phillips:2002:CVR

[Phi02] Peter W. B. Phillips. Contextual values and regulatory decisions. Science and Public Policy, 29(5):398–399, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Phillips:2009:SNE

[Phi09] Tanya Phillips. Shift to a new economy? Science and Public Policy, 36(8):653–655, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Propp:2009:WGE

[PM09] Dr Tilo Propp and Dr Ellen H. M. Moors. Will genomics erode public health and prevention? A scenario of unintended consequences in The Netherlands. Science and Public Policy, 36(3):199–213, April 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porter:2003:IEW

[Por03] Alan L. Porter. Iraqi engineering: Where has all the research gone? Science and Public Policy, 30(2):97–105, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pouris:2005:TRS

[Pou05] Anastassios Pouris. Transport research in South Africa: a quantitative assessment. Science and Public Policy, 32(3): 211–218, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Poveda:2002:GIN

[Pov02] José Ignacio Pradas Poveda.

The geography of innovation: a new model of technology and innovation policies in a decentralised country. Science and Public Policy, 29(5):385–396, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Padilla-Perez:2009:EGF

[PPMP09] Ramón Padilla-Pérez and Jorge Mario Martínez-Piva. Export growth, foreign direct investment and technological capability building under the maquila model: Winding roads, few intersections. Science and Public Policy, 36(4): 301–315, May 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Poti:2000:CDI

[PR00] Bianca Potì and Emanuela Reale. Convergence and differentiation in institutional change among European public research systems: the decreasing role of public research institutes. Science and Public Policy, 27(6):421–431, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Poti:2007:CAM

[PR07] Bianca Potì and Emanuela Reale. Changing allocation models for public research funding: an empirical exploration based on project funding data. Science and Public Policy, 34(6):417–430, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pulzl:2009:IDB

[PR09] Helga Pülzl and Ewald Rametsteiner. Indicator development as ‘boundary spanning’ between scientists and policymakers. Science and Public Policy, 36(10):743–752, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Papaioannou:2006:BPM

[PRB06] Theodoros Papaioannou, Howard Rush, and John Bessant. Benchmarking as a policymaking tool: From the private to the public sector. Science and Public Policy, 33(2):91– 102, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pregernig:2006:TVA

[Pre06] Michael Pregernig. Transdisciplinarity viewed from afar: science-policy assessments as forums for the creation of transdisciplinary knowledge. Science and Public Policy, 33 (6):445–455, July 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Priest:2003:RC

[Pri03] Susanna Hornig Priest. A risky career. Science and Public Policy, 30(6):463–464, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porter:2002:MNE

[PRJN02] Alan L. Porter, J. David Roessner, Xiao-Yin Jin, and Nils C. Newman. Measuring national ‘emerging technology’ capabilities. Science and Public Policy, 29(3):189–200, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Propp:2006:OOO

[Pro06] Dr Tilo Propp. Opposite opinions in one book. Science and Public Policy, 33(3):232– 235, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Polt:2005:TCA

[PS05] Wolfgang Polt and Gerhard Streicher. Trying to capture additionality in Framework Programme 5 — main findings. Science and Public Policy, 32(5):367–373, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pelkonen:2008:APC

[PTW08] Antti Pelkonen, Tuula Teräväinen, and Suvi-Tuuli Waltari. Assessing policy coordination capacity: Higher education, science, and technology policies in Finland. Science and Public Policy, 35(4):241–252, May 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Raman:2001:KR

[Ram01] Sujatha Raman. Kuhn revisited. Science and Public Policy, 28(5):394–395, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Raman:2005:IPS

[Ram05] Dr Sujatha Raman. Institutional perspectives on sciencepolicy boundaries. Science and Public Policy, 32(6):418–422, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Raman:2007:LB

[Ram07] Dr Sujatha Raman. Lack of balance. Science and Public Policy, 34(8):601–602, October 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramlogan:2008:IRA

[Ram08]  Ronnie Ramlogan.  Insti-

tutional reform in American medical research. Science and Public Policy, 35(5):373–374, June 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:2001:EUD

[Rap01]  Brian Rappert.  European

Union drug regulation. Science and Public Policy, 28(6):480– 481, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:2003:ET

[Rap03] Brian Rappert. Engineering terror. Science and Public Policy, 30(2):140–141, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:2008:BRT

[Rap08] Brian Rappert. The benefits, risks, and threats of biotechnology. Science and Public Policy, 35(1):37–43, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rask:2003:PCP

[Ras03] Mikko Rask. The problem of citizens’ participation in Finnish biotechnology policy. Science and Public Policy, 30 (6):441–454, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ravetz:2001:SAK

[Rav01] Jerry Ravetz. Science advice in the knowledge economy. Science and Public Policy, 28(5): 389–393, October 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rayner:2003:DAA

[Ray03] Steve Rayner. Democracy in the age of assessment: Reflections on the roles of expertise and democracy in publicsector decision making. Science and Public Policy, 30(3):

163–170, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Revill:2008:LSN

[RD08] James Revill and Malcolm Dando. Life scientists and the need for a culture of responsibility: After education … what? Science and Public Policy, 35(1):29–35, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reeve:2005:EEU

[Ree05] Neville Reeve. On the evaluation of European Union research: The 2004 Five-Year Assessment. Science and Public Policy, 32(5):335–338, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Renn:2001:RSS

[Ren01] Ortwin Renn. The role of social science in environmental policy making: Experiences and outlook. Science and Public Policy, 28(6):427–437, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rennings:2005:ITM

[Ren05]  Dr Klaus Rennings.  Im-

plementing transition management. Science and Public Policy, 32(6):490–493, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Restivo:2005:BMI

[Res05] Sal Restivo. Bodies in motion in the information age. Science and Public Policy, 32(3):247– 250, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Robitaille:2002:EGS

[RG02] J-P Robitaille and B. Godin.

Evaluation of governments’ scientific output: a bibliometric profile of Canada. Science and Public Policy, 29(1):59–68, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rowe:2004:PPE

[RG04] Gene Rowe and John G. Gammack. Promise and perils of electronic public engagement. Science and Public Policy, 31 (1):39–54, February 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ronde:2007:AJS

[RH07] Patrick Rondé and Caroline Hussler. Is academic judgment sound? Evidence from technological agenda settings by experts. Science and Public Policy, 34(1):15–22, February 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reiss:2004:PEM

[RHL04] Thomas Reiss, Sibylle Hinze, and Iciar Dominguez Lacasa. Performance of European member states in biotechnology. Science and Public Policy, 31(5):344–358, October 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:2002:SUA

[Ric02] Jacques Richardson. Scientists under arms. Science and Public Policy, 29(2):147–149, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:2005:WWE

[Ric05] Jacques Richardson. Water, water, but not everywhere. Science and Public Policy, 32 (4):331–332, August 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:2006:USG

[Ric06] Jacques Richardson. Unesco surveys the globalization of science and technology. Science and Public Policy, 33(4):305– 306, May 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rigby:2007:AYH

[Rig07] John Rigby. From Aarhus to your house: The science and society interface. Science and Public Policy, 34(5):364–367, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rip:2005:PUS

[Rip05] Arie Rip. Public understanding of science. Science and Public Policy, 32(5):408–410, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radosevic:2009:PRF

[RL09] Slavo Radosevic and Benedetto Lepori. Public research funding systems in Central and Eastern Europe: between excellence and relevance: introduction to special section. Science and Public Policy, 36(9): 659–666, November 2009. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:2007:IF

[Rob07] Jody A. Roberts. Innovation for the future. Science and Public Policy, 34(10):756–757, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rod:2006:SUI

[Rod06] Dr Michel Rod. The sealing of university intellectual property boundaries and the ceiling of academic entrepreneurial tolerance. Science and Public Policy, 33(1):3–4, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rodriguez:2007:MZM

[Rod07] Victor Rodriguez. Merton and Ziman’s modes of science: The case of biological and similar material transfer agreements. Science and Public Policy, 34 (5):355–363, June 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rogers:2006:CLP

[Rog06] Juan D. Rogers. Community of learning, practice and collaboration. Science and Public Policy, 33(9):694–695, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rolfstam:2009:PPI

[Rol09] Dr Max Rolfstam. Public procurement as an innovation policy tool: The role of institutions. Science and Public Policy, 36(5):349–360, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosen:2000:PT

[Ros00] Paul Rosen. The personal and the technological. Science and Public Policy, 27(4):306–307, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosen:2001:PTS

[Ros01] Paul Rosen. Public trust in science and technology. Science and Public Policy, 28(4):323– 325, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ross:2002:SWH

[Ros02] David Ross. Scuppering the waves: How they tried to repel clean energy. Science and Public Policy, 29(1):25–35, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ross:2007:PSF

[Ros07]  Kerry Ross.  Providing

“thoughtful feedback”: Public participation in the regulation of Australia’s first genetically modified food crop. Science and Public Policy, 34(3):213– 225, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rod:2003:MSC

[RP03] Michel R. M. Rod and Stanley J. Paliwoda. Multi-sector collaboration: a stakeholder perspective on a government, industry and university collaborative venture. Science and Public Policy, 30(4):273–284, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Robinson:2006:CPE

[RT06] John Robinson and James Tansey. Co-production, emergent properties and strong interactive social research: the Georgia basin futures project. Science and Public Policy, 33 (2):151–160, March 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rubalcaba:2006:WPI

[Rub06] Luis Rubalcaba. Which policy for innovation in services? Science and Public Policy, 33 (10):745–756, December 2006.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ruegg:2007:QPE

[Rue07] Rosalie T. Ruegg. Quantitative portfolio evaluation of US federal research and development programs. Science and Public Policy, 34(10):723–730, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Russell:2006:WSB

[Rus06] Mark C. Russell. Why is the sky blue? Science and Public Policy, 33(8):622–623, October 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ryan:2008:ICN

[Rya08] Camille D. Ryan. Introducing complex networks to social scientists. Science and Public Policy, 35(9):697–699, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saetnan:2002:SDE

[Sae02] Ann Rudinow Saetnan. Scientific? Democratic? Effective? Towards an evaluation of Norway’s first medical consensus conference. Science and Public Policy, 29(3):201–220, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salazar:2008:BIE

[Sal08] Monica Salazar. Building institutions for effective policymaking. Science and Public Policy, 35(7):543–544, August 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanders:2007:EMY

[San07] Barry C. Sanders. Einstein:

a miraculous year one century later. Science and Public Policy, 34(6):449–450, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schoijet:2002:ULS

[Sch02]  Mauricio Schoijet.  Ultra-

left science policy and antimodernization in Argentina: Oscar Varsavsky. Science and Public Policy, 29(1):69–75, February 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schenkelaars:2005:RGC

[Sch05] Piet Schenkelaars. Regulating GM crops in The Netherlands: precaution as societalethical evaluation. Science and Public Policy, 32(4):309–316, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schneider:2009:GPS

[Sch09a] Ingrid Schneider. Governing the patent system in Europe: the EPO’s supranational autonomy and its need for a regulatory perspective. Science and Public Policy, 36(8):619– 629, October 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Scholten:2009:CII

[Sch09b] Peter Scholten. The coproduction of immigrant integration policy and research in The Netherlands: the case of the Scientific Council for Government Policy. Science and Public Policy, 36(7):561–573, August 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Scoones:2007:CPT

[Sco07] Ian Scoones. The contested politics of technology: Biotech in Bangalore. Science and Public Policy, 34(4):261–271, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

SiqueiraRapini:2009:UII

[SdV^{+}09] Márcia Siqueira Rapini, Eduardo da Motta e Albuquerque, Catari Vilela Chave, Leandro Alves Silva, Sara Gonçalves Antunes de Souza, Hérica Morais Righi, and Wellington Marcelo Silva da Cruz. University-industry interactions in an immature system of innovation: Evidence from Minas Gerais, Brazil. Science and Public Policy, 36(5): 373–386, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Seguin:2000:UBC

[Seg00] Dr Eve Seguin. The UK BSE crisis: Strengths and weaknesses of existing conceptual approaches. Science and Public Policy, 27(4):293–301, August 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Senker:2000:ISI

[Sen00] Jacqueline Senker. Introduction to a special issue on changing organisation and structure of European public-sector research systems. Science and Public Policy, 27(6):394–386, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Senker:2001:COP

[Sen01] Jacqueline Senker. Changing organisation of public-sector research in Europe — implications for benchmarking human resources in RTD. Science and Public Policy, 28(4): 277–284, August 2001. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Senker:2003:GSS

[Sen03] Peter Senker. The global STI system — science, technology and inequality. Science and Public Policy, 30(1):2–3, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Senker:2005:HCU

[Sen05] Jacqueline Senker. Hidden constraints on university research. Science and Public Policy, 32(6):493–495, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schmidt:2003:IDP

[SGL03] Evanthia Kalpazidou Schmidt, Ebbe Krogh Graversen, and Kamma Langberg. Innovation and dynamics in public research environments in Denmark: a research-policy perspective. Science and Public Policy, 30(2):107–116, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salazar:2004:DIS

[SH04] Mónica Salazar and Adam Holbrook. A debate on innovation surveys. Science and Public Policy, 31(4):254–266, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:2003:TCK

[Sha03] Naubahar Sharif. Technological change as knowledge change. Science and Public Policy, 30(2):142–143, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:2004:GIS

[Sha04] Naubahar Sharif. Getting inside science and technology studies. Science and Public Policy, 31(4):328–330, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shahin:2006:EHI

[Sha06a] Jamal Shahin. A European history of the Internet. Science and Public Policy, 33(9):681– 693, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:2006:ERD

[Sha06b] Naubahar Sharif. An examination of recent developments in Hong Kong’s innovation system: 1990 to the present. Science and Public Policy, 33(7): 505–518, August 2006. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:2007:CFI

[Sha07] Naubahar Sharif. Challenge faced by innovative firms. Science and Public Policy, 34(7): 532–533, August 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:2009:BDC

[Sha09] Naubahar Sharif. Bringing developing country innovation into the mainstream. Science and Public Policy, 36(2):166– 167, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shen:2005:DDC

[She05] Xiaobai Shen. A dilemma for developing countries in intellectual property strategy? Lessons from a case study of software piracy and Microsoft in China. Science and Public Policy, 32(3):187–198, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shove:2003:PAR

[Sho03]  Elizabeth Shove.  Princi-

pals, agents and research programmes. Science and Public Policy, 30(5):371–381, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simeonova:2004:CMR

[Sim04] Kostadinka Simeonova. Comment on Mirskaya and Rabkin’s article “Russian academic scientists in the first Post-Soviet decade”. Science and Public Policy, 31(6):490, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simeonova:2005:HAI

[Sim05] Dr Kostadinka Simeonova. Holistic approach to innovation in Poland. Science and Public Policy, 32(4):330–331, August 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simeonova:2006:RIB

[Sim06] Kostadinka Simeonova. Research and innovation in Bulgaria. Science and Public Policy, 33(5):351–363, June 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salo:2001:DPT

[SK01] Ahti Salo and Osmo Kuusi.

Developments in parliamentary technology assessment in Finland. Science and Public Policy, 28(6):453–464, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skladany:2004:DSE

[Skl04] Michael Skladany. Democratizing science and environmental health policy. Science and Public Policy, 31(1):81–82, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skoie:2000:DIM

[Sko00a] Hans Skoie. Diversity and identity: The merger of five research councils in Norway. Science and Public Policy, 27(2): 83–96, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skoie:2000:FIR

[Sko00b] Hans Skoie. Faculty involvement in research in mass higher education: current practice and future perspectives in the Scandinavian countries. Science and Public Policy, 27(6): 409–419, December 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Slavin:2001:DEP

[Sla01] Devorah Slavin. Devastating environmental poison. Science and Public Policy, 28(3):232– 233, June 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Slayton:2005:SIF

[Sla05] Rebecca Slayton. Science influenced by funding. Science and Public Policy, 32(2):168– 170, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Slipersaeter:2007:BPS

[SLD07] Stig Slipersæter, Benedetto Lepori, and Michael Dinges. Between policy and science: Research councils’ responsiveness in Austria, Norway and Switzerland. Science and Public Policy, 34(6):401–415, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanz-Menendez:2004:CCS

[SM04] Luis Sanz-Menéndez. Country case studies of S&T and innovation policies. Science and Public Policy, 31(4):328–330, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:2001:CEP

[Smi01] Keith Smith. Comparing economic performance in the presence of diversity. Science and Public Policy, 28(4):267–276, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:2003:TTR

[Smi03] Adrian Smith. Transforming technological regimes for sustainable development: a role for alternative technology niches? Science and Public Policy, 30(2):127–135, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:2005:CEL

[Smi05] Keith Smith. Changing economic landscape: Liberalisation and knowledge infrastructures. Science and Public Policy, 32(5):339–347, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sonnenwald:2003:COE

[Son03] Diane H. Sonnenwald. The conceptual organization: an emergent organizational form for collaborative R&D. Science and Public Policy, 30(4):261– 272, August 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sperling:2004:MPS

[Spe04] Stefan Sperling. Managing potential selves: stem cells, immigrants, and German identity. Science and Public Policy, 31 (2):139–149, April 2004. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spinardi:2000:PDD

[Spi00] Graham Spinardi. Prospects for the Defence Diversification Agency: Technology transfer and the UK defence research establishments. Science and Public Policy, 27(2):123–135, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spinardi:2008:BMD

[Spi08] Graham Spinardi. Ballistic missile defence and the politics of testing: the case of the US ground-based midcourse defence. Science and Public Policy, 35(10):703–715, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salter:2009:BGB

[SQ09] Brian Salter and Ren-Zong Qiu. Bioethical governance and basic stem cell science: China and the global biomedicine economy. Science and Public Policy, 36(1):47–59, February 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shove:2000:UUD

[SR00] Elizabeth Shove and Arie Rip. Users and unicorns: a discussion of mythical beasts in interactive science. Science and Public Policy, 27(3):175–182, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Srinivas:2000:RP

[Sri00] K. Ravi Srinivas. Reconstructing the problem. Science and Public Policy, 27(1):75–76, February 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Srinivas:2004:PDG

[Sri04] K. Ravi Srinivas. Promises and drawbacks of a global IPR regime. Science and Public Policy, 31(1):85–86, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Srinivas:2006:DGE

[Sri06] K. Ravi Srinivas. Does genetic engineering pose a significant risk? Science and Public Policy, 33(2):162–163, March 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schibany:2008:EIS

[SS08] Andreas Schibany and Gerhard Streicher. The European Innovation Scoreboard: drowning by numbers? Science and Public Policy, 35(10):717– 732, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schmoch:2009:WHU

[SS09] Dr Ulrich Schmoch and Dr Torben Schubert. When and how to use bibliometrics as a screening tool for research performance. Science and Public Policy, 36(10):753–762, December 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Siune:2005:IER

[SSA05] Karen Siune, Evanthia Kalpazidou Schmidt, and Kaare Aagaard. Implementation of European research policy. Science and Public Policy, 32(5):375– 384, October 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Standke:2006:STG

[Sta06] Dr Klaus-Heinrich Standke. Science and technology in global cooperation: The case of the United Nations and UN-ESCO. Science and Public Policy, 33(9):627–646, November 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stein:2002:GST

[Ste02a] Josephine Anne Stein. Globalisation, science, technology and policy. Science and Public Policy, 29(6):402–408, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stein:2002:STE

[Ste02b] Josephine Anne Stein. Science, technology and European foreign policy: European integration, global interaction. Science and Public Policy, 29(6): 463–477, December 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stevens:2003:CSI

[Ste03] Sharon McKenzie Stevens.

Community stewardship and inclusive decision-making. Science and Public Policy, 30(5): 383–385, October 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stein:2004:TEK

[Ste04] Josephine Anne Stein. Is there a European knowledge system? Science and Public Policy, 31 (6):435–447, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stewart:2009:PE

[Ste09] Jay Stewart. Painful econometrics. Science and Public Policy, 36(9):739–740, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stilgoe:2002:RTS

[Sti02] Jack Stilgoe. A reflexive turn in social science. Science and Public Policy, 29(2):150–151, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stilgoe:2003:CRD

[Sti03] Jack Stilgoe. Contributing to the risk debate. Science and Public Policy, 30(3):226, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stilgoe:2004:FAP

[Sti04] Jack Stilgoe. Finding out about public illness. Science and Public Policy, 31(1):77–78, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stilgoe:2005:CMP

[Sti05] Jack Stilgoe. Controlling mobile phone health risks in the UK: a fragile discourse of compliance. Science and Public Policy, 32(1):55–64, February 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stowe:2008:HMO

[Sto08] Lisa Stowe. A happy marriage? Organics and genetic engineering? Science and Public Policy, 35(9):696–697, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Suk:2008:ISI

[Suk08] Jonathan Suk. Introduction to special issue on biosecurity governance: Containing biological weapons, constraining biological research? Science and Public Policy, 35(1):2–4, February 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sandhya:2000:RCA

[SV00] G. D. Sandhya and S. Visalakshi. R&D capability and alliance formation in the pharmaceutical industry in India. Science and Public Policy, 27 (2):109–121, April 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schimank:2000:BHR

[SW00a] Dr Uwe Schimank and Markus Winnes. Beyond Humboldt? The relationship between teaching and research in European university systems. Science and Public Policy, 27 (6):397–408, December 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simmons:2000:CRI

[SW00b] Peter Simmons and Dr Gordon Walker. Contract research as interactive social science. Science and Public Policy, 27(3): 193–201, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Soubbotina:2009:NMT

[SW09] Tatyana Soubbotina and Charles Weiss. A new model of technological learning for Russia. Science and Public Policy, 36(4): 271–286, May 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Swedlow:2007:UBS

[Swe07] Brendon Swedlow. Using the boundaries of science to do boundary-work among scientists: Pollution and purity claims. Science and Public Policy, 34(9):633–643, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Siegel:2002:SRP

[SZ02] Donald S. Siegel and Vasilis Zervos. Strategic research partnerships and economic performance: Empirical issues. Science and Public Policy, 29(5):331–343, October 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saad:2008:THS

[SZM08] Dr Mohammed Saad, Dr Girma Zawdie, and Dr Chandra Malairaja. The Triple Helix strategy for universities in developing countries: The experiences in Malaysia and Algeria. Science and Public Policy, 35(6):431–443, July 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teo:2001:SMS

[TA01] Thompson S. H. Teo and James S. K. Ang. Singapore’s manufacturing sector as engine for economic growth: Past, present and future. Science and Public Policy, 28(5):361– 370, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Takeshita:2001:NCS

[Tak01] Chikako Takeshita. Nature conservation and social justice. Science and Public Policy, 28 (6):481–482, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tassava:2002:HAP

[Tas02] Christopher James Tassava.

How America produced ships. Science and Public Policy, 29 (3):239–240, June 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Taysom:2002:UDN

[Tay02] Sophie V. Taysom. Utopian and dystopian narratives of the future. Science and Public Policy, 29(2):151–152, April 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Torgersen:2005:AAB

[TB05] Helge Torgersen and Alexander Bogner. Austria’s agribiotechnology regulation: political consensus despite divergent concepts of precaution. Science and Public Policy, 32 (4):277–284, August 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tait:2002:LSI

[TCW02] Joyce Tait, Joanna Chataway, and David Wield. The life science industry sector: Evolution of agro-biotechnology in Europe. Science and Public Policy, 29(4):253–258, August 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teucher:2003:RGH

[Teu03a] Ulrich Teucher. Reading genes, heavens, and computers. Science and Public Policy, 30(4): 296–297, August 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teucher:2003:RER

[Teu03b] Ulrich Teucher. Relations of equal regard: Science and the humanities. Science and Public Policy, 30(6):465–466, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thorsteinsdottir:2000:PSR

[Tho00] Halla Thorsteinsdóttir. Publicsector research in small countries: does size matter? Science and Public Policy, 27(6): 433–442, December 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tura:2008:BIB

[THP08] Tomi Tura, Vesa Harmaakorpi, and Sanna Pekkola. Breaking inside the black box: towards a dynamic evaluation framework for regional innovative capability. Science and Public Policy, 35(10):733–744, December 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thune:2007:UIC

[Thu07]  Taran Thune.  University-

industry collaboration: The network embeddedness approach. Science and Public Policy, 34(3):158–168, April 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Talukder:2008:ETO

[TK08] Kana Talukder and Jennifer Kuzma. Evaluating technology oversight through multiple frameworks: a case study of genetically engineered cotton in India. Science and Public Policy, 35(2):121–138, March 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Theves:2007:CPP

[TLL07] Jean Thèves, Benedetto Lepori, and Philippe Larédo. Changing patterns of public research funding in France. Science and Public Policy, 34(6): 389–399, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Todt:2001:RP

[Tod01] Oliver Todt. Risk in perspective. Science and Public Policy, 28(5):399–400, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Toft:2005:DRA

[Tof05] Jesper Toft. Denmark’s regulation of agri-biotechnology: coexistence bypassing risk issues. Science and Public Policy, 32 (4):293–300, August 2005. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tran:2009:RSV

[Tra09] Dr Ngoc Ca Tran. Reaching out to society: Vietnamese universities in transition. Science and Public Policy, 36(2): 91–95, March 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tang:2001:EPS

[TS01] Dr Puay Tang and Dr Tom Sinclair. Exploitation practice in social science research. Science and Public Policy, 28(2):131– 137, April 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tsipouri:2001:CWB

[Tsi01] Lena Tsipouri. Can we benchmark the contribution of research and development investment to growth and competitiveness? Science and Public Policy, 28(4):295–302, August 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tsipouri:2004:IEC

[Tsi04] Lena J. Tsipouri. Innovation for European competitiveness and cohesion: Opportunities and difficulties of co-evolution. Science and Public Policy, 31 (6):465–474, December 2004.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Turnbull:2007:VID

[Tur07] David Turnbull. Of vital importance but disappointing. Science and Public Policy, 34 (7):531–532, August 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Turnhout:2009:EBO

[Tur09] Esther Turnhout. The effectiveness of boundary objects: the case of ecological indicators. Science and Public Policy, 36(5):403–412, June 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tickner:2003:PPD

[TW03] Joel A. Tickner and Sara Wright. The precautionary principle and democratizing expertise: a US perspective. Science and Public Policy, 30 (3):213–218, June 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teng-Zeng:2005:SSN

[TZ05] Frank K. Teng-Zeng. The same story or new directions? Science and technology within the framework of the African union and new partnership for Africa’s development. Science and Public Policy, 32(3):231– 246, June 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tyfield:2009:IIC

[TZC09] Dr David Tyfield, Yongguan Zhu, and Jinghua Cao. The importance of the ‘international collaboration dividend’: the case of China. Science and Public Policy, 36(9):723–735, November 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Uyarra:2007:DAC

[Uya07] Elvira Uyarra. Dimensions of the analysis of clusters. Science and Public Policy, 34(9):673– 676, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Valentin:2002:TRC

[Val02] Eva María Mora Valentín. A theoretical review of cooperative relationships between firms and universities. Science and Public Policy, 29 (1):37–46, February 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Valdiva:2009:WDI

[Val09] Walter D. Valdiva. Wage disparity and innovation. Science and Public Policy, 36(7):579– 580, August 2009. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanBuuren:2004:WJK

[vBE04] Arwin van Buuren and Dr Jurian Edelenbos. Why is joint knowledge production such a problem? Science and Public Policy, 31(4):289–299, August 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Viale:2002:ETH

[VC02] Riccardo Viale and Sergio Campodall’Orto. An evolutionary Triple Helix to strengthen academy-industry relations: Suggestions from European regions. Science and Public Policy, 29(3):154–168, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vera-Cruz:2008:VLC

[VCDM^{+}08] Alexandre O. Vera-Cruz, Gabriela Dutrénit, Griselda Martínez, Arturo Torres-Vargas, and Javier Ekboir. Virtues and limits of competitive funds to finance research and innovation: The case of Mexican agriculture. Science and Public Policy, 35(7):501–513, August 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vonortas:2007:ROF

[VD07] Nicholas S. Vonortas and Chintal A. Desai. ‘real options’ framework to assess public research investments. Science and Public Policy, 34(10):699– 708, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanderMeulen:2003:NRS

[vdM03] Barend van der Meulen. New roles and strategies of a research council: Intermediation of the principal-agent relationship. Science and Public Policy, 30(5):323–336, October 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanderSteen:2008:ASP

[vdS08] Martijn van der Steen. Ageing or silvering? Political debate about ageing in The Netherlands. Science and Public Policy, 35(8):575–583, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanEeten:2001:CAD

[vE01] Michel van Eeten. The challenge ahead for deliberative democracy: In reply to Weale. Science and Public Policy, 28 (6):423–426, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Velho:2005:SIL

[Vel05]  Léa Velho.  S&T institu-

tions in Latin America and the Caribbean: an overview. Science and Public Policy, 32(2): 95–108, April 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Verbruggen:2004:MPM

[Ver04] Aviel Verbruggen. A must for policy makers. Science and Public Policy, 31(1):84–85, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Viner:2006:SAR

[VGP06] Dr Neil Viner, Rod Green, and Philip Powell. Segmenting academics: resource targeting of research grants. Science and Public Policy, 33(3):166– 178, April 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gronbaek:2003:ERC

[vHG03] David J. v H. Gronbaek. A European Research Council: an idea whose time has come? Science and Public Policy, 30 (6):391–404, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Viseu:2004:WPM

[Vis04] Ana Viseu. Work practices made visible, connections to history invisible. Science and Public Policy, 31(3):247–248, June 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vogel:2008:FBA

[Vog08a] Kathleen M. Vogel. Framing biosecurity: an alternative to the biotech revolution model? Science and Public Policy, 35 (1):45–54, February 2008. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vogel:2008:IWD

[Vog08b] Kathleen M. Vogel. ‘Iraqi Winnebagos^{T M} of death’: imagined and realized futures of US bioweapons threat assessments. Science and Public Policy, 35(8):561–573, October 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vonortas:2000:TPU

[Von00] Nicholas S. Vonortas. Technology policy in the United States and the European Union: Shifting orientation towards technology users. Science and Public Policy, 27(2):97–108, April 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Voss:2007:IPG

[Voß07] Jan-Peter Voß. Innovation processes in governance: The development of ‘emissions trading’ as a new policy instrument. Science and Public Policy, 34(5):329–343, June 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vonortas:2007:ENL

[VSZ07] Nicholas S. Vonortas, Michael Stampfer, and Klaus Zinöcker. Evidence never lies: Introduction to a special issue on New Frontiers in Evaluation. Science and Public Policy, 34(10): 679–680, December 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vonTunzelmann:2004:TPE

[vTN04] Nick von Tunzelmann and Sussan Nassehi. Technology policy, European Union enlargement, and economic, social and political sustainability. Science and Public Policy, 31(6):475– 483, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wagner:2002:EPS

[Wag02] Caroline S. Wagner. The elusive partnership: Science and foreign policy. Science and Public Policy, 29(6):409–417, December 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watson:2000:MIS

[Wat00] Tony J. Watson. Management and interactive social science: Critical participative research. Science and Public Policy, 27 (3):203–210, June 2000. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Waterton:2005:SCB

[Wat05] Claire Waterton. Scientists’ conceptions of the boundaries between their own research and policy. Science and Public Policy, 32(6):435–444, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weale:2001:SAD

[Wea01] Albert Weale. Science advice, democratic responsiveness and public policy. Science and Public Policy, 28(6):413–421, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weir:2002:HSI

[Wei02] Tony Weir. Hard-to-shake ideas shape energy policy. Science and Public Policy, 29(4): 308–309, August 2002. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weill:2003:CCB

[Wei03] Claire Weill. Can consultation of both experts and the public help developing public policy? Some aspects of the debate in France. Science and Public Policy, 30(3):199–203, June 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wetmore:2007:ISI

[Wet07] Jameson M. Wetmore. Introduction to special issue on science, policy and social inequity. Science and Public Policy, 34(2):83–84, March 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Whitesides:2002:PMD

[Whi02] Greg Whitesides. Are the poor medically disenfranchised? Science and Public Policy, 29(3):234–240, June 2002. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Whitesides:2003:DSM

[Whi03] Greg Whitesides. In defense of the silent majority — the animals. Science and Public Policy, 30(1):67–68, February 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Whitman:2007:GN

[Whi07] Jim Whitman. The governance of nanotechnology. Science and Public Policy, 34(4): 273–283, May 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkin:2004:HWY

[Wil04] James G. Wilkin. How would you like your cluster? Science and Public Policy, 31(6):492– 493, December 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:2006:CCC

[Wil06]  Amanda Williams.  Cana-

dian contribution to the cluster craze. Science and Public Policy, 33(1):84–86, February 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Winskel:2007:GTI

[Win07]  Mark Winskel.  Govern-

ing technology by intellectual property rights. Science and Public Policy, 34(6):450–452, July 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wen:2001:TDC

[WK01] Jiang Wen and Shinichi Kobayashi. Technology diffusion in China: some new evidence in computer-aided design. Science and Public Policy, 28(1):41–47, February 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Winskel:2006:EPI

[WMWW06] Dr Mark Winskel, Dr Andrew McLeod, Dr Robin Wallace, and Robin Williams. Energy policy and institutional context: Marine energy innovation systems. Science and Public Policy, 33(5):365–376, June 2006. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wolfe:2007:CCI

[Wol07] David A. Wolfe. Critical contribution to innovation studies. Science and Public Policy, 34 (9):671–673, November 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Woolgar:2000:SBI

[Woo00] Steve Woolgar. Social basis of interactive social science. Science and Public Policy, 27(3): 165–173, June 2000. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Worner:2001:DII

[WR01] Stefan Wörner and Thomas Reiss. The direct and indirect impacts of new technologies on employment: The example of the German biotechnology sector. Science and Public Policy, 28(5):371–380, October 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wright:2003:RF

[Wri03a] Katharine Wright. Regulating the future. Science and Public Policy, 30(6):464–465, December 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wright:2003:VET

[Wri03b] Katharine Wright. The varieties of experienced time. Science and Public Policy, 30(4): 295–300, August 2003. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Woodhouse:2007:SPR

[WS07] Edward Woodhouse and Daniel Sarewitz. Science policies for reducing societal inequities. Science and Public Policy, 34 (2):139–150, March 2007. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watanabe:2003:NIP

[WT03] Chihiro Watanabe and Shinji Tokumasu. National innovation policies in an IT society: The myth of technology policies focusing on supply sides. Science and Public Policy, 30 (2):70–84, April 2003. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Xielin:2007:VNI

[Xie07] Liu Xielin. Value of national innovation system approach. Science and Public Policy, 34 (10):754–756, December 2007. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zehr:2005:CBW

[Zeh05] Stephen Zehr. Comparative boundary work: US acid rain and global climate change policy deliberations. Science and Public Policy, 32(6):445–456, December 2005. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zeiss:2004:SIW

[Zei04] Ragna Zeiss. Seams in an institutional web. Science and Public Policy, 31(1):82–84, February 2004. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zito:2001:ECE

[Zit01]  Anthony R. Zito.  Epis-

temic communities, European Union governance and the public voice. Science and Public Policy, 28(6):465–476, December 2001. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zhou:2008:EUC

[ZP08] Chunyan Zhou and Xu-Mei Peng. The entrepreneurial university in China: nonlinear paths. Science and Public Policy, 35(9):637–646, November 2008. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).